### v3

# A Mixed-Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling

Abstract.
The Rust programming language restricts aliasing to provide static safety guarantees. However, in certain situations, developers need to bypass these guarantees by using a set of unsafe features. If they are used incorrectly, these features can reintroduce the types of safety issues that Rust was designed to prevent. We seek to understand how current development tools can be improved to better assist developers who find it necessary to interact with unsafe code. To that end, we study how developers reason about foreign function calls, the limitations of the tools that they currently use, their motivations for using unsafe code, and how they reason about encapsulating it. We conducted a mixed-methods investigation consisting of semi-structured interviews with 19 developers, followed by a survey that reached an additional 160 developers. Our participants were motivated to use unsafe code when they perceived that there was no alternative, and most avoided using it. However, limited tooling support for foreign function calls made participants uncertain about their design choices, and certain foreign aliasing and concurrency patterns were difficult to encapsulate. To overcome these challenges, Rust developers need verification tools that can provide guarantees of soundness within multi-language applications.
Rust, unsafe, memory safety, threat safety, interoperation, foreign functions, motivations, tooling, interview, survey, mixed-methods
copyright: noneccs: Software and its engineering
1. Introduction
The Rust programming language has made rapid headway as a safe alternative for systems programming due to its static safety guarantees. However, it is still the case that critical systems are written in languages that do not statically prevent memory and thread safety issues. Developers who interoperate with these programs must use Rust’s unsafe features, which bypass its safety restrictions. To minimize the risks of using these features, the Rust community advocates for keeping unsafe code minimal, well documented, and hidden beneath a safe interface (Astrauskas et al., 2020). However, these practices alone are not enough to ensure that Rust fulfills its promise of static safety. Errors in “safe” encapsulations of unsafe implementations have been a significant source of bugs and security vulnerabilities in the Rust ecosystem (Qin et al., 2020; Xu et al., 2021; Zheng et al., 2023). Applications that rely on interoperation or direct access to system resources may expose developers to sources of unsafety that cannot easily be minimized. Making matters worse, Rust’s semantics are still evolving, so few tools are capable of fully testing or verifying these applications.
Interoperation is a common use case for unsafe code (Evans et al., 2020; Astrauskas et al., 2020; Fulton et al., 2021; Höltervennhoff et al., 2023), but most Rust-specific development tools lack support for foreign function calls. Safe interoperation will be crucial to Rust’s future, since both industry and regulatory stakeholders have advocated for adopting Rust as a replacement for existing systems languages (Wallach, 2024; Rebert and Kern, 2024). Automated translation tools are increasingly effective (Larsen, 2018), but for the moment, interoperation is the most practical solution for adopting Rust within preexisting, large-scale systems (Rebert and Kern, 2024). It can be challenging to use Rust correctly in this context, because it has significant differences from other languages (Höltervennhoff et al., 2023; McCormack et al., 2024). A detailed account of these challenges is missing, and we need this to be able to develop effective tooling for safe interoperation. To provide this perspective, we ask the following research questions:
RQ1 (Interoperation). How do Rust developers reason about memory safety across foreign function boundaries? RQ2 (Tooling). What tools do Rust developers use when contributing to applications that include unsafe code, and how could tooling be improved?
To design effective interventions for developers who use unsafe code, we need to understand what motivates them to use these features and how they reason about the soundness of their designs. However, most large-scale studies of these topics have analyzed patterns in source code instead of directly engaging with developers (Evans et al., 2020; Astrauskas et al., 2020). Interviews and surveys have provided a missing qualitative perspective, but they tend to have a limited focus on unsafe code, a small sample size, or different goals (Fulton et al., 2021; Höltervennhoff et al., 2023). None of these studies have directly investigated how developers use foreign functions—we aim to fill this knowledge gap. In addition, to serve as guidance for the design of future tools, we seek to generalize existing qualitative results to a broader population of developers. To that end, we ask the following additional research questions:
RQ3 (Motivations). What are Rust developers’ motivations for using unsafe code? RQ4 (Encapsulation). How do Rust developers reason about encapsulating unsafe code?
Contribution
We provide the results of an exploratory, mixed method study that addresses the research questions above. We conducted semi-structured interviews with 19 developers and used the results to create a community survey that had 160 valid responses. The majority of our participants used unsafe code to call foreign functions, and they described several aliasing and concurrency patterns that made interoperation difficult. Many developers had used Miri, an interpreter, to find bugs, but most were deterred from using it again due to its slow performance and lack of support for key features. Participants prioritized keeping unsafe code minimal, encapsulated, and well documented, but they rarely audited their dependencies and relied on ad-hoc reasoning to justify their decisions. This correctness gap can be addressed by creating verification and bug finding tools that have robust support for multi-language applications and are capable of providing guarantees of soundness.
Outline
In Section 0, we describe the role of unsafe code in the Rust ecosystem and describe the categories of tools that Rust developers can use for testing, bug finding, and verification. In Section 0, we provide an overview of related studies of unsafe Rust. We describe our methodology in Section 0. We present our results in Section 0 and discuss their implications in Section 0. We indicate threats to validity in Section 0 before concluding in Section 0. Our replication package is available via GitHub [^1] https://github.com/icmccorm/unsafe-study and FigShare. [^2] https://figshare.com/s/aacea766be06aa6deb97
2. Background
Rust’s safety guarantees build on the notion of ownership. Values with “copy” semantics do not have an owner and can be copied freely, while values with “move” semantics have a unique owner. Assignment transfers ownership from one alias to another, but ownership can also be borrowed by creating a reference. Each reference has a lifetime, which is the duration of the program where it is used. Rust’s borrow checker uses this property to enforce that a value can have either a unique, mutable reference or many immutable references, but not both simultaneously (Crichton, 2020). A value’s type can implement traits, which define how it behaves. Certain traits affect a value’s capability to be owned and borrowed. The Copy trait gives a value copy semantics, Send allows values to be transferred between threads, and Sync allows multiple threads to reference a value.
Rust’s aliasing rules are conservative and are not universally compatible with certain idioms of systems programming. When these restrictions become burdensome, developers can use the unsafe keyword to enable features that escape the constraints of Rust’s borrow checker. Within an unsafe context, users can dereference raw pointers, execute inline assembly instructions, modify static mutable state, implement unsafe traits, and call unsafe functions (Klabnik and Nichols, 2022). There are a variety of use cases for these features (Astrauskas et al., 2020). In certain situations, unsafe code can be necessary to implement design patterns that are fundamentally incompatible with the restrictions of safe references, such as shared mutable state. Foreign function calls, inline assembly, and intrinsics allow developers to leverage hardware acceleration and interact with external resources that are not constrained by the borrow checker. These features can also improve performance through bypassing certain runtime assertions, like bounds checks.
However, if unsafe code is used inappropriately, it can reintroduce the same types of safety issues that Rust was designed to prevent, such as data races and accesses-out-of-bounds. Rust’s aliasing restrictions also introduce new forms of undefined behavior. Though references are constrained by the borrow checker, they can also be cast into raw pointers and used freely in unsafe contexts. Rust’s compiler is allowed to assume that programs which use raw pointers still adhere to the rules of Rust’s aliasing model. If a program breaks these rules, then it may be optimized incorrectly, introducing differences in behavior that can constitute security vulnerabilities (Crichton and Vranken, 2023).
2.1. Best Practices & Tooling
Several best practices can help developers mitigate the risks of using unsafe code. When implementing an unsafe feature, developers are encouraged to follow the “interior unsafety” pattern by keeping it minimal, well documented, and hidden beneath a safe interface (Qin et al., 2020). Ideally, this interface should make it impossible for users to trigger undefined behavior from safe code. However certain safety properties are difficult or impossible to encode into Rust’s type system without leveraging external verification tools. If a function or trait places the burden of correctness on its users, then it should be labelled unsafe and documented with any necessary invariants. These and other guidelines are provided by multiple community resources. The Rust Reference (Rust Community, 2023a) and The Rust Programming Language (Klabnik and Nichols, 2022) each include relevant sections on unsafe code, but the Rustonomicon (Rust Community, 2023b) provides the most comprehensive overview. The Unsafe Code Guidelines Reference (Rust Community, 2023c) was intended to be an in-depth guide to the semantics of unsafe features, but it is incomplete and mostly unmaintained. However, the “Issues” section of its repository is still actively used to discuss Rust’s semantics.
In addition to these best practices, dynamic analysis is also regularly used for validating unsafe code. Miri (Rust Community, 2024f), a Rust interpreter, is currently the only method for detecting violations of Rust’s evolving aliasing models, so it is seen as the “de facto” (Gäher et al., 2024) bug finding tool for Rust. Currently, Miri can detect violations of two different models of Rust’s semantics. Stacked Borrows (Jung et al., 2019) was the first iteration of Rust’s aliasing rules, and it is enabled by default. Developers can also switch to using the newer Tree Borrows (Villani et al., 2023) model by providing a configuration flag. Neither of the two models is canonical. Shuttle (Rust Community, 2024h) and Loom (Rust Community, 2024e) are Rust-specific fuzzing tools that permute the scheduling of threads to help detect concurrency bugs. Rust is also compatible with several language agnostic tools, including libFuzzer (LLVM Project, 2022), the LLVM project’s sanitizers (LLVM Project, 2024), and Valgrind (Nethercote and Seward, 2007).
Static analysis is also useful in this context. Clippy (Rust Community, 2024d), a linting tool, can hint at improper use of unsafe features, but it is not a true bug finding tool. The Rust ecosystem does not yet have a de facto solution for static bug finding, but several research efforts have produced promising results. Rudra (Bae et al., 2021) can detect use-after-free errors when a Rust application panics, while MIRChecker (Li et al., 2021) can detect panics and duplicated ownership. The Rust Analysis Platform (RAP) (Cui et al., 2023a, 2024) can detect memory leaks and use-after-free errors. MIRAI aims to provide a unified solution for both static analysis and verification with abstract interpretation (Endor Labs, Rust Community, 2024). Other extensions leverage community knowledge to help enforce safety properties. These have typically been developed as plugins for cargo, Rust’s build tool. Both cargo-audit (Rust Community, 2024g) and cargo-deny (Rust Community, 2024a) detect security vulnerabilities in a project’s dependencies; cargo-deny is more extensible. The cargo-geiger (Rust Community, 2024b) plugin provides statistics on unsafe code, and cargo-vet (Rust Community, 2024c) can be used to track if code has been reviewed by a trusted party.
Verification has been a uniquely active area of research within the Rust community. Here, we provide a high-level overview using several representative tools. Each category of tooling has a different level of compatibility with unsafe code. The first category includes Prusti (Astrauskas et al., 2019), Creusot (Denis et al., 2022), and Aeneas (Ho and Protzenko, 2022), each deductive verifiers, and Flux (Lehmann et al., 2023), a refinement type checker. These tools rely on Rust’s aliasing invariants to encode the properties of programs into their logic, so they do not support unsafe code by construction. The second category includes RefinedRust (Gäher et al., 2024) and Gillian-Rust, which exclusively target unsafe code. RefinedRust encodes encodes programs into a separation logic defined in Iris, which is automatically verified using the Lithium engine. Gillian-Rust relies on symbolic execution. Both approaches are informed by the RustBelt (Jung et al., 2017) model. Remaining tools are targeted at verifying arbitrary Rust programs, but they use unique approaches that provide limited support for unsafe features. Verus (Lattuada et al., 2023), a deductive verifier, uses linear ghost types to verify heap accesses through raw pointers. Kani, a model checker (VanHattum et al., 2022), is capable of detecting spatial and temporal memory safety issues. Though each of these tools have varying levels of support for unsafe code, none can validate programs against Miri’s evolving definitions of Rust’s aliasing model, and only Kani has support for foreign function calls.
3. Related Work
Table 1. A comparison table outlining how we contribute to prior studies of unsafe Rust. The column “Focus” indicates whether a work exclusively focused on unsafe code (direct) or studied the Rust ecosystem more broadly (indirect).
{tblr}
colspec=Q[l,m]Q[l,m]Q[l,m]$Q[c,m]Q[c,m]$ Work & Focus Methodology RQ1 RQ2 RQ3 RQ4
Astrauskas (Astrauskas et al., 2020) Direct Code Analysis
Cui (Cui et al., 2023b) Direct Code Analysis, Survey (n=50)
Evans (Evans et al., 2020) Direct Code Analysis, Survey (n=20)
Fulton (Fulton et al., 2021) Indirect Interviews (n=16), Survey (n=178)
Höltervennhoff (Höltervennhoff et al., 2023) Direct Interviews (n=26)
McCormack (McCormack et al., 2024) Direct Code Analysis
Ozdemir (Ozdemir, 2016) Direct Code Analysis
Qin (Qin et al., 2020) Direct Code Analysis
Xu (Xu et al., 2021) Direct Code Analysis
Zeng (Zeng and Crichton, 2018) Indirect Content Analysis, Interviews (n=3)
Zhu (Zhu et al., 2022) Indirect Content Analysis, Survey (n=101)
Zheng (Zheng et al., 2023) Direct Code Analysis
[…]
[c=1]r Combined:
{tblr}colspec=Q[c,m]Q[l,m] Label Definition
Provides relevant evidence, but does not directly answer our research question.
Gives a partial answer to our research question, but we provide additional parts.
Answers the same research question—we evaluate if their results generalize to a larger population.
Answers the same research question.
Does not address our research question.
Prior studies have used a variety of methods to investigate how Rust developers use unsafe code. Here, we provide an overview of results that are relevant for each of our research questions. In Section 3.1, we describe results from studies that examined how unsafe code is written. We describe studies that engaged directly with Rust developers in Section 0. In Section 3.3, we classify every relevant prior study and outline how we make novel contributions in our answers to each research question.
3.1. Studies of Code
Distribution
Surveys indicate that unsafe code is prevalent throughout the Rust ecosystem, but developers tend to use it minimally and hide it beneath safe interfaces. In Rust, packages and libraries are referred to as “crates,” and they are published on crates.io. In July of 2016, Ozdemir (Ozdemir, 2016) found that 29% of all published crates used unsafe features, but most unsafe blocks were relatively small and tightly scoped to include only unsafe operations. Evans et al. (Evans et al., 2020) observed similar patterns in September of 2018. The number of published crates had grown by more than ten times, but 29% still directly used unsafe code. They also observed that 38% of crates did not use any unsafe features, but they directly or indirectly depended on third-party crates that did. In January of 2020, Astrauskas et al (Astrauskas et al., 2020) found that the proportion of crates using unsafe features had shrunk to 23.6%. However, most instance of unsafe code were relatively straightforward and encapsulated. The majority of unsafe functions called by libraries were standard (e.g. not closures, function pointers, or trait functions) and defined within the same crate or provided by Rust’s standard library. More than 80% of unsafe functions were publicly visible, but most were exposed by crates that provided raw bindings to foreign libraries.
Use Cases
Studies have also examined applications to identify use cases for unsafe code. Qin et al. (Qin et al., 2020) studied a random sample of 600 instances of unsafe code from 10 popular Rust libraries, as well as 250 instances within safe encapsulations provide by Rust’s standard library. They identified three use cases for these operations; 42% were related to interoperation, 22% were for improving performance, and 14% involved sharing memory between threads. Astrauskas et al. (Astrauskas et al., 2020) used “anecdotal examples” and snippets of source code to create a taxonomy of six purposes for using unsafe code. These included implementing data structures with complex sharing, overcoming the incompleteness of Rust’s type system, emphasizing contracts and invariants, calling foreign functions, using intrinsics or inline assembly, and improving performance.
Bugs & Vulnerabilities
Discrepancies between the signatures of safe APIs and their interior unsafe implementations have been a significant source of bugs and vulnerabilities within the Rust ecosystem. Qin et al. (Qin et al., 2020) examined 170 bugs and vulnerabilities from ten high-profile Rust libraries. The Memory safety errors that they encountered were most frequently triggered from within safe contexts due to errors in unsafe implementations elsewhere. They also identified 19 examples of improper encapsulations of unsafe code in both the third-party libraries that they sampled and within Rust’s standard library. Xu et al. (Xu et al., 2021) exclusively studied registered vulnerabilities and identified similar errors in the encapsulations of unsafe code. Zheng et al. (Zheng et al., 2023) conducted the broadest empirical study of security issues in the Rust ecosystem to date. They examined 433 unique vulnerabilities that were discovered between 2014 and 2022, identifying 300 affected repositories. The portions of code that were affected by a vulnerability were significantly more likely to include unsafe blocks or functions. The overwhelming majority of all vulnerabilities affected safe functions.
It can be difficult to correctly implement an encapsulation because unsafe implementations require a variety of safety invariants. Cui et al. (Cui et al., 2023b) manually reviewed 416 unsafe APIs in Rust’s standard library and identify 19 categories of preconditions and postconditions. They evaluated their taxonomy with a survey of 90 Rust developers, and also categorized 186 vulnerabilities based on the safety properties that were violated. The most common violations were related to incorrect uses of Send or Sync traits, incomplete initialization, or accesses-out-of-bounds.
Interoperation
Interoperation has consistently been one of the most prevalent use cases for unsafe code. Astrauskas et al. (Astrauskas et al., 2020) found that 44.6% of all publicly-exported unsafe functions were static bindings to foreign functions. Other surveys of the Rust ecosystem have similar conclusions; Ozdemir (Evans et al., 2020) found that 30% of unsafe blocks were entirely dedicated to foreign function calls, and Evans et al. (Evans et al., 2020) found that 22.5% of unsafe function calls were to foreign functions. Few studies have examined errors that occur due to the interaction between Rust and other languages. Cui et al. (Cui et al., 2023b) explicitly excluded foreign functions from their investigation of safety properties. Xu et al. (Xu et al., 2021) found 12 examples of bugs and vulnerabilities that involved foreign function calls. They were related to application-specific invariants, data races, invalid alignment, and inconsistent layouts. The authors provided few details on the nature of these bugs, indicating that they were “straightforward” and not Rust-specific.
Only two studies have identified bugs that meaningfully involve Rust’s unique aliasing restrictions. In 2023, a set of anonymous authors categorized 22 bugs from five popular Rust encapsulations for foreign libraries (Anonymous, 2023). Most of the errors that they investigated involved incorrect handling of exceptions and heap objects that were improperly shared across foreign boundaries. They identified two bugs related to Rust’s aliasing restrictions, but they did not describe these bugs in terms of Rust’s operational semantics. McCormack et al. (McCormack et al., 2024) are the only group to have studied violations of Rust’s aliasing models in multi-language applications. They created a prototype extension to Miri (Rust Community, 2024f) and used it to find 48 bugs in Rust libraries that called foreign functions. Half of these bugs were violations of Rust’s Tree Borrows (Villani et al., 2023) aliasing model. A few were found in high-profile libraries, including one maintained by the Rust project.
3.2. Studies of Developers
Content Analysis
Studies of online discourse have also provided relevant insights into encapsulation and interoperation. Zeng and Crichton (Zeng and Crichton, 2018) studied the challenges associated with adopting Rust. They interviewed three Rust contributors and analyzed a set of articles and comments posted on Hacker News and the Rust subreddit (/r/rust). They indicated that it was more common for developers to discuss rewriting entire applications in Rust instead of porting them incrementally, suggesting that the difficulties of interoperation may prevent developers from taking this route. They also identified that developers who struggled with Rust’s aliasing rules were often pointed toward safe encapsulations of unsafe operations. Zhu et al. (Zhu et al., 2022) observed a similar phenomenon on Stack Overflow. They used a mixed-methods approach to understand how developers reason about Rust’s aliasing restrictions. In the first stage of their investigation, they examined a random sample of 100 Stack Overflow questions to identify situations where developers asked for help fixing code that had been rejected by the borrow checker. Only three of the 110 borrowing errors in these questions were fixed by writing unsafe code. Developers were referred to safe alternatives in every other situation.
Qualitative & Mixed-Methods Studies
Fulton et al. (Fulton et al., 2021) studied the benefits and drawbacks of adopting Rust as a general-purpose systems programming language. They used an exploratory, mixed-methods design that is similar to ours. First, they interviewed 16 developers in industry who had advocated for using Rust. Then, they distributed a community survey covering the same topics as the interviews, and it received 178 responses. The majority of their participants struggled with Rust’s steep learning curve, but error messages from the compiler and community resources made it easy for them to find solutions to their problems. Participants felt somewhat more productive using Rust, and they were more confident that their implementations were correct. Fulton et al. (Fulton et al., 2021) only briefly discussed unsafe features. Their participants typically used unsafe code to call foreign functions, improve performance, and interact directly with hardware or the operating system.
Höltervennhoff et al. (Höltervennhoff et al., 2023) are the only other group that we are aware of who have used qualitative methods to study how developers use unsafe code. They interviewed 26 developers with experience using unsafe Rust. Except for three who piloted the interview protocol, all of their participants had contributed unsafe code to open-source projects included in the unofficial “Awesome Rust” list (Community, 2022) of applications. We cover similar topics, but our research questions reflect different, complementary goals. They focused on understanding how developers contend with the security implications of using unsafe code, while we are focused on interoperation and tool use. Most of their participants relied on common sense and test coverage to evaluate the security of their code, and few had contributed to projects with formal policies on security or code review. However, most indicated that that they would review unsafe code more carefully than safe code, ensuring that it upheld all necessary safety invariants. Few participants had ever encountered security issues caused by unsafe code, and none of the issues that participants did experience were reported as critical vulnerabilities. We provide additional details on the results that are relevant for answering our research question in Section 0 and throughout Section 0.
3.3. Our Contributions
In Table 0, we use multiple descriptors to classify each of the relevant studies that we presented in previous sections. We label the methods used by each study as code analysis, content analysis, surveys, and interviews. We define code analysis to include mining software repositories or gathering information from bug reports and vulnerabilities. Content analysis includes examining any other type of preexisting online content, such as forum posts, blog posts, articles, or online discussions, to gather insights about Rust developers. Studies with a direct focus on unsafe code are similar to ours in that each of their core research questions are targeted at understanding how developers use unsafe features. Studies with an indirect focus are more broadly concerned with Rust but also happen to provide relevant insights for answering our research questions.
We also use several labels to indicate whether each study can fully answer our research questions. The empty circle ‘ ’ indicates that a study does not answer a question. The label ‘ ’ indicates that a study provides relevant evidence, but it does not address the same question. The label ‘ ’ indicates that a study partially answers our question, but we provide missing parts. The label ‘ ’ indicates that a study fully answers our question, but it does not provide findings that generalize to a population that is comparable to our survey population. Now, we describe our rationale for the labels that we chose and we indicate how we provide new results in our answers to each question.
Interoperation (RQ1)
Studies of bugs, vulnerabilities, and the distribution of unsafe code all indicate that foreign function calls are a potentially significant unchecked source of undefined behavior in the Rust ecosystem. However, none of these studies provide insights into the challenges that developers face when implementing these applications. Most of the participants studied by Fulton et al. (Fulton et al., 2021) indicated that C and Python were easy to interoperate with, but C++ was difficult to handle due to complex aliasing. Participants interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023) also had mixed experiences with interoperation; four described it as being easy, but two struggled with “memory management, e.g. between programming languages.” These studies can only partially answer our research question ( ), since neither provides a detailed description of the patterns that lead to these difficulties, other than that they involved aliasing and initialization (Höltervennhoff et al., 2023; Fulton et al., 2021). We have an explicit focus on interoperation as one of our core research questions, and we provide a missing perspective on the challenges of using Rust in multi-language applications.
Tooling (RQ2)
Rust is a relatively new programming language, so the majority of research efforts have focused on developing new tools instead of evaluating existing ones. Zeng and Crichton (Zeng and Crichton, 2018) hypothesized that a lack of promotion for tooling could prevent Rust from being widely adopted as a systems language. However, their work was published prior to the creation of most Rust-specific bug finding tools. Höltervennhoff et al. (Höltervennhoff et al., 2023) are the only other group that we are aware of who have examined how developers use tooling to assist with writing unsafe code. Key development tools were somewhat underused among their interview population—only 7 of their 26 participants had mentioned using Miri. Since this is a small sample size, it is unclear whether these results are indicative of a broader trend in the Rust community, so they can only partially answer our second research question ( ). We also cover debuggers, static verifiers, and tools for generating bindings to foreign functions, which were not mentioned by Höltervennhoff et al. (Höltervennhoff et al., 2023).
Motivations (RQ3)
Several prior studies have conflated developers’ motivations for using unsafe code with features-oriented use cases. For example, Astrauskas et al. (Astrauskas et al., 2020) include both performance and compiler intrinsics as “motivations” for using unsafe code. In their survey of 20 developers, Evans et al. (Evans et al., 2020) defined performance, ease-of-use, and a variety of specific operations as “reasons” for using unsafe code. Fulton et al. (Fulton et al., 2021) include both performance, interaction with hardware, and foreign function calls as “uses” for unsafe code. We define “motivations” as feature agnostic reasons for using unsafe code. This perspective is similar to Höltervennhoff et al. (Höltervennhoff et al., 2023), who identified necessity and performance, and efficiency as “motivations,” but the separation between features and higher-level reasons was not an explicit design consideration of their study. When viewed together, these results answer our third research question ( ). However, we provide a novel contribution by replicating their findings and evaluating whether they generalize to a large population of developers. Additionally, Höltervennhoff et al. (Höltervennhoff et al., 2023) study developers’ motivations as part of examining the broader topic of the “common practices for deciding for and implementing” unsafe code. We dedicate an entire research question to this topic, so we are able to provide significantly more detail on how specific design decisions relate to each motivation.
Encapsulation (RQ4)
Surveys of the Rust ecosystem suggest that developers typically follow best practices for minimizing and encapsulating unsafe code, but they struggle with ensuring that their interfaces are correct in all situations. Höltervennhoff et al. (Höltervennhoff et al., 2023) are the only prior study to have examined developers experiences with designing applications that include unsafe code, and their results reflect the patterns observed surveys of code and bug reports. The majority of their participants would avoid using unsafe code whenever possible. When unsafe code was necessary, they ensured that it was encapsulated beneath a safe interface of their own design or provided by an external library. However, the majority of participants felt that it was difficult to write unsafe code correctly; they had a “nagging feeling about it”, and half were unsure or undecided about its correctness. Similar to RQ3, these results also answer our research question ( ), but we provide a unique contribution by evaluating whether their results generalize to our survey population. We also use the set of motivations we defined in our answer to RQ3 to investigate developers’ motivations for exposing unsafe APIs, which is not addressed by Holtervennhoff et al. (Höltervennhoff et al., 2023).
4. Methodology
We adapted a sequential, exploratory mixed-methods design (Creswell et al., 2003) consisting of three phases. We describe the first phase in Section 0, where we used a screening survey to recruit developers for semi-structured interviews. In the second phase, after interviewing each eligible candidate, the first three authors conducted a thematic analysis (Miles et al., 2020) of the interview transcripts. We describe this process in Section 0. We describe the third phase in Section 0,where we used the results of our qualitative analysis to create a community survey. We chose this combination of methods because it allowed us to evaluate the generalizability of prior qualitative findings and to provide a missing qualitative perspective on interoperation and tool use.
Figure 1. A sample of materials relevant to each stage of our methodology. This includes the screening survey, interview protocol, an example of a theme from our codebook, and a survey question connected to that theme. Notes [in brackets] are documentation and were not presented to participants.
(a) The screening survey.
  1. 1.
    Do you call foreign functions from Rust?
  2. 2.
    Do you call Rust functions from other languages?
  3. 3.
    Do you use Rust’s intrinsics?
  4. 4.
    Do you use system calls in Rust?
  5. 5.
    Which of the following reference and memory container types have you converted to raw pointers or used to contain raw pointers?
  6. 6.
    Which of the following bug finding tools do you use with codebases containing unsafe Rust?
  7. 7.
    [If yes to #1] - Foreign functions that you call from Rust are written in:
  8. 8.
    [If yes to #1] - Your code that calls foreign Rust functions is written in:
  9. 9.
    [If yes to #1 or #2] - Select the tools and methods you use to create foreign bindings to Rust, or from Rust to other languages.
(b) The interview protocol.
  1. 1.
    What have been your motivations for learning Rust and choosing to use it?
  2. 2.
    What do you use unsafe Rust for?
  3. 3.
    You mentioned using [screening survey #5] in unsafe contexts. What do you use them for? Describe how you reason about memory safety in these situations.
  4. 4.
    Describe a bug you faced that involved unsafe Rust.
  5. 5.
    You mentioned using [screening survey #9] to create foreign bindings. Describe your experiences with these methods.
  6. 6.
    [If yes to screening survey #1 or #2] - How do you navigate the differences between Rust’s memory model and other memory models?
  7. 7.
    You mentioned using [screening survey #6]. Describe your experience using each one.
  8. 8.
    Do your development tools handle all of the problems that you face writing unsafe Rust?
(c) Interview participants’ affiliations ("Affil."), their years of experience using Rust ("Years"), and the number of coding decisions made for their interview (Decs.)
ID Affil. Years Decs.
1 I 2 56
2 I,R 10 76
3 I,O 5 93
4 I,O 8 71
5 I,O 2 61
6 A 3 71
7 I 6 57
8 I 5 49
9 O 2 97
10 O 2 59
11 I 5 86
12 I,A 6 121
13 I 10 92
14 R,O,I 7 94
15 I 1.5 58
16 I 3 20
17 A 4 64
18 I 3 82
19 I 4 13
I = Industry  R = Rust Team
A = Academia  O = Open Source
(d) Codes for the theme "Why unsafe?", which describes participants’ motivations for using unsafe code.
Code Definition Participant Quote
Increase Performance The user is choosing to use unsafe because they perceive that it would be more performant than safe code. It does not matter if they actually measured performance or not. P3 "I can use unsafe to speed up my code a bit and then wrap that in some sort of safer, larger abstraction."
Easier or More Ergonomic There is a stated or implied choice between using a safe pattern and using an unsafe pattern, but the participant chose the unsafe pattern due to the perception that it was easier to implement or validate. P4 "It was simpler to use the raw allocation and deallocation API directly instead of trying to use the existing higher level primitives."
No Other Choice When there is no safe option or existing encapsulation to accomplish something. P10 "I use unsafe Rust for mostly to create new data structures or types that are not really possible to express with safe Rust."
(e) A survey question with a direct correspondence to the codes shown in Figure 0.
Think about the situations where you have typically used unsafe. Which of the following reasons have motivated you to do so?
  • I could use a safe pattern but unsafe is faster or more space-efficient.
  • I am not aware of a safe alternative at any level of ease-of-use or performance.
  • I could use a safe pattern but unsafe is easier to implement or more ergonomic.
  • Other [short answer]
4.1. Interviews
On May 3rd, 2023, we posted a call to participate in semi-structured interviews on the Rust subreddit (/r/rust), the Rust Programming Language forums, and the #dark-arts channel of the Rust Community Discord server. We also contacted acquaintances with relevant experience. To be eligible to participate, candidates needed to have at least one year of experience using Rust and had to have “regularly written or edited Rust code within an unsafe block or function.” Eligible candidates completed a series of multiple choice and short answer questions about their use of unsafe features and development tools. We provide a sample of these questions in Figure 0. We also collected participants’ years of experience using Rust and we asked them to describe their affiliation. The first and third authors collaboratively coded each affiliation as one or more of the following categories: “Industry,” “Rust Team,” “Academia,” and “Open Source.”
We used participants’ responses to the screening survey to guide each interview. We provide a sample of our interview protocol in Figure 0. Certain questions were gated based on prior responses. If an candidate answered "No" to questions #0 and #0 of Figure 0, then they did not use Rust in multi-language applications, so we did not ask them questions #0-0 of the screening survey or question #0 of the interview protocol, since each of these questions involve foreign functions.
The first author conducted all 19 interviews remotely over Zoom between May and July of 2023. We used snowball sampling (Patton, 1990) to expand our pool of participants beyond the group that responded to our initial calls to participate. If a participant shared new information that we had not yet learned, then we prompted them to invite any acquaintances with relevant experience to complete our screening survey. We stopped recruiting new participants after we had reached “data saturation” as defined by Saunders et al. (Saunders et al., 2018); we began to hear the same or similar responses to each of our questions. We transcribed the audio recordings of the interviews using Whisper (Radford et al., 2023), and the first author reviewed every transcript to correct errors and label whether utterances came from the interviewer or the interviewee.
4.2. Qualitative Analysis
We conducted an inductive, thematic analysis of our interview transcripts following best practices by Miles et al. (Miles et al., 2020). To begin, the first three authors conducted open coding of random samples of quotes taken from each interview. To create a sample, the first author picked random character indices within each of the interview transcripts and chose a snippet of one of the interviewees’ responses near each index. If all responses at that index had been sampled before, then the first author generated a new random index and repeated this process until a new quote was found. We coded three samples of quotes, and we met after each sample to resolve disagreements and record new codes within a shared codebook.
Next, the first and second authors used a card-sorting (Braun and Clarke, 2006) exercise to identify higher-level themes among the codes that we had generated in the first phase. We created cards for each of the quotes randomly sampled in the first phase, as well as for a set of quotes that were highlighted by the first author when reviewing each transcript. We used the themes that we created from this process to organize our codebook, and we removed redundant or irrelevant codes. Then, the first and second authors used this codebook to conduct closed coding of 7 additional random samples of quotes, meeting after coding each sample to resolve disagreements and update the codebook. We used the same process as before to create samples, but instead of treating each interview transcript separately, we sampled quotes from a single document containing the concatenated text of every interview transcript. Table 0 shows one of the themes that we created during this process, as well as its constituent codes.
After this refinement stage, the first and second authors divided the set of transcripts between each other to code in full, independently, using ATLAS.ti Web (GmbH., 2024). The first author coded 11 transcripts, while the second author coded the remaining 8 transcripts. Throughout this process, both authors continued to update a shared codebook with additional codes and themes to describe new observations. After all transcripts had been coded by one of the two authors, the first author conducted an additional coding pass for every transcript and resolved each of their disagreements with the second author. The fourth column of Table 0 shows the final number of coding decisions made for each participants’ transcript. The complete codebook and all coding decisions are included in our replication package.
4.3. Survey
We constructed the community survey after we had completed coding every interview, but before the first author had finished auditing coding decisions. Each of the first three authors independently assembled a list of survey questions, which the first author assembled into a single survey in Qualtrics (Qualtrics, 2024). Excluding consent and eligibility, the survey had 85 yes/no, multiple-choice, and Likert scale questions, which were grouped into 26 sections. To avoid priming participants to select certain responses, we randomly shuffled the order of multiple choice and yes/no options and randomly reversed the order of Likert scale items. Similar to our screening survey, certain questions were gated based on prior responses. We provide a complete copy of our survey, including its branching logic, in Section 4 of our Appendix.
Questions were either descriptive or exploratory. Figure 0 shows an example of a descriptive question. Each of its response options correspond to the codes shown in Figure 0 for the theme “Why unsafe?”. Exploratory questions arose after we had finished conducting interviews, and they did not have a direct correspondence to the themes that we identified during qualitative analysis. For example, Höltervennhoff et al. (Höltervennhoff et al., 2023)’s study was published concurrent to our investigation, and three of their participants mentioned using either cargo-audit or cargo-deny to detect vulnerabilities in their dependencies. None of our interview participants had mentioned using these tools, but we included a survey question asking whether respondents had used them before.
We also relaxed our eligibility criteria for the survey to include individuals who would have been eligible for Höltervenhoff et al. (Höltervennhoff et al., 2023)’s study. Candidates were eligible to participate in their study if they had committed unsafe code to GitHub, but there was no indication if they had done so regularly. Our interview participants needed to have “regularly” engaged with unsafe Rust, but our survey respondents were eligible if they had “written, edited, read, audited, or engaged with unsafe Rust code in any way.”
Two of our interview participants piloted an initial draft of the survey, and we incorporated their feedback into the final version. We distributed our community survey on September 22nd, 2023 to the same acquaintances and communities where we distributed our screening survey. We also published a link to our survey in the September 28th edition of the “This Week in Rust” newsletter, and we advertised the survey during a presentation and poster sessions at an industry partners conference hosted by our institution. At the end of the survey, participants could provide a link to their profile on either GitHub or the Rust Programming Language Forum to have a chance to receive one of two \$250 gift cards. Only participants with public account activity prior to the start date of the survey were eligible for compensation. We closed the survey on October 9th, 2023.
5. Ethics & Threats to Validity
Ethics
All procedures in this study were approved by our institution’s IRB. Prior to each intervention, participants were read or presented a script outlining the procedures of the study and they were asked to affirm consent. Participants were reminded to avoid sharing personally identifying information in their responses, and interview participants were asked to remain in a private space. The first author reviewed the contents of the replication package and redacted all direct (e.g. name, affiliation) and indirect (e.g. position, project) personal identifiers.
External Validity
Recruitment via snowball sampling and through online communities may bias our results toward certain areas of the Rust community. We attempted to maintain a diverse survey and interview population by sampling from multiple forums. However, both Reddit and “This Week in Rust” were over-represented compared to the other communities where we distributed our survey. We cannot determine if our interview population has a similar bias, since we did not record where interview participants had first learned about our study. Additionally, the gift cards that we offered for compensation are most easily redeemable in North America, which may have dissuaded certain individuals from participating in either stage of our study. We did not collect location data from participants, so we cannot determine if our results are biased toward individuals from any particular area.
Respondents were eligible to complete our survey if they had one year of experience and had “engaged” with unsafe Rust in any way. A subset may have been eligible without having any significant practical experience with unsafe code. We mitigated this by gating questions based on whether participants had used certain features. When relevant, we compare results from the entire population with results from the subset who “regularly” wrote or edited unsafe code, which matches the eligibility criteria for our interviews. This criteria may have introduced self-selection bias among our interview participants, since individual who frequently use unsafe code may be more likely to perceive that it is necessary for their purposes.
Construct Validity
The validity of our results depends on participants’ ability to accurately self-report their behavior, and what developers perceive about their use of unsafe code may differ from how it is used in practice. Our qualitative results are unavoidably influenced by the individual biases of each investigator. We attempted to mitigate this through auditing coding decisions and meeting frequently to resolve disagreements. When we distributed our community survey, we configured Qualtrics (Qualtrics, 2024) to reject multiple responses from the same individual using IP information (which was not recorded or visible to investigators), RelevantID (RelevantID, 2024) and reCaptcha (Liu, 2018). We did not enable these features for the initial screening survey. The overwhelming majority of interview participants enabled their camera during interviews, but a few did not; they were not prompted or required to do so. We rely on the baseline assumption that every participant was honest about their identity, but we could not independently verify this in every situation.
6. Results
We found that participants used unsafe code for ergonomics, performance, and when they perceived that there was no safe alternative. They attempted to keep unsafe code minimal and encapsulated, but many participants were still uncertain about its correctness. Most participants had called foreign functions, and non-trivial differences between Rust and other languages made these calls difficult to encapsulate in certain situations. Miri was the most popular bug finding tool. However, while more than 60% of respondents had used Miri before, its performance and lack of support for key features deterred 69% of these respondents from using it again. Our participants would benefit from tools that can provide guarantees of soundness and that have robust support for multi-language applications.
We report the demographics of interview and survey participants in Section 0. Then, we describe how these developers reason about foreign functions in Section 0 (RQ1), and we discuss the tooling they use in Section 0 (RQ2). We describe developers’ motivations for using unsafe code in Section 0 (RQ3), and we report how they reason about encapsulation in Section 0 (RQ4). We summarize key results and provide suggestions for tool designers at the end of each section.
6.1. Demographics
[…]
Years of Experience
Location # SE C C++ Rust
Rust Forums 16 17.3 ± 10.1 7.5 ± 6.8 5.6 ± 6.2 4.5 ± 2.4
Rust Discord 10 8 ± 3.7 2.7 ± 2.8 3.4 ± 2.3 3.6 ± 2
Reddit (/r/rust) 83 13.9 ± 9.1 7.6 ± 7.2 6 ± 6.3 4.1 ± 2.2
"This Week in Rust" 52 15.1 ± 9.7 6.6 ± 7.7 5.7 ± 6.9 4.1 ± 2.3
Other 3 23.7 ± 15.8 17.3 ± 14.6 9 ± 5.6 3.7 ± 1.5
All 160 14.5 ± 9.5 7.2 ± 7.5 5.8 ± 6.3 4.1 ± 2.2
We invited all 42 eligible candidates to participate in interviews, and we interviewed 19 who responded to our invitation. Table 0 summarizes interview participants’ demographic information. There were 15 who had experience in industry, 6 in open source development, 3 in academia, and 2 were members of the Rust Team. They had 4.7 years of experience with Rust on average, and the most experienced participant had been using Rust for 10 years. We did not collected any additional demographic information from our interview participants.
We received 370 survey responses, of which 242 were complete, 203 met our eligibility criteria, and 160 passed each of Qualtrics’ (Qualtrics, 2024) measures for fraud detection. Hereafter, when we refer to “respondents,” we include only the 160 that passed every one of our checks. We collected several categories of demographic information from survey respondents. The majority were young, educated, identified as male, and had experience in either industry or open-source development. Nearly 66% had completed at least a bachelor’s degree, and 30% had completed a graduate degree. Most respondents (83%) were younger than 40 years old, and 56% were between the ages of 18 and 29. We followed Spiel et al.’s (Spiel et al., 2019) guidelines for ethically surveying respondents’ gender; 80% identified as “Man”, 6% identified as “Non-binary”, 3% identified as “Woman” and 9% did not disclose their gender. Table 0 summarizes respondents’ years of experience with the field of software engineering, Rust, and other systems languages. On average, respondents had more than a decade of experience in software engineering, more than 5 years of experience in both C and C++, and just over 4 years of experience using Rust.
Table 3. How frequently survey respondents “engaged with” and wrote unsafe code. Percentages on the left include “Less than once a year” and “Yearly”, while those on the right include “Weekly” and “Daily”. The number of respondents in the sample is listed in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
[…]
[Uncaptioned image]23%41%(160)
3b How frequently do you write new Rust code or edit existing Rust code within an unsafe block or function?
[Uncaptioned image]31%26%(160)
[…]
Participants had a variety of roles within the Rust community and differing levels of engagement with unsafe code. Among survey respondents, 74% had committed unsafe code to a public GitHub repository, and 66% had contributed to a published crate. Our interview population had regularly written or edited unsafe code, but 47% of our survey population did not. Since “regularly” is subjective, we asked the questions shown in Table 0 to provide a more specific time-frame and to distinguish editing unsafe code from practices that do not require code contributions, such as auditing. Only 26% (6.1) of survey respondents would write or edit unsafe code more than once a month, but 41% (6.1) would engage with it at this frequency.
6.2. RQ1 - Interoperation
The majority of interview participants and 70% of survey respondents used unsafe code to call foreign functions. The five most popular languages that respondents interoperated with were C (90%), C++ (57%), JavaScript or TypeScript (22%), assembly (21%), and Python (17%). With the exception of Assembly, the relative popularity of each language is similar to results from Fulton et al. (Fulton et al., 2021). Assembly would have been included as part of the category "Other" in their survey, which was selected by only 7% of respondents. Additionally, 72% of our respondents had converted part or all of an application written in another language to Rust, and 16% of these respondents had not called foreign functions. This supports Zeng and Crichton (Zeng and Crichton, 2018)’s notion that some developers may prefer to rewrite entire applications in Rust instead of relying on interoperation to incrementally translate them. However, none of our evidence suggests that this is a dominant preference.
Memory Management
Rust’s standard library provides several types that can be used to manage ownership of heap allocations. Each of these types can be converted to raw pointers when necessary, but similar to Rust’s safe references, they have distinct invariants that must be maintained to avoid undefined behavior. Participants frequently used these types to allocate memory for foreign function calls, converting them into raw pointers in the process. In particular, 54% of survey respondents who had called foreign functions also used Rust’s memory container types for this purpose. Rust’s thread-safe Arc—an atomic, reference-counted allocation—was used by 28% of these participants, while only 8% had used the non-atomic Rc. Rust’s Box provides unique ownership over a heap allocation, and it was used by 95% of these participants to allocate memory for foreign calls. One interview participant used Box extensively when porting an online multiplayer text-adventure game from C to Rust. The game’s logic was written in C, but all heap memory was allocated in Rust using Box. When a Box was unwrapped and exposed as a raw pointer to C, its address was added to a table. The table was checked and updated to ensure that each allocation was only freed once.
Several interview participants mentioned having issues with Box, noting that when a Box is moved, any raw pointers to its contents are invalidated. However, participants with a background in C++ development had not expected Box to behave this way. They were accustomed to using unique_ptr from the C++ standard library, which has a similar role but different semantics:
You give the raw pointer off to the interface, and then you relocate the unique_ptr, but you still rely on that pointer having valid access tags. (P0)
Survey participants did not encounter this issue as frequently. Of the 51% of respondents had used Box in any unsafe context, only 17% had encountered issues when a pointer because invalid after a Box was moved. However, we did not determine if respondents had actually unwrapped a Box or otherwise retained a raw pointer to its contents, so this 17% could still represent a significant portion of the respondents who would have been able to encounter this problem.
One participant observed several aliasing patterns in foreign libraries that were at odds with Rust’s restrictions. They struggled to encapsulate foreign functions that would read from a “source” pointer and write to a “destination” pointer. It was difficult to use safe references as parameters for the Rust encapsulations of these functions when the source and destination pointers needed to be aliases. The borrow checker would prevent the mutable destination reference from aliasing with the immutable source reference. Interview participants also noted that C allows unrestricted integer-to-pointer conversion, but Rust does not have an agreed-upon semantics for this behavior (Jung, 2024). Another participant observed that C libraries would often use on “a rat’s nest of pointers” (P0) to encode self-referential structures, which were difficult to represent in Rust encapsulations. Each of the aliasing violations from the most popular libraries tested by McCormack et al. (McCormack et al., 2024) were caused by incorrect encapsulations of self-referential structures.
To manage these differences, participants reviewed the documentation and implementation of foreign libraries to determine their safety requirements. However, one participant found that documentation was frequently missing, so they had to manually reason about these properties.
It’s really difficult to figure out a lot of the properties of a given codebase…are you actually mutating something?…if I give you some input data, are you going to retain a pointer to that and keep operating on it later?…it’s a very human issue of lack of documentation and lack of correctness within a codebase. (P0)
The reverse of this situation occurred when participants exposed Rust libraries to languages with fewer safety guarantees. Properties that were implicitly enforced by Rust’s type system became documentation in other languages.
In Rust, you can trivially say, I’m going to return to you a thing, which just borrows my memory, and then you just can’t access me while you’re using that …this is not very easily mappable into a C API, or rather you could do it, but then you have to read all this documentation… (P0)
Some foreign APIs were structured in a way that made memory management easier. One participant found that “there [were] no complex lifetimes involved” (P0) in their encapsulation of an embedded library. Others observed that the properties they typically associated with a “well-crafted C and C++ API” (P0) made it easier to create a Rust encapsulation. For example, one participant observed that C APIs tend to expose pairs of initialization and cleanup functions for each type, which could be linked into the behavior of the Drop trait for its Rust encapsulation. This trait provides a cleanup function that is called to free the resources associated with an object when it leaves scope.
Concurrency
Interview participants typically used more sophisticated multithreading patterns in the Rust components of multi-language applications. One participant used granular locking within the Rust codebase, while the C codebase had a singular global lock. For another participant, the C++ component of their application was structured along a single main thread, while the Rust component used multithreading freely—“we were throwing threads around all the time” (P0). It was easy for this participant to accidentally call into C++ from the wrong Rust thread, but they saw this as less challenging to manage than it would have been to implement the entire application in C++:
It was a validation of a belief…that this would have been impossible to do in pure C++…we had to hit this [problem] for the occasional times we had to call out to C++ from Rust…but if we were only writing C++ code… we would have been hitting this constantly. (P0)
However, Rust’s approach to concurrency also made it difficult for one participant to encapsulate a foreign library. In Rust, a type can automatically derive the traits Send and Sync if all of its components implement these traits. However this participant observed that in C and C++, a single type can have methods and components with varying thread-safety properties: “Rust enforces thread safety based on types… C tends to do it based on functions” (P0). They found that it was more difficult to encapsulate foreign types that were only partially thread-safe.
Table 4. How often survey respondents would avoid passing abstract data types by value to foreign functions and converting foreign raw pointers into safe references. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in the sample is listed in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
[…]
[Uncaptioned image]49%24%(112)
4b How often do you intentionally avoid converting raw pointers to memory allocated by FFI calls into safe references, such as &T or &mut T?
[Uncaptioned image]23%43%(112)
[…]
Information-Hiding
Participants attempted to keep interoperation with foreign code as straightforward as possible. When declaring bindings, one participant preferred using primitive types and pointers instead of abstract data types like Option and NonNull, even though Rust allows these types to be implicitly cast into raw pointers (Community, 2024). They felt that these casts hid useful contextual information, so they preferred to use raw pointers explicitly. A few participants indicated that they sometimes avoided passing structs by value across foreign boundaries. One of these participants observed that certain Rust programs would behave incorrectly when structs were moved across a foreign boundary instead of copied, or if they had a destructor on the C++ side of the boundary. They typically avoided passing structs by value in either situation, and they recommended that “…if you’re doing FFI, the only things you should pass by value are Copy types” (P0). Other interview participants also indicated that they felt comfortable passing structs by value as long as they had copy semantics and equivalent types on each side of the boundary.
In addition to limiting the use of certain types at boundaries, interview participants also attempted to minimize the interactions between each side of the foreign function boundary. They thought that a “very chatty API” (P0) would be difficult to encapsulate correctly. One participant described this complexity in terms of the heap object graph that a foreign library exposes to Rust. When the structure of the foreign heap was exposed to Rust in detail, safe encapsulations became difficult to use:
It’s hard to keep those objects around for very long unless you want all of your structs to end up with a bunch of lifetimes… you don’t want that… it’ll scare away new programmers for sure. (P0)
They typically avoided casting foreign pointers into Rust’s safe reference types since “applying lifetime semantics to foreign code” (P0) required difficult, manual reasoning. Another participant used opaque types to hide the underlying layout of foreign values. This design pattern is useful in situations where Rust needs to be able to reference a foreign value, but the value does not need to be accessed on the Rust side of the boundary. Instead of declaring the structure of its type on each side of the boundary, a value can be given a zero-size, “opaque” placeholder type in Rust.
We investigated these information-hiding practices using the survey questions shown in Table 0, where respondents indicated if they “intentionally avoid” passing structs by value or casting foreign pointers into Rust’s references. We added “Unsure” as a response option for these questions since they indirectly rely on participants having a notion of why they would “intentionally” avoid these behaviors. Neither practice was consistently followed, but more respondents would avoid passing abstract data types by value most or all of the time (49%, 6.2), than would avoid casting foreign pointers into references (43%, 6.2). Key Findings - RQ1 (Interoperation).
How do Rust developers reason about memory safety across foreign function boundaries? The majority of participants called foreign functions, which were typically written in C and C++. Certain foreign libraries used aliasing and concurrency patterns that conflicted with the expectations of Rust’s type system, which made them difficult to encapsulate. To make interoperation easier, participants minimized their interaction with foreign libraries, but they frequently used Rust’s memory containers as allocators, and few actively avoided casting foreign pointers into safe references. Developers identified safety properties and design patterns of foreign APIs that match the characteristics of libraries where McCormack et al. (McCormack et al., 2024) detected aliasing violations, indicating that foreign function calls may be a significant unchecked source of undefined behavior in the Rust ecosystem. We expect that developers who write encapsulations for foreign APIs would benefit from analysis tools that can infer their safety properties. This information could be leveraged to automatically generate sound, safe encapsulations.
6.3. RQ2 - Tooling
Table 5. How often survey respondents audited their dependencies, used a debugger, wrote tests, and executed tests in Miri (if they had used Miri at least once). Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[l,m]$Q[c,m]$
ID Question Sample Distribution
[…]
[Uncaptioned image]61%30%(160)
Used Miri
[Uncaptioned image]72%19%(98)
5b How often do you run test cases in Miri? Used Miri
[Uncaptioned image]38%56%(98)
[…]
[Uncaptioned image]6%88%(160)
Used Auditing Tools
[Uncaptioned image]10%80%(81)
[…]
Table 6. Whether or not survey respondents who had used foreign functions would trust the correctness of foreign function bindings written by hand and generated using a tool. Percentages on the left include “Definitely yes” and “Probably yes”, while those on the right include “Probably not” and “Definitely not”. The number of respondents the each sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
[…]
[Uncaptioned image]45%7%(112)
6b Do you trust FFI bindings that are generated by a tool?
[Uncaptioned image]79%2%(112)
[…]
Participants used a wide variety of development tools to assist with validating their design choices. We focused primarily on dynamic tools for bug finding and static tools for generating bindings to foreign functions. However, certain participants also described their experiences using debuggers, verifiers, and auditing tools.
Dynamic Tools
Most participants used dynamic bug finding tools. Interview participants often mentioned Miri, AddressSanitizer, and Valgrind. We prompted survey respondents to indicate which dynamic tools they had used from a list of options mentioned by our interview participants. Survey respondents most frequently used Miri (61%), Valgrind (44%), AddressSanitizer (26%), and cargo-fuzz (26%). Each of the remaining tools listed in our survey was used by less than 15% of respondents. Three interview participants and 31% of survey respondents used some form of fuzzing tools. One interview participant leveraged industry sponsorship to fuzz a JIT compiler for hours-on-end. Another participant implemented a BTree with optimizations for high performance computing, and they used libfuzzer to perform differential testing (McKeeman, 1998) against the implementation provided by Rust’s standard library. Of the 26 developers interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023) only seven had used Miri, seven used fuzzing tools, and three used Valgrind. Our results indicate that this is not representative of a broader trend in the Rust community.
The majority of survey respondents also wrote tests for applications that used unsafe code. As shown in Table 0, 61% (6.3) of all participants wrote tests most of the time or always. This was equally true for the subset of participants who had used Miri; 72% would write test cases at least most of the time. However, 56% (6.3) of survey respondents who had used Miri only sometimes used it to run their test cases, if at all. Höltervennhoff et al. (Höltervennhoff et al., 2023) also had similar findings. The majority of their participants wrote unit tests for unsafe code, but a subset claimed that they would infrequently run these tests because foreign function calls or low-level hardware interaction made testing “impossible.” While notable, we cannot make any causal claims between Miri’s limitations and this lack of testing in our survey population. Additionally, we did not ask if survey respondents regularly ran their test cases. Interview participants frequently mentioned that Miri’s slow performance and lack of support for foreign function calls prevented them from using it to find bugs in certain categories of applications. Survey respondents who had used Miri encountered similar issues; 62% were also deterred from using it due to its lack of support for foreign functions (43%), slow performance (26%), or lack of support for inline assembly (19%). Developers interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023) also struggled with Miri’s lack of support for key features, but none had mentioned its performance.
Notably, one of our interview participant had neglected to use Miri since their application relied heavily on foreign function calls: “we’ve assumed that we are a JIT and Miri can’t run JIT code, so therefore it’s useless” (P0). However, after a vulnerability was discovered in their codebase, they decided it that it would be worthwhile to run Miri on the subset of their test cases that did not require foreign function calls. They discovered that Miri would have been able to detect the vulnerability earlier, since it was caused by a violation of Rust’s aliasing model. Another interview participant managed to circumvent Miri’s limitations by creating mocks of foreign functions, but this solution did not scale to programs with “50,000 lines of code” (P0). This participant and a few others were aware of the Krabcake project (Klock and Garza, 2023), which aims to extend Valgrind with Miri’s borrow tracking features. Each of these participants was already using Valgrind with Rust codebases, and they felt that Krabcake would be an ideal solution for overcoming Miri’s limitations.
Foreign Function Bindings
Interview participants encountered several issues with tools that generate bindings to foreign functions. Participants reported that bindgen added significant runtime overhead during build times and was not as expressive as they desired. One participant observed that bindgendid some weird stuff for anonymous unions and structs” (P0) but that this was easy to fix by modifying the tool’s configuration. However, another participant found it necessary to switch from using bindgen to writing bindings by hand, because it was not capable of expanding the macros that they had used to define core data structures. Others felt that the bindings they wrote by hand were more idiomatic than the output of a tool. The majority (58%) of survey respondents who created bindings used multiple methods, while 28% only generated bindings and 14% only wrote them manually.
Interview participants who wrote bindings by hand viewed this process as error-prone, noting that bindings become a “fiddly mess” (P0) when you need to ensure that interfaces remain consistent on each side of the foreign boundary. These errors clearly do occur in practice, as 35% of survey respondents who had called foreign function had also encountered incorrect bindings. McCormack et al. (McCormack et al., 2024) also found several instances of incorrect bindings in Rust libraries. Some were intentional—developers neglected to add return types for values they did not use—but others introduced aliasing violations or uninitialized reads. Survey respondents who had called foreign functions were somewhat more likely to trust generated bindings. As shown in Table 0, only 45% (6.3) of respondents would at least “probably” trust handwritten bindings, compared to 79% (6.3) who would trust generated bindings. However, developers were still more likely to trust handwritten bindings than to distrust them.
Debugging
Few interview participants reported using a debugger with Rust code. Only 10% of survey respondents used a debugger daily, 29% used one monthly, and 44% used one yearly at most, if at all. One interview participant had used both MSVC’s debugger and LLDB, and while both worked, neither matched the quality of the other parts of the Rust toolchain, “where you get all the bells and whistles that you want, and things just work nicely together” (P0). Mozilla’s RR debugger (O’Callahan et al., 2017) was useful for one participant to resolve bugs found from fuzzing, but it was generally more common for participants to report using “printf-style” (P0) debugging.
Formal Methods
Few interview participants and only 10 survey respondents used tools that apply formal methods for static and dynamic verification. Interview participants mentioned using Kani, Prusti, and Crucible. Kani was useful for one participant, who needed to prove that a particular routine would never execute more than twice for any input. Another participant who had “tried a bunch of the tools” in this category found that Prusti was the most complete, but they felt that it would be more effective if these tools were “inherently parts of Rust” so that they could “hoist a lot of currently unsafe code into safe code” (P0). Of the ten survey respondents who had used “formal methods tools,” five had used Kani, three had used Prusti, three had used Creusot, one had used Flux, and one had used Crucible. We provided a short-response option for this question to include any tools that were missing from our list.
Auditing
Only a few interview participants indicated that they had ever audited their dependencies. Höltervennhoff et al. (Höltervennhoff et al., 2023) had similar findings; half of the developers that they interviewed had deliberately included dependencies with unsafe code, but the majority did not audit their dependencies. One of our participants would examine their dependencies if they were curious about how they functioned, and they actively avoided adding dependencies that used outdated versions of Rust or deprecated features. Another participant was in the process of having a third-party institution examine their codebase to determine if they could reduce their amount of unsafe code.
Survey respondents rarely audited their dependencies’ use of unsafe code. More than 80% (6.3) of respondents only sometimes audited their dependencies’ use of unsafe code, if at all. It is possible that a subset of the participants who answered “Never” to this question did not have any dependencies to audit. However, 51% of our respondents had used automated auditing tools, compared to four interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023). Of this subset, 63% had used cargo-audit, 42% had used cargo-update, 31% had used cargo-deny, 22% had used cargo-geiger, and only 7% had used cargo-vet.
Key Findings - RQ2 (Tooling).
What tools do Rust developers use when contributing to applications that include unsafe code, and how could tooling be improved? Miri appears to be the “de facto” (Gäher et al., 2024) bug finding tool for the Rust ecosystem, since it was used by 61% of all respondents. However, 62% of the respondents who had used Miri before were deterred from using it again due to its performance and lack of support for key features—predominately, foreign function calls. Tools for generating foreign function bindings were often slow and required additional configuration, so a subset of developers still preferred writing bindings by hand. Participants infrequently used debuggers, and only 10 survey respondents had used formal methods tools. Survey respondents infrequently audited their dependencies, but 51% had used one or more auditing tools. Improvements to Miri may be the most effective way to provide stronger safety guarantees for the majority of our audience, since it was the most widely used tool in our survey.
6.4. RQ3 - Motivations
Table 7. Survey respondents’ motivations for using unsafe code, as measured by the question shown in Figure 0. All but one participant chose to answer this question. Each cell shows the percentage of participants who identified with both of the motivations for that row and column. The cell at the bottom right hand corner shows the percentage who selected all three motivations.
Necessity Performance Ergonomics
Necessity 77% 92% 86%
Performance 47% 53%
Ergonomics 18%
Participants used unsafe code because they perceived that it was more performant or ergonomic than safe alternatives, or that there were no safe alternatives at all. Here, we refer to these motivations as “Performance,” “Ergonomics,” and “Necessity,” which correspond to the codes “Increase Performance”, “Easier or More Ergonomic,” and “No Other Choice” from Figure 0, respectively. We used the question shown in Figure 0 to measure the distribution of each of these motivations in our survey population. All but one participant chose to answer this question. Results are shown in Table 0. Necessity was the most common motivation, followed by performance and ergonomics.
Table 8. How frequently participants measured the impact of their changes when using unsafe code to improve performance. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in the sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
8a When you choose to use unsafe because it performs faster or is more space efficient, how often do you measure the difference? Used unsafe for performance
[Uncaptioned image]46%36%(76)
[…]
Performance
Some interview participants and 47% of survey respondents indicated that they would use unsafe code when they perceived that it was faster or more space-efficient than an equivalent safe API. Participants usually identified a safe alternative in these situations, but they felt that it would be too expensive to use: “so you could do it, right? But that would be some performance overhead” (P0). Most interview participants who attempted to use unsafe to improve performance had implemented small-scale optimizations in existing applications. These typically involved eliminating runtime checks, which were seen as unnecessary due to what they perceived were local or global invariants. For example, one participant used an unsafe API to consume a pointer to a string without checking if it contained valid UTF-8 characters. This pointer was provided by a foreign call to Python, and the participant thought that it would only ever provide text in a valid format. Another participant chose to store a heap allocation as a raw pointer in a static, mutable variable instead of using one of Rust’s safe static encapsulations, since they felt confident that the allocation would remain valid for the duration of the program.
However, a few interview participants used unsafe code to create entire components that were purpose-built to achieve performance. In each situation, performance was seen as a functional requirement for the application domain: “I write a serialization framework…and obviously a goal for it is to be extremely performant” (P0). A participant who took this approach found that it significantly increased the amount of unsafe code in their application, but they felt that this was a reasonable compromise to achieve greater performance. Survey respondents who reported using unsafe code to improve performance did not have a strong tendency toward either large-scale or small-scale use. We found that 26% only pursued “small-scale-optimizations in existing applications,” 20% only built new, “large-scale components,” and 40% used unsafe code for performance in either situation.
Participants did not consistently measure the performance impact of unsafe code. In some situations, they relied on their intuition to determine which design decisions would be best: “I’ll totally admit that…I think this is going to be a hot thing, so I’m going to kind of prematurely optimize” (P0). Survey respondents who used unsafe to increase performance were also inconsistent about measuring its impact. As shown in Table 0, 46% of respondents reported measuring performance at least most of the time, while 36% only sometimes measured performance, if at all. In contrast, Höltervennhoff et al. (Höltervennhoff et al., 2023) found that 6 participants would only use unsafe code to improve performance when the change was “noticeable.” We did not make a distinction between profiling and subjective measures of performance in our survey. However, one of our interview participants performed sophisticated profiling on three versions of a component, each of which had varying amounts of unsafe. The most unsafe-heavy version performed the best, but they opted for one with a moderate amount of unsafe code, balancing safety and performance concerns.
Ergonomics
Participants also chose to use unsafe code when they felt it would be easier or more ergonomic than using a safe API. This was the least common motivation cited by interview participants, and only 18% of survey respondents identified with this motivation. Höltervennhoff et al. (Höltervennhoff et al., 2023) indicate that some of their participants would use unsafe code if it “saved them effort”, but it was unclear if this motivation was typical. Our participants usually stated or implied a choice between using safe or unsafe design patterns. For example, one participant chose to use only Rust’s unsafe allocator API rather than a mix of unsafe and safe allocation operations. They found that this approach was easier to reason about, since “if there’s too much abstraction, you lose the information that you need to make it actually sound” (P0). Several participants used the unsafe function transmute to perform unrestricted type conversion. Since types have differing invariants, transmute can only be used in unsafe contexts. One participant used transmutation to simplify the implementation of chess engine, which used three enumerations to represent the position of a chess piece. They found that it was easier to use integer conversion, bitwise operations, and transmutation to convert between these enumerations instead of safely matching on their values. Another participant used transmutation when working with a Rust encapsulation of a C font library. This library exposed C heap allocations as Rust s with short lifetimes—even though these objects would remain valid on the heap for the duration of the program. This participant found it was easier to transmute the Rust objects to have the 'static lifetime, which is indefinite, instead of adapting to the restrictions of the encapsulation.
I just…transmuted that one to 'static… another use case for unsafe is working around a bad API when you know that you can use it correctly. (P0)
Table 9. How frequently participants who felt that they had “no clear alternative” to using unsafe code were certain it would be impossible to avoid it. The percentage on the left includes “Always” and “Most of the time”, while the one the right includes “Sometimes” and “Never”. The number of respondents in the sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
9a When you feel that you have no clear alternative other than using unsafe, how often are you certain that it would be completely impossible to accomplish this task using a safe design pattern? Had no alternative to unsafe code
[Uncaptioned image]65%17%(123)
[…]
Necessity
In most situations, participants perceived that they had “no clear alternative” to using unsafe code. The majority of interview participants and 77% of survey respondents identified with this motivation. Höltervennhoff et al. (Höltervennhoff et al., 2023) observed this to a similar extent; 18 out of their 26 participants claimed to use unsafe code by necessity. Our interview participants often reported being constrained by a combination of Rust’s type system and the properties of their applications. For example, one participant found that it was necessary to implement Send and Sync in a single-threaded application to maintain compatibility with an existing API. The perceived that the implementation was sound by construction. However, each of these traits are declared as unsafe, so they can only be implemented in an unsafe context.
Participants who contributed to just-in-time (JIT) compilers and operating systems found that access memory through raw pointers was unavoidable. One participant found that this was necessary to implement an unwinding table: “I have to jump to that address and I really can’t do anything to verify it…” (P0). In other situations, developers felt that there could be a safe alternative, but they were unaware of how to implement it: “there’s probably a better way that I just don’t know about yet or didn’t know about at the time and I haven’t thought about it” (P0). However, as shown in Table 0, 65% of survey respondents were certain at least most of the time, if not always, that it would be impossible to implement an equivalent safe pattern.
Co-occurrence
These motivations were not mutually exclusive. One participant implemented a zero-copy deserialization pattern using unsafe code, and they could relate to each of the reasons we identified. Performance was a domain-specific constraint of the internationalization library that they contributing to, which motivated them to use zero-copy deserialization. We reviewed the library’s documentation, and it indicated that contributors had considered using an existing crate, but its API was not ergonomic for their use case. The participant perceived that unsafe code was inherently necessary to be able to implement their own version.
Key Findings - RQ3 (Motivations).
What are Rust developers’ motivations for using unsafe code? Participants were most often motivated to use unsafe code because they felt that there was no safe alternative. In other situations, participants used unsafe code because it performed better or was more ergonomic. For performance, developers pursued both large-scale abstractions and small-scale optimizations, but they did not consistently measure the impact of their changes. Each of these motivations can influence a design decision simultaneously. We expect that interventions that consider the intersection of multiple motivations will be more effective. For example, profiling tools could leverage information on the distribution of unsafe code to assist developers in making a compromise between safety and performance.
6.5. RQ4 - Encapsulation
Table 10. Survey questions relevant to RQ4. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
10a When you use an unsafe API, how often do you look for documentation to ensure that you meet all of its requirements for safety and correctness? Used unsafe APIs
[Uncaptioned image]85%8%(152)
10b When you expose a safe API for unsafe code, how often do you include runtime checks to ensure that its requirements for correctness and safety are met? Exposed a safe API for unsafe code
[Uncaptioned image]58%29%(141)
[…]
[Uncaptioned image]90%3%(68)
10d When you expose an unsafe API to other users, and it is their responsibility to ensure that certain requirements are met, how often do you include runtime checks for these requirements?
[Uncaptioned image]18%71%(68)
[…]
[Uncaptioned image]6%88%(160)
Regularly wrote unsafe code
[Uncaptioned image]7%85%(85)
[…]
[Uncaptioned image]82%10%(160)
Used unsafe APIs
[Uncaptioned image]83%10%(152)
[…]
Invariants
Participants reasoned about the soundness and correctness of their unsafe code in terms of invariants. When developing operating systems and embedded applications, these invariants were provided by hardware specifications. For example, one participant who developed an embedded application would write a “magic value” (P0) to RAM to indicate which operation had triggered a reboot. The microcontroller that they used did not erase its memory on reboot, so they could rely on the value being initialized. Some participants perceived hardware-level correctness properties as “orthogonal to Rust” (P0), since they were related to the properties of the language. However, other participants who developed embedded applications relied on third-party libraries to encode hardware requirements into the type system, so that their program would fail to compile if a device was misconfigured.
Developers leveraged Rust’s type system to uphold invariants that they perceived as necessary for soundness. In operating systems, these typically began as operation-specific refinement properties on primitive types, but they quickly built up into system-level invariants; “it gets to a high-level really quickly” (P0). Rust’s aliasing restrictions were a natural fit:
We only allow… someone who has access to one of these mapped memory regions to borrow it… That’s a great example of taking something that’s unsafe inherently and then kind of wrapping it up in a safe abstraction, using the power of Rust to do the borrow checking for us. (P0)
Another participant experimented with using ghost permissions to reason about pointer arithmetic. This method was popularized for Rust by Yanovski et al.’s GhostCell type (Yanovski et al., 2021). Other participants relied on ad-hoc reasoning to ensure that their programs were correct and free of undefined behavior. Some appealed to their prior experience: “[in] C++, you do this all the time” (P0). One participant was accustomed to reasoning about aliasing rules in compiler development, so they felt confident that they were adhering to Rust’s restrictions in unsafe contexts. Other participants validated their decisions by auditing their code, but most had some level of implicit trust; “that’s sort of another… I got to trust it type thing” (P0). This aligns with Höltervennhoff et al. (Höltervennhoff et al., 2023)—a majority of their participants reasoned about unsafe code in terms of contracts or invariants, but some reported using it carelessly, and most felt that it was difficult to write correctly due to its context-sensitivity.
Isolation
The majority of interview participants made a conscious effort to minimize and isolate unsafe code, which made it easier to document and reason about.
That’s generally the most important thing, just being self-contained, being well-isolated, being well-encapsulated… (P0)
Most survey respondents (61%) predicted that it would be at least somewhat easy for another developer at their skill level to understand a random unsafe block or function from their code. Interview participants who isolated their unsafe code had more confidence that its requirements were met, especially if they went beyond what Rust’s type system could verify. However, participants reported that unsafe code was prevalent in JIT compilers and operating systems. Domain-specific, non-local reasoning was necessary to understand any arbitrary unsafe code snippet: “you have to know the innards of the JIT-compiled function…so it’s just not encapsulated” (P0). Though unsafe code may have been isolated, it is unclear if participants used it minimally. As shown in Table 0, more than 80% (6.5) of survey respondents only sometimes refactored their code to remove unsafe features, if at all. This distribution was similar for the subset of respondents who had regularly written unsafe code.
Safe Interfaces
Most interview participants and 88% of survey respondents reported that they would attempt to encapsulate unsafe code beneath safe APIs. This was also common for Höltervennhoff et al. (Höltervennhoff et al., 2023), who found that 22 of their 26 participants mentioned safe interfacing. Our interview participants would create encapsulations where they perceived that the requirements for their unsafe code were satisfied by either Rust’s type system or what they reasoned to be local invariants. However 43% of our survey respondents reported exposing unsafe APIs, while only 6 of Höltervenhoff et al. (Höltervennhoff et al., 2023)’s 26 participants did so.
Interview participants cited the same motivations for exposing unsafe APIs as they did for using unsafe code in any capacity. One participant exposed unsafe versions of safe API endpoints which had the same behavior, but they did not include runtime checks. Another interview participant exposed unsafe APIs so that users could circumvent their safe encapsulations, in case they became difficult to use. Each of these practices aligns with another participant’s perception of a cultural value among Rust developers to make API surfaces as detailed as possible: “Rust has a tendency to… make the API as complicated as it needs to be to fully represent what is actually happening behind the scenes” (P0). Other participants indicated that they would only expose an unsafe API when it would be impossible to encapsulate without placing the burden of correctness on the user. Our survey respondents also identified with these motivations; 76% exposed unsafe APIs because it would be impossible to encapsulate them without preconditions, while 51% exposed unsafe APIs for performance and 13% did so for ergonomics. Höltervennhoff et al. (Höltervennhoff et al., 2023) also found that participants would expose unsafe APIs by necessity or to improve performance, but it was unclear if ergonomics was also a motivation.
Regardless of motivation, 61% of respondents claimed that they would only expose unsafe APIs when they had requirements beyond what Rust’s type system could verify, and they typically documented these requirements. Results for Question 6.5 indicate that 90% of respondents who exposed unsafe APIs would document their safety requirements most of the time, if not always. However, respondents did not typically add any additional safeguards to prevent unsafe APIs from being used incorrectly, since 71% (6.5) only sometimes inserted run-time checks within these APIs, if at all. This was more common when exposing safe APIs; 58% (6.5) inserted checks at least most of the time when exposing a safe API for unsafe code. When using unsafe APIs, 85% (6.5) of participants would usually look for documentation to inform their design decisions.
Interview participants perceived that the Rust community has a strong preference for safety, which affects tool design and development practices. Participants avoided exposing unsafe APIs in libraries partly because of the perception that other Rust developers would otherwise avoid using their crate: “everyone wants a safe interface… no one really wants to go use the unsafe heavy ones” (P0). One library author viewed safe encapsulation as an inevitable requirement. Even though exposing an unsafe trait places the burden of responsibility on users, they would still blame the library.
So in theory, the burden of proof is on them, but also if they implement something improperly, then they’re going to create a ticket because they’ll be like, oh, “I use your library and [this behavior is] a fault.” (P0)
An overwhelming majority of survey respondents also preferred safe APIs. More than 80% (6.5) of respondents claimed that they would choose a safe API over an unsafe alternative most of the time, if not always—regardless of whether or not they had ever used an unsafe API before.
Table 11. Survey questions relevant to RQ4. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
11a When you expose a safe API for unsafe code, how often are all of its requirements for correctness and safety satisfied by the properties of Rust’s type system? Exposed a safe API for unsafe code
[Uncaptioned image]62%18%(141)
11b How often is the Rust community’s guidance and documentation adequate for you to know how to use unsafe correctly? All
[Uncaptioned image]70%16%(160)
[…]
Uncertainty
Many interview participants who created safe APIs for unsafe code were uncertain if their encapsulations were sound in all possible situations.
No library author knows when they’re going to trigger [undefined behavior]… we do our best to give you a sound interface, but god knows if there’s a hole in it… (P0)
Höltervennhoff et al. (Höltervennhoff et al., 2023) had similar findings; 16 of their 26 participants felt some degree of uncertainty about whether their unsafe code was correct. Our interview participants were most often uncertain when unsafe was pervasive or when they called foreign functions. When describing their uncertainty about encapsulation, our participants compared the “theory” of safety to the “practice” of how an API would be used. When these concepts didn’t align, errors occurred. One participant observed this mismatch with a Rust crate that encapsulated a foreign library. The library used a memory-mapped file, and the API made it possible to unmap the file while the user retained a reference to it, leading to a use-after-free error. Another participant struggled with discrepancies between the interface and the implementation of Rust’s Layout API, which describes the size and alignment of an allocation. The design of Layout was changed to “tweak the safety requirements” (P0), which regretfully caused previously correct code to exhibit undefined behavior.
Interview participants also had a collective sense of uncertainty about Rust’s semantics. They perceived that Rust lacked a formal specification and that the guidelines for unsafe code were incomplete.
I think the biggest issue with Rust’s unsafe isn’t necessarily the tools, but just the spec—or, rather, the absence of the spec. If you write unsafe code today, you write it against the void. (P0)
A few interview participants encountered situations where the precise definition of undefined behavior was either unclear or overly restrictive. For example, one participant felt that it would be necessary to create a mutable reference to uninitialized memory. This is considered undefined behavior by Rust’s current semantics (Durham, 2022), but the participant was uncertain about how this pattern would be problematic in practice. Another participant wished to be able to take a reference to the first field of a product type and use it to create valid references to adjacent fields using integer offsets. This was considered undefined behavior under Rust’s Stacked Borrows (Jung et al., 2019) aliasing model, but it is accepted under the newer Tree Borrows model (Villani et al., 2023). Additionally, multiple participants were uncertain about how to handle the case when a program panics while a value is being dropped.
Survey respondents were somewhat more certain about their decisions. As shown in Table 0, more than 60% (6.5) of respondents who had exposed a safe API for unsafe code were certain at least most of the time, if not always, that its safety properties were satisfied by Rust’s type system. Additionally, 70% (6.5) of respondents found that the Rust community’s documentation and guidance was adequate for them to know how to use unsafe code correctly at least most of the time. However, only 6% felt that it was always adequate, suggesting that improved documentation would be helpful for the overwhelming majority of respondents. Similar to the questions in Tables 6.3 and  6.4, we added “Unsure” as a response option for the questions shown in Table 0 since they rely on developers’ subjective notions of certainty and adequacy.
Key Findings - RQ4 (Encapsulation).
How do Rust developers reason about encapsulating unsafe code? Interview participants derived requirements for unsafe code from external specifications and the inherent guarantees of the borrow checker. However, they also relied on ad-hoc reasoning and tacit knowledge. Many were uncertain if their encapsulations were sound in all situations. Participants attempted to minimize and isolate unsafe code, and they avoided using unsafe APIs in favor of safer alternatives. However, this was not always possible in certain domain-specific contexts, such as JIT compilers and operating systems, where unsafe code was used pervasively.
Survey respondents who exposed unsafe APIs documented their safety properties, and they relied on documentation to use unsafe APIs correctly. We expect that this documentation could be leveraged by tool designers to create formal specifications for these APIs. When participants chose to expose unsafe APIs, we found that they were motivated by performance, ergonomics, and necessity to the same extent as for any arbitrary use of unsafe code. Respondents rarely refactored unsafe code, even if they used it regularly. Static analysis tools could encourage refactoring by suggesting safe, idiomatic replacements, but it is also possible the the best practices for writing unsafe code make frequently refactoring unnecessary.
7. Discussion
The majority of our participants attempted to follow several best practices when using unsafe code, such as testing (6.3), writing documentation (6.5), and using safe interfaces whenever possible (6.5). As shown in Table 0, more than 70% of respondents were motivated to use unsafe code when they perceived that there was no safe alternative. However, only 14% (6.4) of these respondents were always certain that it would be impossible to implement a safe alternative. Likewise, 62% (6.5) of respondents who exposed safe APIs for unsafe code were certain “most of the time” that their interfaces were sound, but only 23% were always certain. We cannot rely on developers’ perceptions to accurately reflect the soundness of their code, but this uncertainty suggests that there is still a gap between Rust’s promise of static safety and the reality of how developers use unsafe features.
Multiple factors contribute to this uncertainty. Foreign function calls presented unique challenges due to the differences between Rust and other languages. Rust’s aliasing model is still evolving, so certain features do not yet have an agreed-upon semantics in every possible situation (Jung, 2020). Developers who implement design patterns that sit at the cutting-edge of Rust’s evolving aliasing model will inevitably be uncertain if their implementations are sound. However, all developers who write unsafe code will need a reliable way to determine if their implementations follow the rules of whichever aliasing model becomes canonical. Miri is not entirely adequate for this task, since 61% of the respondents who had used it were deterred by its slow performance and lack of support for key features. However, even a fast, feature-complete version of Miri will not be enough to ensure that developers can always be certain that their implementations are sound, since dynamic analysis is inherently incomplete.
To address the challenges faced by our participants, the next generation of Rust-specific development tools should be capable of providing guarantees of soundness and should support multi-language applications. Static analysis will be particularly helpful for encapsulating foreign functions and guiding developers toward refactoring unsafe code. However, any new approaches to tooling should consider the role of performance and ergonomics as motivations for using unsafe features.
Multi-language Support
Participants who called foreign functions would pass pointers to heap objects in both directions between Rust and other languages. Many used Rust’s memory container types as allocators, and few actively avoided casting foreign pointers into Rust’s safe references. Each of these actions can lead to undefined or undesirable behavior, since foreign code can break Rust’s aliasing rules. However, no production-ready tools are capable of detecting aliasing violations in this context. McCormack et al. (McCormack et al., 2024) were able to extend Miri with support for calling foreign functions, but the limitations of interpretation will prevent their approach from being effective beyond the context of their investigation.
A production-ready tool would need to be capable of instrumenting either native binaries or a shared intermediate format. One promising solution is the Krabcake (Klock and Garza, 2023) project, which aims to extend Valgrind (Nethercote and Seward, 2007) with a plugin for detecting aliasing violations. Valgrind provides a unique advantage compared to other dynamic analysis platforms, since it can instrument native libraries without recompilation. This solution would also be appealing for a significant portion of our respondents, since 44% had used Valgrind. However, Valgrind incurs significant runtime overhead, so it may also be worthwhile to explore developing a similar tool using LLVM’s sanitizer APIs.
Multi-language applications also present a challenge for static verification tools. Several verifiers rely on Rust’s ownership model to provide aliasing information that can simplify the process of specifying higher-order, functional properties (Lehmann et al., 2023; Astrauskas et al., 2019). However, this is only effective when unsafe code is rare. Multi-language applications are exposed to significant sources of unsafety, since C and C++ lack static aliasing restrictions. Our interview participants would benefit from static analysis and inference tools to ensure that static verification remains usable in this context. Techniques such as bi-abduction (Calcagno et al., 2011) would be useful for inferring specifications from unannotated foreign libraries. Contributors to Clang have already proposed a form of lifetime inferences for C++, which would provide annotations that could be used to generate Rust encapsulation with equivalent lifetime semantics (Brænne et al., 2022). Similar analyses would be helpful for determining if foreign types can implement the Send or Sync traits and to warn when functions retain copies of pointers derived from reference-type arguments.
Sound Verification
Despite the need for guarantees of soundness, only 10 survey respondents had used formal methods tools. One possible explanation for the lack of responses to this question is that only Verus (Lattuada et al., 2023) and Kani (VanHattum et al., 2022) had any significant support for unsafe code when we distributed the survey, and Verus had been published only several months prior. Since then, Verus has added support for programs that manipulate raw pointers, and both Gillian-Rust (Élie Ayoun et al., 2024) and RefinedRust (Gäher et al., 2024) have provided alternative methods of verification that specifically target unsafe code. However, none of these tools use formal models that are directly compatible with the aliasing models implemented in Miri, which was used by the majority of our respondents.
To ensure widespread adoption, verification tools could build on the methods and best practices that the majority of our participants already used. Improvements to Miri are likely to be the most effective way to reach our audience. For example, if Miri was extended to support symbolic execution, it provide up to 61% of our survey respondents with guarantees of soundness without requiring them to learn to use a new tool. The Seer (Renshaw and Canciani, 2018) symbolic execution engine was capable of this at one point, but it was created before Miri had support for detecting aliasing violations (Jung et al., 2019), and the project appears to have been abandoned. It would be worthwhile for the Rust community to resurrect this project or implement an equivalent analysis within Miri. Additionally, certain approaches to static verification will inevitably require developers to write specifications, which can be burdensome. However, 90% (6.5) of our respondents who exposed unsafe APIs claim to have already provided informal specifications as comments or documentation. There may be a niche for a lightweight, formal specification language targeted at annotating unsafe APIs with their core safety properties.
Education
Only 6% (6.5) of survey respondents indicated that the Rust community’s guidance and documentation were always enough to determine how to use unsafe code correctly. While the Rustonomicon (Rust Community, 2023b) is an invaluable entry point for learning about Rust’s unsafe features, we expect that its one-size-fits-all approach will inevitably leave certain users wanting more targeted information. We agree with Höltervennhoff et al. (Höltervennhoff et al., 2023) that one of the most impactful ways to improve Rust’s documentation would be to provide a dedicated resource that succinctly summarizes points of contention about Rust’s semantics.
Future interventions should also consider that necessity, performance, and ergonomics can each contribute to a single use case for unsafe code. Understanding these motivations may be helpful for effectively communicating safe alternatives. For example, a profiling tool that integrates information on the distribution and coverage of unsafe code could make it easier to perform a cost-benefit analysis on the tradeoffs between safety and performance. We also expect that lints or static analysis tools could help developers eliminate unnecessary uses of unsafe code, since 88% (6.5) of participants only sometimes refactored their code to remove unsafe features. However, refactoring may not have been necessary for these participants if they had only minimally used unsafe code.
8. Conclusion
The Rust programming language provides static safety guarantees. However, Rust is frequently used to interoperate with other languages. To call foreign functions, developers need to use Rust’s unsafe features, which bypass its safety restrictions. These features can introduce bugs and security vulnerabilities if they are used incorrectly. Few Rust-specific development tools provide comprehensive support for unsafe code, and even fewer support multi-language applications. Developers will need new solutions for testing, bug finding, and verification in this context. We conducted a mixed-methods, exploratory study to determine which interventions will be effective for developers who use unsafe code. We interviewed 19 developers who regularly wrote or edited unsafe code and created a survey that reached an additional 160 developers who engaged with unsafe code in some capacity. We examined the challenges associated with interoperation, limitations in tooling, developers’ motivations for using unsafe code, and how they reason about encapsulation.
The majority of participants used unsafe code to call foreign functions. Certain foreign libraries had aliasing and concurrency patterns that conflicted with the expectations of Rust’s type system, making them difficult to encapsulate. However, participants did not have a reliable solution for determining if their encapsulations were sound. Miri, a Rust interpreter, was the de facto solution for detecting undefined behavior, but participants who had used it before were deterred from using again it due to its slow performance and lack of support for key features. Most participants were motivated to use unsafe code by necessity, but they also cited performance and ergonomics as motivations. The majority followed best practices such as encapsulating and documenting unsafe code, but many were somewhat uncertain about their design patterns, and few felt that Rust’s documentation was always adequate for understanding how to use unsafe code correctly.
To resolve this uncertainty, our participants will need verification and analysis tools that are capable of providing guarantees of soundness and have robust support for multi-language applications. We expect that tools which integrate with Miri or use preexisting documentation or can generate specifications from documentation will be uniquely effective at reaching the majority of our audience. Any interventions should consider that developers may be motivated by a combination of necessity, performance, and ergonomics.

### v4

@@
-A Mixed-Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling
+A Mixed Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling
 Abstract.
 The Rust programming language restricts aliasing to provide static safety guarantees.
-However, in certain situations, developers need to bypass these guarantees by using a set of unsafe features.
-If they are used incorrectly, these features can reintroduce the types of safety issues that Rust was designed to prevent.
+However, in certain situations, developers need to bypass these guarantees by using a set of “unsafe” features.
+When these features are used incorrectly, they can reintroduce the types of safety issues that Rust was designed to prevent.
 We seek to understand how current development tools can be improved to better assist developers who find it necessary to interact with unsafe code.
 To that end, we study how developers reason about foreign function calls, the limitations of the tools that they currently use, their motivations for using unsafe code, and how they reason about encapsulating it.
-We conducted a mixed-methods investigation consisting of semi-structured interviews with 19 developers, followed by a survey that reached an additional 160 developers.
-Our participants were motivated to use unsafe code when they perceived that there was no alternative, and most avoided using it.
+We conducted a mixed methods investigation consisting of semi-structured interviews with 19 developers, followed by a survey that reached an additional 160 developers.
+Our participants were motivated to use unsafe code when they perceived that there was no alternative, and most claimed that they would avoid using it.
 However, limited tooling support for foreign function calls made participants uncertain about their design choices, and certain foreign aliasing and concurrency patterns were difficult to encapsulate.
-To overcome these challenges, Rust developers need verification tools that can provide guarantees of soundness within multi-language applications.
-Rust, unsafe, memory safety, threat safety, interoperation, foreign functions, motivations, tooling, interview, survey, mixed-methods
-copyright: noneccs: Software and its engineering
+To overcome these challenges, developers will need analysis tools that can find Rust-specific forms of undefined behavior within multilanguage applications.
+$c$s: Software and its engineering
 1. Introduction
 The Rust programming language has made rapid headway as a safe alternative for systems programming due to its static safety guarantees.
-However, it is still the case that critical systems are written in languages that do not statically prevent memory and thread safety issues.
-Developers who interoperate with these programs must use Rust’s unsafe features, which bypass its safety restrictions.
-To minimize the risks of using these features, the Rust community advocates for keeping unsafe code minimal, well documented, and hidden beneath a safe interface (Astrauskas et al., 2020).
-However, these practices alone are not enough to ensure that Rust fulfills its promise of static safety.
-Errors in “safe” encapsulations of unsafe implementations have been a significant source of bugs and security vulnerabilities in the Rust ecosystem (Qin et al., 2020; Xu et al., 2021; Zheng et al., 2023).
+However, it is still the case that critical systems are written in languages that do not provide inherent safety guarantees (Rust for Linux, 2026; Jansens, 2023).
+Rust developers who interoperate with these applications must use a special subset of “unsafe” features, including calling foreign functions and accessing memory through raw “C-style” pointers.
+These features can only be used within a block or function marked with the unsafe keyword.
+To minimize the risks of unsafe code, the Rust community advocates for keeping it minimal, well documented, and hidden beneath a safe interface (Astrauskas et al., 2020).
+However, these practices alone are not enough to ensure that Rust delivers on its promise of static safety.
+Errors in safe encapsulations of unsafe implementations have been a significant source of bugs and security vulnerabilities in the Rust ecosystem (Qin et al., 2020; Xu et al., 2021; Zheng et al., 2023).
 Applications that rely on interoperation or direct access to system resources may expose developers to sources of unsafety that cannot easily be minimized.
-Making matters worse, Rust’s semantics are still evolving, so few tools are capable of fully testing or verifying these applications.
+Making matters worse, Rust’s static and dynamic semantics are still evolving (Villani et al., 2025; Rémy Rakic, 2023), so few tools are capable of fully testing or verifying these applications.
 Interoperation is a common use case for unsafe code (Evans et al., 2020; Astrauskas et al., 2020; Fulton et al., 2021; Höltervennhoff et al., 2023), but most Rust-specific development tools lack support for foreign function calls.
-Safe interoperation will be crucial to Rust’s future, since both industry and regulatory stakeholders have advocated for adopting Rust as a replacement for existing systems languages (Wallach, 2024; Rebert and Kern, 2024).
-Automated translation tools are increasingly effective (Larsen, 2018), but for the moment, interoperation is the most practical solution for adopting Rust within preexisting, large-scale systems (Rebert and Kern, 2024).
-It can be challenging to use Rust correctly in this context, because it has significant differences from other languages (Höltervennhoff et al., 2023; McCormack et al., 2024).
+Safe interoperation will be crucial for Rust’s future, since both industry and regulatory stakeholders have advocated for adopting Rust as a replacement for existing systems languages (Wallach, 2024; Rebert and Kern, 2024).
+Automated translation tools are increasingly effective (Larsen, 2018), but for the moment, interoperation is the most practical method for adopting Rust within existing, large-scale systems (Rebert and Kern, 2024).
+It can be challenging to use Rust correctly in this context, because it has significant differences from other languages (Höltervennhoff et al., 2023; McCormack et al., 2025b).
 A detailed account of these challenges is missing, and we need this to be able to develop effective tooling for safe interoperation.
@@
 What tools do Rust developers use when contributing to applications that include unsafe code, and how could tooling be improved?
-To design effective interventions for developers who use unsafe code, we need to understand what motivates them to use these features and how they reason about the soundness of their designs.
-However, most large-scale studies of these topics have analyzed patterns in source code instead of directly engaging with developers (Evans et al., 2020; Astrauskas et al., 2020).
-Interviews and surveys have provided a missing qualitative perspective, but they tend to have a limited focus on unsafe code, a small sample size, or different goals (Fulton et al., 2021; Höltervennhoff et al., 2023).
-None of these studies have directly investigated how developers use foreign functions—we aim to fill this knowledge gap.
-In addition, to serve as guidance for the design of future tools, we seek to generalize existing qualitative results to a broader population of developers.
+Developers who interoperate with Rust will need to use a variety of unsafe features in addition to foreign function calls to be able to reconcile the differences between Rust and other languages.
+Having a broader understanding of what motivates Rust developers to use unsafe code in any capacity and how they reason about soundness will be helpful to identify if there are factors beyond these differences that impact the safety of multilanguage applications.
+However, most large-scale studies on these topics have analyzed patterns in source code instead of directly engaging with developers (Evans et al., 2020; Astrauskas et al., 2020).
+Interviews and surveys have provided a missing qualitative perspective, but tend to have a limited focus on unsafe code (Fulton et al., 2021), a small sample size (Evans et al., 2020), or different goals (Höltervennhoff et al., 2023).
+In addition, to serve as guidance for the design of future tools, we seek to generalize prior qualitative results to a broader population of developers.
 To that end, we ask the following additional research questions:
-RQ3 (Motivations). What are Rust developers’ motivations for using unsafe code? RQ4 (Encapsulation). How do Rust developers reason about encapsulating unsafe code?
+RQ3 (Motivations).
+What are Rust developers’ motivations for using unsafe code?
+RQ4 (Encapsulation).
+How do Rust developers reason about encapsulating unsafe code?
 Contribution
@@
 We conducted semi-structured interviews with 19 developers and used the results to create a community survey that had 160 valid responses.
-The majority of our participants used unsafe code to call foreign functions, and they described several aliasing and concurrency patterns that made interoperation difficult.
-Many developers had used Miri, an interpreter, to find bugs, but most were deterred from using it again due to its slow performance and lack of support for key features.
-Participants prioritized keeping unsafe code minimal, encapsulated, and well documented, but they rarely audited their dependencies and relied on ad-hoc reasoning to justify their decisions.
-This correctness gap can be addressed by creating verification and bug finding tools that have robust support for multi-language applications and are capable of providing guarantees of soundness.
+Most of our participants used unsafe code to call foreign functions, and they described several aliasing and concurrency patterns that made interoperation difficult.
+Many developers had used Miri (Jung et al., 2026), a Rust interpreter, to find bugs, but most were deterred from using it again due to its slow performance and lack of support for key features.
+Participants prioritized keeping unsafe code minimal, encapsulated, and well documented, but rarely audited their dependencies and relied on ad hoc reasoning to justify their decisions.
+Expanded tooling support for multilanguage applications and improved documentation will be necessary to support developers in maintaining Rust’s safety guarantees across foreign function boundaries.
 Outline
-In Section 0, we describe the role of unsafe code in the Rust ecosystem and describe the categories of tools that Rust developers can use for testing, bug finding, and verification.
-In Section 0, we provide an overview of related studies of unsafe Rust.
-We describe our methodology in Section 0.
-We present our results in Section 0 and discuss their implications in Section 0.
-We indicate threats to validity in Section 0 before concluding in Section 0.
-Our replication package is available via GitHub [^1] https://github.com/icmccorm/unsafe-study and FigShare.
-[^2] https://figshare.com/s/aacea766be06aa6deb97
+In Section 0, we describe the role of unsafe code in the Rust ecosystem and the challenges associated with interoperating with other languages.
+We describe our methodology in Section 0, and we indicate threats to validity in Section 0.
+We present our results in Section 0 and compare our findings with prior work in Section 0.
+We discuss the implications of our results in Section 0 before concluding in Section 0.
+Our replication package is available via Zenodo (McCormack et al., 2026).
 2. Background
-Rust’s safety guarantees build on the notion of ownership.
-Values with “copy” semantics do not have an owner and can be copied freely, while values with “move” semantics have a unique owner.
+Rust’s safety guarantees are based on the notion of ownership.
+Values with copy semantics do not have an owner and can be copied freely, while values with move semantics have a unique owner.
 Assignment transfers ownership from one alias to another, but ownership can also be borrowed by creating a reference.
 Each reference has a lifetime, which is the duration of the program where it is used.
-Rust’s borrow checker uses this property to enforce that a value can have either a unique, mutable reference or many immutable references, but not both simultaneously (Crichton, 2020).
-A value’s type can implement traits, which define how it behaves.
+Rust’s borrow checker enforces that a value can have either a unique mutable reference or many immutable references, but not both simultaneously (Crichton, 2020).
+A value’s type can implement traits that define how it behaves.
 Certain traits affect a value’s capability to be owned and borrowed.
-The Copy trait gives a value copy semantics, Send allows values to be transferred between threads, and Sync allows multiple threads to reference a value.
+Values with the Copy trait have copy semantics, values that are Send can be moved between threads, and values that are Sync can be borrowed by multiple threads.
 Rust’s aliasing rules are conservative and are not universally compatible with certain idioms of systems programming.
-When these restrictions become burdensome, developers can use the unsafe keyword to enable features that escape the constraints of Rust’s borrow checker.
-Within an unsafe context, users can dereference raw pointers, execute inline assembly instructions, modify static mutable state, implement unsafe traits, and call unsafe functions (Klabnik and Nichols, 2022).
-There are a variety of use cases for these features (Astrauskas et al., 2020).
+When these restrictions become burdensome, developers can use the unsafe keyword to enable features that are not constrained by the borrow checker.
+Within an unsafe context, users can dereference raw pointers, execute inline assembly instructions, modify static mutable state, implement unsafe traits, and call unsafe functions (Klabnik et al., 2026).
+There are a variety of uses for these features (Astrauskas et al., 2020).
 In certain situations, unsafe code can be necessary to implement design patterns that are fundamentally incompatible with the restrictions of safe references, such as shared mutable state.
-Foreign function calls, inline assembly, and intrinsics allow developers to leverage hardware acceleration and interact with external resources that are not constrained by the borrow checker.
-These features can also improve performance through bypassing certain runtime assertions, like bounds checks.
-However, if unsafe code is used inappropriately, it can reintroduce the same types of safety issues that Rust was designed to prevent, such as data races and accesses-out-of-bounds.
+Foreign function calls, inline assembly, and intrinsics allow developers to leverage hardware acceleration and interact with external resources.
+These features may also improve performance by skipping certain run-time assertions, such as bounds checks.
+However, if unsafe code is used inappropriately, it can reintroduce the same types of safety issues that Rust was designed to prevent, such as data races and accesses out-of-bounds.
 Rust’s aliasing restrictions also introduce new forms of undefined behavior.
-Though references are constrained by the borrow checker, they can also be cast into raw pointers and used freely in unsafe contexts.
-Rust’s compiler is allowed to assume that programs which use raw pointers still adhere to the rules of Rust’s aliasing model.
-If a program breaks these rules, then it may be optimized incorrectly, introducing differences in behavior that can constitute security vulnerabilities (Crichton and Vranken, 2023).
-2.1. Best Practices & Tooling
-Several best practices can help developers mitigate the risks of using unsafe code.
+Safe references are constrained by the borrow checker, but they can also be cast as raw pointers, which can only be used in unsafe contexts.
+The Rust compiler is allowed to assume that programs which use these features are still following Rust’s aliasing rules.
+Programs that break these rules may be optimized incorrectly, introducing differences in behavior that can constitute security vulnerabilities.
+Several best practices can help mitigate the risks of using unsafe code.
 When implementing an unsafe feature, developers are encouraged to follow the “interior unsafety” pattern by keeping it minimal, well documented, and hidden beneath a safe interface (Qin et al., 2020).
 Ideally, this interface should make it impossible for users to trigger undefined behavior from safe code.
-However certain safety properties are difficult or impossible to encode into Rust’s type system without leveraging external verification tools.
-If a function or trait places the burden of correctness on its users, then it should be labelled unsafe and documented with any necessary invariants.
-These and other guidelines are provided by multiple community resources.
-The Rust Reference (Rust Community, 2023a) and The Rust Programming Language (Klabnik and Nichols, 2022) each include relevant sections on unsafe code, but the Rustonomicon (Rust Community, 2023b) provides the most comprehensive overview.
-The Unsafe Code Guidelines Reference (Rust Community, 2023c) was intended to be an in-depth guide to the semantics of unsafe features, but it is incomplete and mostly unmaintained.
-However, the “Issues” section of its repository is still actively used to discuss Rust’s semantics.
-In addition to these best practices, dynamic analysis is also regularly used for validating unsafe code.
-Miri (Rust Community, 2024f), a Rust interpreter, is currently the only method for detecting violations of Rust’s evolving aliasing models, so it is seen as the “de facto” (Gäher et al., 2024) bug finding tool for Rust.
+However, certain safety properties are difficult or impossible to encode into Rust’s type system without leveraging external verification tools.
+Miri (Jung et al., 2026), a Rust interpreter, is currently the only tool that can detect when unsafe operations lead to violations of Rust’s newest aliasing rules, so it is typically seen as the “de facto” (Gäher et al., 2024) bug-finding tool for applications that use unsafe code.
 Currently, Miri can detect violations of two different models of Rust’s semantics.
-Stacked Borrows (Jung et al., 2019) was the first iteration of Rust’s aliasing rules, and it is enabled by default.
-Developers can also switch to using the newer Tree Borrows (Villani et al., 2023) model by providing a configuration flag.
+The Stacked Borrows (Jung et al., 2019) model was the first iteration of these rules; it is enabled by default.
+Developers can also switch to using the newer Tree Borrows (Villani et al., 2025) model by providing a configuration flag.
 Neither of the two models is canonical.
-Shuttle (Rust Community, 2024h) and Loom (Rust Community, 2024e) are Rust-specific fuzzing tools that permute the scheduling of threads to help detect concurrency bugs.
-Rust is also compatible with several language agnostic tools, including libFuzzer (LLVM Project, 2022), the LLVM project’s sanitizers (LLVM Project, 2024), and Valgrind (Nethercote and Seward, 2007).
-Static analysis is also useful in this context.
-Clippy (Rust Community, 2024d), a linting tool, can hint at improper use of unsafe features, but it is not a true bug finding tool.
-The Rust ecosystem does not yet have a de facto solution for static bug finding, but several research efforts have produced promising results.
-Rudra (Bae et al., 2021) can detect use-after-free errors when a Rust application panics, while MIRChecker (Li et al., 2021) can detect panics and duplicated ownership.
-The Rust Analysis Platform (RAP) (Cui et al., 2023a, 2024) can detect memory leaks and use-after-free errors.
-MIRAI aims to provide a unified solution for both static analysis and verification with abstract interpretation (Endor Labs, Rust Community, 2024).
-Other extensions leverage community knowledge to help enforce safety properties.
-These have typically been developed as plugins for cargo, Rust’s build tool.
-Both cargo-audit (Rust Community, 2024g) and cargo-deny (Rust Community, 2024a) detect security vulnerabilities in a project’s dependencies; cargo-deny is more extensible.
-The cargo-geiger (Rust Community, 2024b) plugin provides statistics on unsafe code, and cargo-vet (Rust Community, 2024c) can be used to track if code has been reviewed by a trusted party.
-Verification has been a uniquely active area of research within the Rust community.
-Here, we provide a high-level overview using several representative tools.
-Each category of tooling has a different level of compatibility with unsafe code.
-The first category includes Prusti (Astrauskas et al., 2019), Creusot (Denis et al., 2022), and Aeneas (Ho and Protzenko, 2022), each deductive verifiers, and Flux (Lehmann et al., 2023), a refinement type checker.
-These tools rely on Rust’s aliasing invariants to encode the properties of programs into their logic, so they do not support unsafe code by construction.
-The second category includes RefinedRust (Gäher et al., 2024) and Gillian-Rust, which exclusively target unsafe code.
-RefinedRust encodes encodes programs into a separation logic defined in Iris, which is automatically verified using the Lithium engine.
-Gillian-Rust relies on symbolic execution.
-Both approaches are informed by the RustBelt (Jung et al., 2017) model.
-Remaining tools are targeted at verifying arbitrary Rust programs, but they use unique approaches that provide limited support for unsafe features.
-Verus (Lattuada et al., 2023), a deductive verifier, uses linear ghost types to verify heap accesses through raw pointers.
-Kani, a model checker (VanHattum et al., 2022), is capable of detecting spatial and temporal memory safety issues.
-Though each of these tools have varying levels of support for unsafe code, none can validate programs against Miri’s evolving definitions of Rust’s aliasing model, and only Kani has support for foreign function calls.
-3. Related Work
-Table 1.
-A comparison table outlining how we contribute to prior studies of unsafe Rust.
-The column “Focus” indicates whether a work exclusively focused on unsafe code (direct) or studied the Rust ecosystem more broadly (indirect).
-{tblr}
-colspec=Q[l,m]Q[l,m]Q[l,m]$Q[c,m]Q[c,m]$ Work & Focus Methodology RQ1 RQ2 RQ3 RQ4
-Astrauskas (Astrauskas et al., 2020) Direct Code Analysis
-Cui (Cui et al., 2023b) Direct Code Analysis, Survey (n=50)
-Evans (Evans et al., 2020) Direct Code Analysis, Survey (n=20)
-Fulton (Fulton et al., 2021) Indirect Interviews (n=16), Survey (n=178)
-Höltervennhoff (Höltervennhoff et al., 2023) Direct Interviews (n=26)
-McCormack (McCormack et al., 2024) Direct Code Analysis
-Ozdemir (Ozdemir, 2016) Direct Code Analysis
-Qin (Qin et al., 2020) Direct Code Analysis
-Xu (Xu et al., 2021) Direct Code Analysis
-Zeng (Zeng and Crichton, 2018) Indirect Content Analysis, Interviews (n=3)
-Zhu (Zhu et al., 2022) Indirect Content Analysis, Survey (n=101)
-Zheng (Zheng et al., 2023) Direct Code Analysis
-[…]
-[c=1]r Combined:
-{tblr}colspec=Q[c,m]Q[l,m] Label Definition
-Provides relevant evidence, but does not directly answer our research question.
-Gives a partial answer to our research question, but we provide additional parts.
-Answers the same research question—we evaluate if their results generalize to a larger population.
-Answers the same research question.
-Does not address our research question.
-Prior studies have used a variety of methods to investigate how Rust developers use unsafe code.
-Here, we provide an overview of results that are relevant for each of our research questions.
-In Section 3.1, we describe results from studies that examined how unsafe code is written.
-We describe studies that engaged directly with Rust developers in Section 0.
-In Section 3.3, we classify every relevant prior study and outline how we make novel contributions in our answers to each research question.
-3.1. Studies of Code
-Distribution
-Surveys indicate that unsafe code is prevalent throughout the Rust ecosystem, but developers tend to use it minimally and hide it beneath safe interfaces.
-In Rust, packages and libraries are referred to as “crates,” and they are published on crates.io.
-In July of 2016, Ozdemir (Ozdemir, 2016) found that 29% of all published crates used unsafe features, but most unsafe blocks were relatively small and tightly scoped to include only unsafe operations.
-Evans et al.
-(Evans et al., 2020) observed similar patterns in September of 2018.
-The number of published crates had grown by more than ten times, but 29% still directly used unsafe code.
-They also observed that 38% of crates did not use any unsafe features, but they directly or indirectly depended on third-party crates that did.
-In January of 2020, Astrauskas et al (Astrauskas et al., 2020) found that the proportion of crates using unsafe features had shrunk to 23.6%.
-However, most instance of unsafe code were relatively straightforward and encapsulated.
-The majority of unsafe functions called by libraries were standard (e.g.
-not closures, function pointers, or trait functions) and defined within the same crate or provided by Rust’s standard library.
-More than 80% of unsafe functions were publicly visible, but most were exposed by crates that provided raw bindings to foreign libraries.
-Use Cases
-Studies have also examined applications to identify use cases for unsafe code.
-Qin et al.
-(Qin et al., 2020) studied a random sample of 600 instances of unsafe code from 10 popular Rust libraries, as well as 250 instances within safe encapsulations provide by Rust’s standard library.
-They identified three use cases for these operations; 42% were related to interoperation, 22% were for improving performance, and 14% involved sharing memory between threads.
-Astrauskas et al.
-(Astrauskas et al., 2020) used “anecdotal examples” and snippets of source code to create a taxonomy of six purposes for using unsafe code.
-These included implementing data structures with complex sharing, overcoming the incompleteness of Rust’s type system, emphasizing contracts and invariants, calling foreign functions, using intrinsics or inline assembly, and improving performance.
-Bugs & Vulnerabilities
-Discrepancies between the signatures of safe APIs and their interior unsafe implementations have been a significant source of bugs and vulnerabilities within the Rust ecosystem.
-Qin et al.
-(Qin et al., 2020) examined 170 bugs and vulnerabilities from ten high-profile Rust libraries.
-The Memory safety errors that they encountered were most frequently triggered from within safe contexts due to errors in unsafe implementations elsewhere.
-They also identified 19 examples of improper encapsulations of unsafe code in both the third-party libraries that they sampled and within Rust’s standard library.
-Xu et al.
-(Xu et al., 2021) exclusively studied registered vulnerabilities and identified similar errors in the encapsulations of unsafe code.
-Zheng et al.
-(Zheng et al., 2023) conducted the broadest empirical study of security issues in the Rust ecosystem to date.
-They examined 433 unique vulnerabilities that were discovered between 2014 and 2022, identifying 300 affected repositories.
-The portions of code that were affected by a vulnerability were significantly more likely to include unsafe blocks or functions.
-The overwhelming majority of all vulnerabilities affected safe functions.
-It can be difficult to correctly implement an encapsulation because unsafe implementations require a variety of safety invariants.
-Cui et al.
-(Cui et al., 2023b) manually reviewed 416 unsafe APIs in Rust’s standard library and identify 19 categories of preconditions and postconditions.
-They evaluated their taxonomy with a survey of 90 Rust developers, and also categorized 186 vulnerabilities based on the safety properties that were violated.
-The most common violations were related to incorrect uses of Send or Sync traits, incomplete initialization, or accesses-out-of-bounds.
-Interoperation
-Interoperation has consistently been one of the most prevalent use cases for unsafe code.
-Astrauskas et al.
-(Astrauskas et al., 2020) found that 44.6% of all publicly-exported unsafe functions were static bindings to foreign functions.
-Other surveys of the Rust ecosystem have similar conclusions; Ozdemir (Evans et al., 2020) found that 30% of unsafe blocks were entirely dedicated to foreign function calls, and Evans et al.
-(Evans et al., 2020) found that 22.5% of unsafe function calls were to foreign functions.
-Few studies have examined errors that occur due to the interaction between Rust and other languages.
-Cui et al.
-(Cui et al., 2023b) explicitly excluded foreign functions from their investigation of safety properties.
-Xu et al.
-(Xu et al., 2021) found 12 examples of bugs and vulnerabilities that involved foreign function calls.
-They were related to application-specific invariants, data races, invalid alignment, and inconsistent layouts.
-The authors provided few details on the nature of these bugs, indicating that they were “straightforward” and not Rust-specific.
-Only two studies have identified bugs that meaningfully involve Rust’s unique aliasing restrictions.
-In 2023, a set of anonymous authors categorized 22 bugs from five popular Rust encapsulations for foreign libraries (Anonymous, 2023).
-Most of the errors that they investigated involved incorrect handling of exceptions and heap objects that were improperly shared across foreign boundaries.
-They identified two bugs related to Rust’s aliasing restrictions, but they did not describe these bugs in terms of Rust’s operational semantics.
-McCormack et al.
-(McCormack et al., 2024) are the only group to have studied violations of Rust’s aliasing models in multi-language applications.
-They created a prototype extension to Miri (Rust Community, 2024f) and used it to find 48 bugs in Rust libraries that called foreign functions.
-Half of these bugs were violations of Rust’s Tree Borrows (Villani et al., 2023) aliasing model.
-A few were found in high-profile libraries, including one maintained by the Rust project.
-3.2. Studies of Developers
-Content Analysis
-Studies of online discourse have also provided relevant insights into encapsulation and interoperation.
-Zeng and Crichton (Zeng and Crichton, 2018) studied the challenges associated with adopting Rust.
-They interviewed three Rust contributors and analyzed a set of articles and comments posted on Hacker News and the Rust subreddit (/r/rust).
-They indicated that it was more common for developers to discuss rewriting entire applications in Rust instead of porting them incrementally, suggesting that the difficulties of interoperation may prevent developers from taking this route.
-They also identified that developers who struggled with Rust’s aliasing rules were often pointed toward safe encapsulations of unsafe operations.
-Zhu et al.
-(Zhu et al., 2022) observed a similar phenomenon on Stack Overflow.
-They used a mixed-methods approach to understand how developers reason about Rust’s aliasing restrictions.
-In the first stage of their investigation, they examined a random sample of 100 Stack Overflow questions to identify situations where developers asked for help fixing code that had been rejected by the borrow checker.
-Only three of the 110 borrowing errors in these questions were fixed by writing unsafe code.
-Developers were referred to safe alternatives in every other situation.
-Qualitative & Mixed-Methods Studies
-Fulton et al.
-(Fulton et al., 2021) studied the benefits and drawbacks of adopting Rust as a general-purpose systems programming language.
-They used an exploratory, mixed-methods design that is similar to ours.
-First, they interviewed 16 developers in industry who had advocated for using Rust.
-Then, they distributed a community survey covering the same topics as the interviews, and it received 178 responses.
-The majority of their participants struggled with Rust’s steep learning curve, but error messages from the compiler and community resources made it easy for them to find solutions to their problems.
-Participants felt somewhat more productive using Rust, and they were more confident that their implementations were correct.
-Fulton et al.
-(Fulton et al., 2021) only briefly discussed unsafe features.
-Their participants typically used unsafe code to call foreign functions, improve performance, and interact directly with hardware or the operating system.
-Höltervennhoff et al.
-(Höltervennhoff et al., 2023) are the only other group that we are aware of who have used qualitative methods to study how developers use unsafe code.
-They interviewed 26 developers with experience using unsafe Rust.
-Except for three who piloted the interview protocol, all of their participants had contributed unsafe code to open-source projects included in the unofficial “Awesome Rust” list (Community, 2022) of applications.
-We cover similar topics, but our research questions reflect different, complementary goals.
-They focused on understanding how developers contend with the security implications of using unsafe code, while we are focused on interoperation and tool use.
-Most of their participants relied on common sense and test coverage to evaluate the security of their code, and few had contributed to projects with formal policies on security or code review.
-However, most indicated that that they would review unsafe code more carefully than safe code, ensuring that it upheld all necessary safety invariants.
-Few participants had ever encountered security issues caused by unsafe code, and none of the issues that participants did experience were reported as critical vulnerabilities.
-We provide additional details on the results that are relevant for answering our research question in Section 0 and throughout Section 0.
-3.3. Our Contributions
-In Table 0, we use multiple descriptors to classify each of the relevant studies that we presented in previous sections.
-We label the methods used by each study as code analysis, content analysis, surveys, and interviews.
-We define code analysis to include mining software repositories or gathering information from bug reports and vulnerabilities.
-Content analysis includes examining any other type of preexisting online content, such as forum posts, blog posts, articles, or online discussions, to gather insights about Rust developers.
-Studies with a direct focus on unsafe code are similar to ours in that each of their core research questions are targeted at understanding how developers use unsafe features.
-Studies with an indirect focus are more broadly concerned with Rust but also happen to provide relevant insights for answering our research questions.
-We also use several labels to indicate whether each study can fully answer our research questions.
-The empty circle ‘ ’ indicates that a study does not answer a question.
-The label ‘ ’ indicates that a study provides relevant evidence, but it does not address the same question.
-The label ‘ ’ indicates that a study partially answers our question, but we provide missing parts.
-The label ‘ ’ indicates that a study fully answers our question, but it does not provide findings that generalize to a population that is comparable to our survey population.
-Now, we describe our rationale for the labels that we chose and we indicate how we provide new results in our answers to each question.
-Interoperation (RQ1)
-Studies of bugs, vulnerabilities, and the distribution of unsafe code all indicate that foreign function calls are a potentially significant unchecked source of undefined behavior in the Rust ecosystem.
-However, none of these studies provide insights into the challenges that developers face when implementing these applications.
-Most of the participants studied by Fulton et al.
-(Fulton et al., 2021) indicated that C and Python were easy to interoperate with, but C++ was difficult to handle due to complex aliasing.
-Participants interviewed by Höltervennhoff et al.
-(Höltervennhoff et al., 2023) also had mixed experiences with interoperation; four described it as being easy, but two struggled with “memory management, e.g.
-between programming languages.” These studies can only partially answer our research question ( ), since neither provides a detailed description of the patterns that lead to these difficulties, other than that they involved aliasing and initialization (Höltervennhoff et al., 2023; Fulton et al., 2021).
-We have an explicit focus on interoperation as one of our core research questions, and we provide a missing perspective on the challenges of using Rust in multi-language applications.
-Tooling (RQ2)
-Rust is a relatively new programming language, so the majority of research efforts have focused on developing new tools instead of evaluating existing ones.
-Zeng and Crichton (Zeng and Crichton, 2018) hypothesized that a lack of promotion for tooling could prevent Rust from being widely adopted as a systems language.
-However, their work was published prior to the creation of most Rust-specific bug finding tools.
-Höltervennhoff et al.
-(Höltervennhoff et al., 2023) are the only other group that we are aware of who have examined how developers use tooling to assist with writing unsafe code.
-Key development tools were somewhat underused among their interview population—only 7 of their 26 participants had mentioned using Miri.
-Since this is a small sample size, it is unclear whether these results are indicative of a broader trend in the Rust community, so they can only partially answer our second research question ( ).
-We also cover debuggers, static verifiers, and tools for generating bindings to foreign functions, which were not mentioned by Höltervennhoff et al.
-(Höltervennhoff et al., 2023).
-Motivations (RQ3)
-Several prior studies have conflated developers’ motivations for using unsafe code with features-oriented use cases.
-For example, Astrauskas et al.
-(Astrauskas et al., 2020) include both performance and compiler intrinsics as “motivations” for using unsafe code.
-In their survey of 20 developers, Evans et al.
-(Evans et al., 2020) defined performance, ease-of-use, and a variety of specific operations as “reasons” for using unsafe code.
-Fulton et al.
-(Fulton et al., 2021) include both performance, interaction with hardware, and foreign function calls as “uses” for unsafe code.
-We define “motivations” as feature agnostic reasons for using unsafe code.
-This perspective is similar to Höltervennhoff et al.
-(Höltervennhoff et al., 2023), who identified necessity and performance, and efficiency as “motivations,” but the separation between features and higher-level reasons was not an explicit design consideration of their study.
-When viewed together, these results answer our third research question ( ).
-However, we provide a novel contribution by replicating their findings and evaluating whether they generalize to a large population of developers.
-Additionally, Höltervennhoff et al.
-(Höltervennhoff et al., 2023) study developers’ motivations as part of examining the broader topic of the “common practices for deciding for and implementing” unsafe code.
-We dedicate an entire research question to this topic, so we are able to provide significantly more detail on how specific design decisions relate to each motivation.
-Encapsulation (RQ4)
-Surveys of the Rust ecosystem suggest that developers typically follow best practices for minimizing and encapsulating unsafe code, but they struggle with ensuring that their interfaces are correct in all situations.
-Höltervennhoff et al.
-(Höltervennhoff et al., 2023) are the only prior study to have examined developers experiences with designing applications that include unsafe code, and their results reflect the patterns observed surveys of code and bug reports.
-The majority of their participants would avoid using unsafe code whenever possible.
-When unsafe code was necessary, they ensured that it was encapsulated beneath a safe interface of their own design or provided by an external library.
-However, the majority of participants felt that it was difficult to write unsafe code correctly; they had a “nagging feeling about it”, and half were unsure or undecided about its correctness.
-Similar to RQ3, these results also answer our research question ( ), but we provide a unique contribution by evaluating whether their results generalize to our survey population.
-We also use the set of motivations we defined in our answer to RQ3 to investigate developers’ motivations for exposing unsafe APIs, which is not addressed by Holtervennhoff et al.
-(Höltervennhoff et al., 2023).
+These practices are standard for the Rust community, but Rust is also increasingly being adopted in interoperation with preexisting large-scale C and C++ applications, such as Chromium (Jansens, 2023) and the Linux Kernel (Rust for Linux, 2026).
+These applications will expose Rust components that may be “safe” in isolation to significant external sources of unsafety, which will not be as easy to minimize or document.
+In addition to these challenges, Miri does not fully support finding undefined behavior in this context.
+Miri can execute and trace certain foreign functions from shared libraries, but it cannot detect aliasing violations that are triggered in foreign code, and its slow performance prevents it from being used at scale.
+These obstacles will make it difficult for developers to avoid introducing undefined behavior in multilanguage Rust applications.
+Projects such as the DARPA’s TRACTOR program (Wallach, 2024) and the Rust/C++ Interoperability Initiative (Bauman and The Rust Foundation, 2024) seek to provide new development tools to assist with these challenges.
+We need a broader understanding of how Rust developers use unsafe code to be able to design interventions that are effective in practice, since developers who use Rust in this context will necessarily engage with a variety of unsafe operations in order to reconcile differences in semantics across language boundaries.
 3. Methodology
-We adapted a sequential, exploratory mixed-methods design (Creswell et al., 2003) consisting of three phases.
+We investigated how Rust developers engage with unsafe code by adapting a sequential, exploratory mixed methods design (Creswell et al., 2003), which consisted of three phases.
 We describe the first phase in Section 0, where we used a screening survey to recruit developers for semi-structured interviews.
@@
     Do your development tools handle all of the problems that you face writing unsafe Rust?
-(c) Interview participants’ affiliations ("Affil."), their years of experience using Rust ("Years"), and the number of coding decisions made for their interview (Decs.)
-ID Affil. Years Decs.
-1 I 2 56
-2 I,R 10 76
-3 I,O 5 93
-4 I,O 8 71
-5 I,O 2 61
-6 A 3 71
-7 I 6 57
-8 I 5 49
-9 O 2 97
-10 O 2 59
-11 I 5 86
-12 I,A 6 121
-13 I 10 92
-14 R,O,I 7 94
-15 I 1.5 58
-16 I 3 20
-17 A 4 64
-18 I 3 82
-19 I 4 13
-I = Industry  R = Rust Team
-A = Academia  O = Open Source
-(d) Codes for the theme "Why unsafe?", which describes participants’ motivations for using unsafe code.
+(c) Codes for the theme "Why unsafe?", which describes participants’ motivations for using unsafe code.
 Code Definition Participant Quote
@@
 P10 "I use unsafe Rust for mostly to create new data structures or types that are not really possible to express with safe Rust."
-(e) A survey question with a direct correspondence to the codes shown in Figure 0.
+(d) A survey question with a direct correspondence to the codes shown in Figure 0.
 Think about the situations where you have typically used unsafe. Which of the following reasons have motivated you to do so?
@@
 Certain questions were gated based on prior responses.
-If an candidate answered "No" to questions #0 and #0 of Figure 0, then they did not use Rust in multi-language applications, so we did not ask them questions #0-0 of the screening survey or question #0 of the interview protocol, since each of these questions involve foreign functions.
+If an candidate answered "No" to questions #0 and #0 of Figure 0, then they did not use Rust in multilanguage applications, so we did not ask them questions #0-0 of the screening survey or question #0 of the interview protocol, since each of these questions involve foreign functions.
 The first author conducted all 19 interviews remotely over Zoom between May and July of 2023.
+Participants were not required to enable their camera.
 We used snowball sampling (Patton, 1990) to expand our pool of participants beyond the group that responded to our initial calls to participate.
-If a participant shared new information that we had not yet learned, then we prompted them to invite any acquaintances with relevant experience to complete our screening survey.
-We stopped recruiting new participants after we had reached “data saturation” as defined by Saunders et al.
-(Saunders et al., 2018); we began to hear the same or similar responses to each of our questions.
-We transcribed the audio recordings of the interviews using Whisper (Radford et al., 2023), and the first author reviewed every transcript to correct errors and label whether utterances came from the interviewer or the interviewee.
+If a participant shared new information that we had not learned yet, then we prompted them to invite any acquaintances with relevant experience to complete our screening survey.
+We stopped recruiting new participants after we had reached saturation.
+Our concept of saturation corresponds to Saunders’ et al (Saunders et al., 2018)’s definition of data saturation as a subjective judgement made during data collection before analysis.
+We knew that we had reached this point after we began to “hear the same comments again and again” (Grady, 1998) and had informally identified several themes from our interviews.
+Two themes were particularly notable at this stage; certain participants felt that Rust’s Box could be cumbersome due to its noalias semantics (Section 0), and that Miri was difficult to use in multilanguage applications (Section 0).
+Our sample size (19) is comparable to other interview studies with Rust developers (Fulton et al., 2021; Höltervennhoff et al., 2023).
+We transcribed the audio recordings of the interviews using Whisper (Radford et al., 2023), and the first author reviewed every transcript to correct errors and label whether the utterances came from the interviewer or the interviewee.
 3.2. Qualitative Analysis
-We conducted an inductive, thematic analysis of our interview transcripts following best practices by Miles et al.
-(Miles et al., 2020).
+We conducted an inductive, thematic analysis of our interview transcripts following best-practices by Miles, Huberman, and Saldaña (Miles et al., 2020).
 To begin, the first three authors conducted open coding of random samples of quotes taken from each interview.
@@
 If all responses at that index had been sampled before, then the first author generated a new random index and repeated this process until a new quote was found.
-We coded three samples of quotes, and we met after each sample to resolve disagreements and record new codes within a shared codebook.
+We coded three samples of quotes and met after each sample to resolve disagreements and record new codes within a shared codebook.
 Next, the first and second authors used a card-sorting (Braun and Clarke, 2006) exercise to identify higher-level themes among the codes that we had generated in the first phase.
@@
 We used the themes that we created from this process to organize our codebook, and we removed redundant or irrelevant codes.
-Then, the first and second authors used this codebook to conduct closed coding of 7 additional random samples of quotes, meeting after coding each sample to resolve disagreements and update the codebook.
+The first and second authors used this codebook to conduct closed coding of seven additional random samples of quotes, meeting after coding each sample to resolve disagreements and update the codebook.
 We used the same process as before to create samples, but instead of treating each interview transcript separately, we sampled quotes from a single document containing the concatenated text of every interview transcript.
@@
 After all transcripts had been coded by one of the two authors, the first author conducted an additional coding pass for every transcript and resolved each of their disagreements with the second author.
-The fourth column of Table 0 shows the final number of coding decisions made for each participants’ transcript.
 The complete codebook and all coding decisions are included in our replication package.
@@
 Similar to our screening survey, certain questions were gated based on prior responses.
-We provide a complete copy of our survey, including its branching logic, in Section 4 of our Appendix.
+We provide a complete copy of our survey, including its branching logic, in our Appendix (McCormack et al., 2026).
 Questions were either descriptive or exploratory.
 Figure 0 shows an example of a descriptive question.
-Each of its response options correspond to the codes shown in Figure 0 for the theme “Why unsafe?”.
+Each response option corresponds to the codes shown in Figure 0 for the theme “Why unsafe?”.
 Exploratory questions arose after we had finished conducting interviews, and they did not have a direct correspondence to the themes that we identified during qualitative analysis.
@@
 (Höltervennhoff et al., 2023)’s study was published concurrent to our investigation, and three of their participants mentioned using either cargo-audit or cargo-deny to detect vulnerabilities in their dependencies.
-None of our interview participants had mentioned using these tools, but we included a survey question asking whether respondents had used them before.
+None of our interview participants had mentioned using these tools, but we included a survey question asking if respondents had used them before.
 We also relaxed our eligibility criteria for the survey to include individuals who would have been eligible for Höltervenhoff et al.
@@
 Candidates were eligible to participate in their study if they had committed unsafe code to GitHub, but there was no indication if they had done so regularly.
-Our interview participants needed to have “regularly” engaged with unsafe Rust, but our survey respondents were eligible if they had “written, edited, read, audited, or engaged with unsafe Rust code in any way.”
+Our interview participants needed to have “regularly” engaged with unsafe code, but our survey respondents were eligible if they had “written, edited, read, audited, or engaged with unsafe Rust code in any way.”
 Two of our interview participants piloted an initial draft of the survey, and we incorporated their feedback into the final version.
 We distributed our community survey on September 22nd, 2023 to the same acquaintances and communities where we distributed our screening survey.
-We also published a link to our survey in the September 28th edition of the “This Week in Rust” newsletter, and we advertised the survey during a presentation and poster sessions at an industry partners conference hosted by our institution.
+We also published a link to our survey in the September 28th edition of the “This Week in Rust” newsletter and we advertised the survey during a presentation and poster sessions at an industry partners conference hosted by our institution.
 At the end of the survey, participants could provide a link to their profile on either GitHub or the Rust Programming Language Forum to have a chance to receive one of two \$250 gift cards.
-Only participants with public account activity prior to the start date of the survey were eligible for compensation.
-We closed the survey on October 9th, 2023.
+We randomly selected two participants who had provided a link to their profile with account activity prior to the start date of the survey.
+We configured Qualtrics (Qualtrics, 2024) to reject multiple responses from the same individual.
+We also enabled reCaptcha (Liu, 2018) and RelevantID (RelevantID, 2026) to detect possibly fraudulent submissions.
+RelevantID uses several forms of metadata, including location and IP information; none of this information was retained or visible to the investigators.
+In our analysis, we excluded responses that had a reCaptcha score of less than 0.5, indicating that the response was likely from a bot.
+For RelevantID, we excluded responses that had a fraud score greater than 0.3 or were marked as likely to be duplicate.
 4. Ethics & Threats to Validity
@@
 All procedures in this study were approved by our institution’s IRB.
-Prior to each intervention, participants were read or presented a script outlining the procedures of the study and they were asked to affirm consent.
+Before each intervention, participants were read or presented a script outlining the procedures of the study and they were asked to affirm consent.
 Participants were reminded to avoid sharing personally identifying information in their responses, and interview participants were asked to remain in a private space.
@@
 We attempted to maintain a diverse survey and interview population by sampling from multiple forums.
-However, both Reddit and “This Week in Rust” were over-represented compared to the other communities where we distributed our survey.
-We cannot determine if our interview population has a similar bias, since we did not record where interview participants had first learned about our study.
+However, both Reddit and “This Week in Rust” were overrepresented compared to the other communities where we distributed our survey.
+Our sample does not include developers who are not active in these communities.
+We cannot determine if our interview population has a similar bias, since we did not record where these participants had first learned about our study.
+The "silent" developers excluded from these populations may have different experiences and perspectives that are not represented in the current findings.
+Rust’s unsafe “superpowers” are typically advertised as an advanced feature that developers “might run into every once in a while, but may not use every day” (Rust Community, 2023a).
+Developers who have engaged with unsafe code may more likely to have had more experience using Rust in any capacity than those who have not used unsafe code.
+We suspect that committed Rust experts are overrepresented in our interview population, since many were affiliated with prominent open source projects or had experience using the language within security-critical applications.
+It is more difficult to evaluate the capabilities of our survey respondents.
+There was significant variance in the reported number of years of experience.
+However, on average, our survey population had nearly the same number of years of experience using Rust as our interview population, and more than a decade of experience within the field of software engineering.
+Most respondents reported behaviors that indicate non-trivial experience with Rust, such as using bug-finding or auditing tools and committing unsafe code to a published library.
+For this reason, we expect that our current results reflect the “best-case scenario” for developers’ experiences using unsafe Rust.
+Any aspects of Rust development that our population found to be difficult seem likely to be even more challenging for the broader Rust community.
 Additionally, the gift cards that we offered for compensation are most easily redeemable in North America, which may have dissuaded certain individuals from participating in either stage of our study.
 We did not collect location data from participants, so we cannot determine if our results are biased toward individuals from any particular area.
-Respondents were eligible to complete our survey if they had one year of experience and had “engaged” with unsafe Rust in any way.
-A subset may have been eligible without having any significant practical experience with unsafe code.
+Construct Validity
+Respondents were eligible to complete our survey if they had one year of experience and had “engaged” with unsafe code in any way.
+A subset may have been eligible without having significant practical experience with unsafe code.
 We mitigated this by gating questions based on whether participants had used certain features.
 When relevant, we compare results from the entire population with results from the subset who “regularly” wrote or edited unsafe code, which matches the eligibility criteria for our interviews.
-This criteria may have introduced self-selection bias among our interview participants, since individual who frequently use unsafe code may be more likely to perceive that it is necessary for their purposes.
-Construct Validity
+This criteria may have introduced self-selection bias among our interview participants, since individuals who frequently use unsafe code may be more likely to perceive that it is necessary for their purposes.
+We did not enable reCaptcha (Liu, 2018), RelevantID (RelevantID, 2026), or any of Qualtrics’ other safeguards when recruiting participants for interviews.
+These safeguards are subject to false positives (Zhang et al., 2022), and may have flagged legitimate participants.
 The validity of our results depends on participants’ ability to accurately self-report their behavior, and what developers perceive about their use of unsafe code may differ from how it is used in practice.
-Our qualitative results are unavoidably influenced by the individual biases of each investigator.
-We attempted to mitigate this through auditing coding decisions and meeting frequently to resolve disagreements.
-When we distributed our community survey, we configured Qualtrics (Qualtrics, 2024) to reject multiple responses from the same individual using IP information (which was not recorded or visible to investigators), RelevantID (RelevantID, 2024) and reCaptcha (Liu, 2018).
-We did not enable these features for the initial screening survey.
-The overwhelming majority of interview participants enabled their camera during interviews, but a few did not; they were not prompted or required to do so.
-We rely on the baseline assumption that every participant was honest about their identity, but we could not independently verify this in every situation.
+We did not evaluate whether participants were able to reason correctly about unsafe code; only how they perceived their usage of these features.
+Höltervennhoff et al.
+(Höltervennhoff et al., 2023) found that developers who use unsafe code have varying degrees of awareness about its implications for correctness; 10 of their interview participants indicated that unsafe contexts entirely disable Rust’s static safety restrictions, which is incorrect.
+Our qualitative results are also unavoidably influenced by the individual biases of each investigator.
+We tried to mitigate this by auditing coding decisions and meeting frequently to resolve disagreements.
+Although most of our interview protocol was written to be neutral toward unsafe code, certain questions may have biased participants toward providing responses that favored safe coding practices, such as asking participants to describe how they “reason about memory safety” and “navigate the differences” between Rust and other languages.
 5. Results
-We found that participants used unsafe code for ergonomics, performance, and when they perceived that there was no safe alternative.
-They attempted to keep unsafe code minimal and encapsulated, but many participants were still uncertain about its correctness.
-Most participants had called foreign functions, and non-trivial differences between Rust and other languages made these calls difficult to encapsulate in certain situations.
-Miri was the most popular bug finding tool.
-However, while more than 60% of respondents had used Miri before, its performance and lack of support for key features deterred 69% of these respondents from using it again.
-Our participants would benefit from tools that can provide guarantees of soundness and that have robust support for multi-language applications.
 We report the demographics of interview and survey participants in Section 0.
@@
 We describe developers’ motivations for using unsafe code in Section 0 (RQ3), and we report how they reason about encapsulation in Section 0 (RQ4).
-We summarize key results and provide suggestions for tool designers at the end of each section.
+We summarize key results at the end of each section.
 5.1. Demographics
-[…]
-Years of Experience
-Location # SE C C++ Rust
-Rust Forums 16 17.3 ± 10.1 7.5 ± 6.8 5.6 ± 6.2 4.5 ± 2.4
-Rust Discord 10 8 ± 3.7 2.7 ± 2.8 3.4 ± 2.3 3.6 ± 2
-Reddit (/r/rust) 83 13.9 ± 9.1 7.6 ± 7.2 6 ± 6.3 4.1 ± 2.2
-"This Week in Rust" 52 15.1 ± 9.7 6.6 ± 7.7 5.7 ± 6.9 4.1 ± 2.3
-Other 3 23.7 ± 15.8 17.3 ± 14.6 9 ± 5.6 3.7 ± 1.5
-All 160 14.5 ± 9.5 7.2 ± 7.5 5.8 ± 6.3 4.1 ± 2.2
+Table 1. Demographics of interview and survey participants.
+(a) Interview participants’ affiliations ("Affil.") and years of experience using Rust ("Years") by ID.
+ID Affil. Years
+1 I 2
+2 I,R 10
+3 I,O 5
+4 I,O 8
+5 I,O 2
+6 A 3
+7 I 6
+8 I 5
+9 O 2
+10 O 2
+ID Affil. Years
+11 I 5
+12 I,A 6
+13 I 10
+14 R,O,I 7
+15 I 1.5
+16 I 3
+17 A 4
+18 I 3
+19 I 4
+I = Industry   R = Rust Team
+A = Academia   O = Open Source
+(b) The mean ($\mu$) and standard deviation ($\sigma$) of survey respondents’ years of experience in the field of software engineering ("SE"), as well as using Rust, C and C++.
+Entries are grouped by where respondents indicated that they had first heard about the survey.
+Respondents could select multiple options.
+SE C C++ Rust
+Location # $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
+Rust Forums 16 17.3 10.1 7.5 6.8 5.6 6.2 4.5 2.4
+Rust Discord 10 8.0 3.7 2.7 2.8 3.4 2.3 3.6 2.0
+Reddit (/r/rust) 83 13.9 9.1 7.6 7.2 6.0 6.3 4.1 2.2
+"This Week in Rust" 52 15.1 9.7 6.6 7.7 5.7 6.9 4.1 2.3
+Other 3 23.7 15.8 17.3 14.6 9.0 5.6 3.7 1.5
+All 160 14.5 9.5 7.2 7.5 5.8 6.3 4.1 2.2
+Table 0 provides summary statistics on interview and survey participants.
 We invited all 42 eligible candidates to participate in interviews, and we interviewed 19 who responded to our invitation.
 Table 0 summarizes interview participants’ demographic information.
-There were 15 who had experience in industry, 6 in open source development, 3 in academia, and 2 were members of the Rust Team.
-They had 4.7 years of experience with Rust on average, and the most experienced participant had been using Rust for 10 years.
-We did not collected any additional demographic information from our interview participants.
-We received 370 survey responses, of which 242 were complete, 203 met our eligibility criteria, and 160 passed each of Qualtrics’ (Qualtrics, 2024) measures for fraud detection.
-Hereafter, when we refer to “respondents,” we include only the 160 that passed every one of our checks.
+There were 15 who had experience in industry, six in open source development, three in academia, and two were members of the Rust Team.
+They had 4.7 years of experience with Rust on average, and the most experienced participants had been using Rust for 10 years.
+We did not collect any additional demographic information from our interview participants.
+We received 368 survey responses, of which 240 were complete, 203 met our eligibility criteria, and 160 passed each of our measures for fraud detection.
+Hereafter, when we refer to “respondents,” we include only the 160 that passed each of our checks.
 We collected several categories of demographic information from survey respondents.
-The majority were young, educated, identified as male, and had experience in either industry or open-source development.
+The majority were young, educated, identified as male, and had experience in either industry or open source development.
 Nearly 66% had completed at least a bachelor’s degree, and 30% had completed a graduate degree.
-Most respondents (83%) were younger than 40 years old, and 56% were between the ages of 18 and 29.
-We followed Spiel et al.’s (Spiel et al., 2019) guidelines for ethically surveying respondents’ gender; 80% identified as “Man”, 6% identified as “Non-binary”, 3% identified as “Woman” and 9% did not disclose their gender.
+Most of the respondents (83%) were under 40 years of age, and 56% were between the ages of 18 and 29.
+We followed Spiel et al.’s (Spiel et al., 2019) guidelines for ethically surveying respondents’ gender; 82% identified as “Man”, 6% identified as “Non-binary”, 3% identified as “Woman” and 9% did not disclose their gender.
 Table 0 summarizes respondents’ years of experience with the field of software engineering, Rust, and other systems languages.
-On average, respondents had more than a decade of experience in software engineering, more than 5 years of experience in both C and C++, and just over 4 years of experience using Rust.
+On average, respondents had more than a decade of experience in software engineering, more than 5 years of experience with C and C++, and just over 4 years of experience using Rust.
 Table 2.
@@
 Percentages on the left include “Less than once a year” and “Yearly”, while those on the right include “Weekly” and “Daily”.
-The number of respondents in the sample is listed in parenthesis on the righthand side.
-{tblr}
-width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
-[…]
-[Uncaptioned image]23%41%(160)
-3b How frequently do you write new Rust code or edit existing Rust code within an unsafe block or function?
-[Uncaptioned image]31%26%(160)
-[…]
+The number of respondents in the sample is listed in parentheses on the righthand side.
+ID Question Sample Distribution
+2a How frequently do you engage with unsafe Rust code in any way? All 23%41%(160)
+2b How frequently do you write new Rust code or edit existing Rust code within an unsafe block or function? 31%26%(160)
+Less than once a year    Yearly    Monthly    Weekly    Daily
 Participants had a variety of roles within the Rust community and differing levels of engagement with unsafe code.
@@
 Since “regularly” is subjective, we asked the questions shown in Table 0 to provide a more specific time-frame and to distinguish editing unsafe code from practices that do not require code contributions, such as auditing.
-Only 26% (6.1) of survey respondents would write or edit unsafe code more than once a month, but 41% (6.1) would engage with it at this frequency.
+Only 26% (2) of survey respondents would write or edit unsafe code more than once a month, but 41% (2) would engage with it at this frequency.
 5.2. RQ1 - Interoperation
@@
 (Fulton et al., 2021).
-Assembly would have been included as part of the category "Other" in their survey, which was selected by only 7% of respondents.
-Additionally, 72% of our respondents had converted part or all of an application written in another language to Rust, and 16% of these respondents had not called foreign functions.
-This supports Zeng and Crichton (Zeng and Crichton, 2018)’s notion that some developers may prefer to rewrite entire applications in Rust instead of relying on interoperation to incrementally translate them.
-However, none of our evidence suggests that this is a dominant preference.
-Memory Management
+Their survey included Assembly as part of the "Other" category, which only 7% of their respondents had selected.
+Although 72% of our respondents had converted part or all of an application written in another language to Rust, 25% of these respondents had not called foreign functions.
+This subset might have ported these applications wholesale instead of opting for an incremental approach.
+Aliasing & Memory Management
 Rust’s standard library provides several types that can be used to manage ownership of heap allocations.
@@
 Participants frequently used these types to allocate memory for foreign function calls, converting them into raw pointers in the process.
-In particular, 54% of survey respondents who had called foreign functions also used Rust’s memory container types for this purpose.
+In particular, 54% of survey respondents who had called foreign functions also used one of Rust’s memory container types for this purpose.
 Rust’s thread-safe Arc—an atomic, reference-counted allocation—was used by 28% of these participants, while only 8% had used the non-atomic Rc.
@@
 When a Box was unwrapped and exposed as a raw pointer to C, its address was added to a table.
-The table was checked and updated to ensure that each allocation was only freed once.
-Several interview participants mentioned having issues with Box, noting that when a Box is moved, any raw pointers to its contents are invalidated.
-However, participants with a background in C++ development had not expected Box to behave this way.
-They were accustomed to using unique_ptr from the C++ standard library, which has a similar role but different semantics:
-You give the raw pointer off to the interface, and then you relocate the unique_ptr, but you still rely on that pointer having valid access tags. (P0)
+The table was checked and updated to prevent double-free errors.
+Several interview participants mentioned having issues with this type, noting that when a Box is moved, any raw pointers to its contents become invalid.
+Participants with a background in C++ development had not expected Box to behave this way.
+They were accustomed to using unique_ptr from the C++ standard library.
+This has a similar role, but different semantics:
+You give the raw pointer off to the interface, and then you relocate the unique pointer, but you still rely on that pointer having valid access tags. (P9)
 Survey participants did not encounter this issue as frequently.
-Of the 51% of respondents had used Box in any unsafe context, only 17% had encountered issues when a pointer because invalid after a Box was moved.
+Of the 51% of respondents had used Box in any unsafe context, only 17% had encountered issues when a pointer became invalid after a Box was moved.
 However, we did not determine if respondents had actually unwrapped a Box or otherwise retained a raw pointer to its contents, so this 17% could still represent a significant portion of the respondents who would have been able to encounter this problem.
+Our population is most representative of experienced Rust developers, so it is possible that the average Rust developer is even more likely to have misused Rust’s Box type in these contexts.
 One participant observed several aliasing patterns in foreign libraries that were at odds with Rust’s restrictions.
@@
 Another participant observed that C libraries would often use on “a rat’s nest of pointers” (P14) to encode self-referential structures, which were difficult to represent in Rust encapsulations.
-Each of the aliasing violations from the most popular libraries tested by McCormack et al.
-(McCormack et al., 2024) were caused by incorrect encapsulations of self-referential structures.
 To manage these differences, participants reviewed the documentation and implementation of foreign libraries to determine their safety requirements.
@@
 (P11)
+This pattern matches two of the aliasing violations observed by McCormack et al.
+(McCormack et al., 2025b), where "unique" mutable references were cast into raw pointers and stored into the heap by foreign functions.
 The reverse of this situation occurred when participants exposed Rust libraries to languages with fewer safety guarantees.
@@
 Concurrency
-Interview participants typically used more sophisticated multithreading patterns in the Rust components of multi-language applications.
+Interview participants typically used more sophisticated multithreading patterns in the Rust components of multilanguage applications.
 One participant used granular locking within the Rust codebase, while the C codebase had a singular global lock.
@@
 Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”.
-The number of respondents in the sample is listed in parenthesis on the righthand side.
-{tblr}
-width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
-[…]
-[Uncaptioned image]49%24%(112)
-4b How often do you intentionally avoid converting raw pointers to memory allocated by FFI calls into safe references, such as &T or &mut T?
-[Uncaptioned image]23%43%(112)
-[…]
+The number of respondents in the sample is listed in parentheses on the righthand side.
+ID Question Sample Distribution
+3a How often do you intentionally avoid passing Rust’s abstract data types (structs, enums) by value across FFI boundaries? Used FFI 49%24%(112)
+3b How often do you intentionally avoid converting raw pointers to memory allocated by FFI calls into safe references, such as &T or &mut T? 23%43%(112)
+Always    Most of the time    About half the time    Unsure    Sometimes    Never
 Information-Hiding
 Participants attempted to keep interoperation with foreign code as straightforward as possible.
-When declaring bindings, one participant preferred using primitive types and pointers instead of abstract data types like Option and NonNull, even though Rust allows these types to be implicitly cast into raw pointers (Community, 2024).
+When declaring bindings, one participant preferred using primitive types and pointers instead of abstract data types like Option and NonNull, even though Rust allows these types to be implicitly cast into raw pointers (Community, 2026b).
 They felt that these casts hid useful contextual information, so they preferred to use raw pointers explicitly.
-A few participants indicated that they sometimes avoided passing structs by value across foreign boundaries.
+A few participants indicated that they had avoided passing structs by value across foreign boundaries in certain situations.
 One of these participants observed that certain Rust programs would behave incorrectly when structs were moved across a foreign boundary instead of copied, or if they had a destructor on the C++ side of the boundary.
@@
 We added “Unsure” as a response option for these questions since they indirectly rely on participants having a notion of why they would “intentionally” avoid these behaviors.
-Neither practice was consistently followed, but more respondents would avoid passing abstract data types by value most or all of the time (49%, 6.2), than would avoid casting foreign pointers into references (43%, 6.2).
+Neither practice was consistently followed, but slightly more respondents would avoid passing abstract data types by value most or all of the time (49%, 3), than would avoid casting foreign pointers into references (43%, 3).
 Key Findings - RQ1 (Interoperation).
@@
 Developers identified safety properties and design patterns of foreign APIs that match the characteristics of libraries where McCormack et al.
-(McCormack et al., 2024) detected aliasing violations, indicating that foreign function calls may be a significant unchecked source of undefined behavior in the Rust ecosystem.
-We expect that developers who write encapsulations for foreign APIs would benefit from analysis tools that can infer their safety properties.
-This information could be leveraged to automatically generate sound, safe encapsulations.
+(McCormack et al., 2025b) detected aliasing violations, indicating that foreign function calls may be a significant unchecked source of undefined behavior in the Rust ecosystem.
 5.3. RQ2 - Tooling
@@
 Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”.
-The number of respondents in each sample is included in parenthesis on the righthand side.
-{tblr}
-width=colspec=Q[c,m]X[l,m]$Q[c,m]$
+The number of respondents in each sample is included in parentheses on the righthand side.
 ID Question Sample Distribution
-[…]
-[Uncaptioned image]61%30%(160)
-Used Miri
-[Uncaptioned image]72%19%(98)
-5b How often do you run test cases in Miri? Used Miri
-[Uncaptioned image]38%56%(98)
-[…]
-[Uncaptioned image]6%88%(160)
-Used Auditing Tools
-[Uncaptioned image]10%80%(81)
-[…]
+4a How often do you write tests for Rust applications that use unsafe? All 61%30%(160)
+Used Miri 72%19%(98)
+4b How often do you run test cases in Miri? Used Miri 38%56%(98)
+4c How often do you audit your dependencies’ use of unsafe? All 6%88%(160)
+Used Auditing Tools 10%80%(81)
+Always    Most of the time    About half the time    Sometimes    Never
 Table 5.
@@
 Percentages on the left include “Definitely yes” and “Probably yes”, while those on the right include “Probably not” and “Definitely not”.
-The number of respondents the each sample is included in parenthesis on the righthand side.
-{tblr}
-width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
-[…]
-[Uncaptioned image]45%7%(112)
-6b Do you trust FFI bindings that are generated by a tool?
-[Uncaptioned image]79%2%(112)
-[…]
+The number of respondents the each sample is included in parentheses on the righthand side.
+ID Question Sample Distribution
+5a Do you trust FFI bindings that are written by hand? Used FFI 45%7%(112)
+5b Do you trust FFI bindings that are generated by a tool? 79%2%(112)
+Definitely yes    Probably yes    Might or might not    Probably not    Definitely not
 Participants used a wide variety of development tools to assist with validating their design choices.
-We focused primarily on dynamic tools for bug finding and static tools for generating bindings to foreign functions.
+We focused primarily on dynamic bug-finding tools and static tools for generating bindings to foreign functions.
 However, certain participants also described their experiences using debuggers, verifiers, and auditing tools.
@@
 Most participants used dynamic bug finding tools.
-Interview participants often mentioned Miri, AddressSanitizer, and Valgrind.
+Interview participants often mentioned Miri (Jung et al., 2026), AddressSanitizer (Serebryany et al., 2012), and Valgrind (Nethercote and Seward, 2007).
 We prompted survey respondents to indicate which dynamic tools they had used from a list of options mentioned by our interview participants.
-Survey respondents most frequently used Miri (61%), Valgrind (44%), AddressSanitizer (26%), and cargo-fuzz (26%).
+Survey respondents most frequently used Miri (61%), Valgrind (44%), AddressSanitizer (26%), and cargo-fuzz (Rust Community, 2025a) (26%).
 Each of the remaining tools listed in our survey was used by less than 15% of respondents.
@@
 One interview participant leveraged industry sponsorship to fuzz a JIT compiler for hours-on-end.
-Another participant implemented a BTree with optimizations for high performance computing, and they used libfuzzer to perform differential testing (McKeeman, 1998) against the implementation provided by Rust’s standard library.
+Another participant implemented a BTree data structure with optimizations for high performance computing, and they used libfuzzer (LLVM Project, 2026) to perform differential testing (McKeeman, 1998) against the implementation provided by Rust’s standard library.
 Of the 26 developers interviewed by Höltervennhoff et al.
@@
 The majority of survey respondents also wrote tests for applications that used unsafe code.
-As shown in Table 0, 61% (6.3) of all participants wrote tests most of the time or always.
+As shown in Table 0, 61% (4) of all participants wrote tests most of the time or always.
 This was equally true for the subset of participants who had used Miri; 72% would write test cases at least most of the time.
-However, 56% (6.3) of survey respondents who had used Miri only sometimes used it to run their test cases, if at all.
+However, 56% (4) of survey respondents who had used Miri only sometimes used it to run their test cases, if at all.
 Höltervennhoff et al.
@@
 Additionally, we did not ask if survey respondents regularly ran their test cases.
-Interview participants frequently mentioned that Miri’s slow performance and lack of support for foreign function calls prevented them from using it to find bugs in certain categories of applications.
-Survey respondents who had used Miri encountered similar issues; 62% were also deterred from using it due to its lack of support for foreign functions (43%), slow performance (26%), or lack of support for inline assembly (19%).
+However, interview participants frequently mentioned that Miri’s slow performance and lack of support for foreign function calls prevented them from using it to find bugs in certain categories of applications, and survey respondents who had used Miri encountered similar issues.
+The majority (62%) were deterred from using it due to one or more of the following issues: lack of support for foreign functions (43%), slow performance (26%), or lack of support for inline assembly (19%).
 Developers interviewed by Höltervennhoff et al.
-(Höltervennhoff et al., 2023) also struggled with Miri’s lack of support for key features, but none had mentioned its performance.
+(Höltervennhoff et al., 2023) had also struggled with Miri’s lack of support for key features, but none had mentioned its performance.
 Notably, one of our interview participant had neglected to use Miri since their application relied heavily on foreign function calls: “we’ve assumed that we are a JIT and Miri can’t run JIT code, so therefore it’s useless” (P13).
@@
 Interview participants encountered several issues with tools that generate bindings to foreign functions.
-Participants reported that bindgen added significant runtime overhead during build times and was not as expressive as they desired.
+Participants reported that bindgen (Rust Community, 2026a) added significant runtime overhead during build times and was not as expressive as they desired.
 One participant observed that bindgendid some weird stuff for anonymous unions and structs” (P17) but that this was easy to fix by modifying the tool’s configuration.
@@
 McCormack et al.
-(McCormack et al., 2024) also found several instances of incorrect bindings in Rust libraries.
+(McCormack et al., 2025b) also found several instances of incorrect bindings in Rust libraries.
 Some were intentional—developers neglected to add return types for values they did not use—but others introduced aliasing violations or uninitialized reads.
 Survey respondents who had called foreign functions were somewhat more likely to trust generated bindings.
-As shown in Table 0, only 45% (6.3) of respondents would at least “probably” trust handwritten bindings, compared to 79% (6.3) who would trust generated bindings.
+As shown in Table 0, only 45% (5) of respondents would at least “probably” trust handwritten bindings, compared to 79% (5) who would trust generated bindings.
 However, developers were still more likely to trust handwritten bindings than to distrust them.
@@
 Only 10% of survey respondents used a debugger daily, 29% used one monthly, and 44% used one yearly at most, if at all.
-One interview participant had used both MSVC’s debugger and LLDB, and while both worked, neither matched the quality of the other parts of the Rust toolchain, “where you get all the bells and whistles that you want, and things just work nicely together” (P0).
+One interview participant had used both MSVC’s debugger and LLDB (LLVM Project, 2025), and while both worked, neither matched the quality of the other parts of the Rust toolchain, “where you get all the bells and whistles that you want, and things just work nicely together” (P3).
 Mozilla’s RR debugger (O’Callahan et al., 2017) was useful for one participant to resolve bugs found from fuzzing, but it was generally more common for participants to report using “printf-style” (P7) debugging.
@@
 Few interview participants and only 10 survey respondents used tools that apply formal methods for static and dynamic verification.
-Interview participants mentioned using Kani, Prusti, and Crucible.
+Interview participants mentioned using Kani (VanHattum et al., 2022), Prusti (Astrauskas et al., 2019), and Crucible (Galois, 2025).
 Kani was useful for one participant, who needed to prove that a particular routine would never execute more than twice for any input.
 Another participant who had “tried a bunch of the tools” in this category found that Prusti was the most complete, but they felt that it would be more effective if these tools were “inherently parts of Rust” so that they could “hoist a lot of currently unsafe code into safe code” (P11).
-Of the ten survey respondents who had used “formal methods tools,” five had used Kani, three had used Prusti, three had used Creusot, one had used Flux, and one had used Crucible.
+Of the ten survey respondents who had used “formal methods tools,” five had used Kani, three had used Prusti, three had used Creusot (Denis et al., 2022), one had used Flux (Lehmann et al., 2023), and one had used Crucible.
 We provided a short-response option for this question to include any tools that were missing from our list.
@@
 Survey respondents rarely audited their dependencies’ use of unsafe code.
-More than 80% (6.3) of respondents only sometimes audited their dependencies’ use of unsafe code, if at all.
+More than 80% (4) of respondents only sometimes audited their dependencies’ use of unsafe code, if at all.
 It is possible that a subset of the participants who answered “Never” to this question did not have any dependencies to audit.
@@
 (Höltervennhoff et al., 2023).
-Of this subset, 63% had used cargo-audit, 42% had used cargo-update, 31% had used cargo-deny, 22% had used cargo-geiger, and only 7% had used cargo-vet.
+Of this subset, 63% had used cargo-audit (Rust Community, 2025f), 42% had used cargo-update (Rust Community, 2025e), 31% had used cargo-deny (Rust Community, 2025c), 22% had used cargo-geiger (Rust Community, 2025d), and only 7% had used cargo-vet (Rust Community, 2025b).
 Key Findings - RQ2 (Tooling).
 What tools do Rust developers use when contributing to applications that include unsafe code, and how could tooling be improved?
-Miri appears to be the “de facto” (Gäher et al., 2024) bug finding tool for the Rust ecosystem, since it was used by 61% of all respondents.
+Miri appears to be the de facto bug finding tool for the Rust ecosystem, since it was used by 61% of all respondents.
 However, 62% of the respondents who had used Miri before were deterred from using it again due to its performance and lack of support for key features—predominately, foreign function calls.
@@
 Each cell shows the percentage of participants who identified with both of the motivations for that row and column.
-The cell at the bottom right hand corner shows the percentage who selected all three motivations.
+All three motivations were selected by 7% of participants.
 Necessity Performance Ergonomics
@@
 Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”.
-The number of respondents in the sample is included in parenthesis on the righthand side.
-{tblr}
-width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
-8a When you choose to use unsafe because it performs faster or is more space efficient, how often do you measure the difference? Used unsafe for performance
-[Uncaptioned image]46%36%(76)
-[…]
+The number of respondents in the sample is included in parentheses on the righthand side.
+ID Question Sample Distribution
+7a When you choose to use unsafe because it performs faster or is more space efficient, how often do you measure the difference? Used unsafe for performance 46%36%(76)
+Always    Most of the time    About half the time    Sometimes    Never
 Performance
@@
 Survey respondents who reported using unsafe code to improve performance did not have a strong tendency toward either large-scale or small-scale use.
-We found that 26% only pursued “small-scale-optimizations in existing applications,” 20% only built new, “large-scale components,” and 40% used unsafe code for performance in either situation.
+We found that 26% only pursued “small-scale-optimizations in existing applications,” 20% only built new, “large-scale components,” and 42% used unsafe code for performance in either situation.
 Participants did not consistently measure the performance impact of unsafe code.
 In some situations, they relied on their intuition to determine which design decisions would be best: “I’ll totally admit that…I think this is going to be a hot thing, so I’m going to kind of prematurely optimize” (P11).
-Survey respondents who used unsafe to increase performance were also inconsistent about measuring its impact.
+Survey respondents who used unsafe code to increase performance were also inconsistent about measuring its impact.
 As shown in Table 0, 46% of respondents reported measuring performance at least most of the time, while 36% only sometimes measured performance, if at all.
@@
 Another participant used transmutation when working with a Rust encapsulation of a C font library.
-This library exposed C heap allocations as Rust s with short lifetimes—even though these objects would remain valid on the heap for the duration of the program.
+This library exposed C heap allocations to Rust as references with short lifetimes—even though these objects would remain valid on the heap for the duration of the program.
 This participant found it was easier to transmute the Rust objects to have the 'static lifetime, which is indefinite, instead of adapting to the restrictions of the encapsulation.
@@
 The percentage on the left includes “Always” and “Most of the time”, while the one the right includes “Sometimes” and “Never”.
-The number of respondents in the sample is included in parenthesis on the righthand side.
-{tblr}
-width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
-9a When you feel that you have no clear alternative other than using unsafe, how often are you certain that it would be completely impossible to accomplish this task using a safe design pattern?
-Had no alternative to unsafe code
-[Uncaptioned image]65%17%(123)
-[…]
+The number of respondents in the sample is included in parentheses on the righthand side.
+ID Question Sample Distribution
+8a When you feel that you have no clear alternative other than using unsafe, how often are you certain that it would be completely impossible to accomplish this task using a safe design pattern?
+Had no alternative to unsafe code 65%17%(123)
+Always    Most of the time    About half the time    Unsure    Sometimes    Never
 Necessity
@@
 For example, one participant found that it was necessary to implement Send and Sync in a single-threaded application to maintain compatibility with an existing API.
-The perceived that the implementation was sound by construction.
-However, each of these traits are declared as unsafe, so they can only be implemented in an unsafe context.
-Participants who contributed to just-in-time (JIT) compilers and operating systems found that access memory through raw pointers was unavoidable.
+They perceived that the implementation was sound by construction.
+However, both of these traits are unsafe and only be implemented in an unsafe context.
+Participants who contributed to just-in-time (JIT) compilers and operating systems found it difficult to avoid accessing memory through raw pointers.
 One participant found that this was necessary to implement an unwinding table: “I have to jump to that address and I really can’t do anything to verify it…” (P12).
@@
 Each of these motivations can influence a design decision simultaneously.
-We expect that interventions that consider the intersection of multiple motivations will be more effective.
-For example, profiling tools could leverage information on the distribution of unsafe code to assist developers in making a compromise between safety and performance.
 5.5. RQ4 - Encapsulation
@@
 Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”.
-The number of respondents in each sample is included in parenthesis on the righthand side.
-{tblr}
-width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
-10a When you use an unsafe API, how often do you look for documentation to ensure that you meet all of its requirements for safety and correctness? Used unsafe APIs
-[Uncaptioned image]85%8%(152)
-10b When you expose a safe API for unsafe code, how often do you include runtime checks to ensure that its requirements for correctness and safety are met? Exposed a safe API for unsafe code
-[Uncaptioned image]58%29%(141)
-[…]
-[Uncaptioned image]90%3%(68)
-10d When you expose an unsafe API to other users, and it is their responsibility to ensure that certain requirements are met, how often do you include runtime checks for these requirements?
-[Uncaptioned image]18%71%(68)
-[…]
-[Uncaptioned image]6%88%(160)
-Regularly wrote unsafe code
-[Uncaptioned image]7%85%(85)
-[…]
-[Uncaptioned image]82%10%(160)
-Used unsafe APIs
-[Uncaptioned image]83%10%(152)
-[…]
+The number of respondents in each sample is included in parentheses on the righthand side.
+ID Question Sample Distribution
+9a When you use an unsafe API, how often do you look for documentation to ensure that you meet all of its requirements for safety and correctness? Used unsafe APIs 85%8%(152)
+9b When you expose a safe API for unsafe code, how often do you include runtime checks to ensure that its requirements for correctness and safety are met?
+Exposed a safe API for unsafe code 58%29%(141)
+9c When you expose an unsafe API to users, and it is their responsibility to ensure that certain requirements are met, how often do you document these requirements? Exposed an unsafe API 90%3%(68)
+9d When you expose an unsafe API to other users, and it is their responsibility to ensure that certain requirements are met, how often do you include runtime checks for these requirements? 18%71%(68)
+9e How often do you refactor Rust applications to remove unsafe code? All 6%88%(160)
+Regularly wrote unsafe code 7%85%(85)
+9f How often do you choose to avoid using an unsafe API when a safe alternative exists? All 82%10%(160)
+Used unsafe APIs 83%10%(152)
+Always    Most of the time    About half the time    Sometimes    Never
 Invariants
@@
 Some participants perceived hardware-level correctness properties as “orthogonal to Rust” (P1), since they were related to the properties of the language.
-However, other participants who developed embedded applications relied on third-party libraries to encode hardware requirements into the type system, so that their program would fail to compile if a device was misconfigured.
-Developers leveraged Rust’s type system to uphold invariants that they perceived as necessary for soundness.
+However, other participants who developed embedded applications relied on third-party libraries to encode hardware requirements into the type system so that their programs would fail to compile if a device had been misconfigured.
+Developers leveraged Rust’s type system to uphold invariants that they believed were necessary for soundness.
 In operating systems, these typically began as operation-specific refinement properties on primitive types, but they quickly built up into system-level invariants; “it gets to a high-level really quickly” (P3).
@@
 This method was popularized for Rust by Yanovski et al.’s GhostCell type (Yanovski et al., 2021).
-Other participants relied on ad-hoc reasoning to ensure that their programs were correct and free of undefined behavior.
+Other participants relied on ad hoc reasoning to ensure that their programs were correct and free of undefined behavior.
 Some appealed to their prior experience: “[in] C++, you do this all the time” (P15).
@@
 However, participants reported that unsafe code was prevalent in JIT compilers and operating systems.
-Domain-specific, non-local reasoning was necessary to understand any arbitrary unsafe code snippet: “you have to know the innards of the JIT-compiled function…so it’s just not encapsulated” (P0).
+Domain-specific, non-local reasoning was necessary to understand any arbitrary unsafe snippet: “you have to know the innards of the JIT-compiled function…so it’s just not encapsulated” (P11).
 Though unsafe code may have been isolated, it is unclear if participants used it minimally.
-As shown in Table 0, more than 80% (6.5) of survey respondents only sometimes refactored their code to remove unsafe features, if at all.
+As shown in Table 0, more than 80% (9) of survey respondents only sometimes refactored their code to remove unsafe features, if at all.
 This distribution was similar for the subset of respondents who had regularly written unsafe code.
 Safe Interfaces
-Most interview participants and 88% of survey respondents reported that they would attempt to encapsulate unsafe code beneath safe APIs.
+Most interview participants and 88% of survey respondents reported that they had exposed a safe API for unsafe code.
 This was also common for Höltervennhoff et al.
@@
 Our interview participants would create encapsulations where they perceived that the requirements for their unsafe code were satisfied by either Rust’s type system or what they reasoned to be local invariants.
-However 43% of our survey respondents reported exposing unsafe APIs, while only 6 of Höltervenhoff et al.
-(Höltervennhoff et al., 2023)’s 26 participants did so.
-Interview participants cited the same motivations for exposing unsafe APIs as they did for using unsafe code in any capacity.
+However 43% of our survey respondents reported exposing unsafe APIs, relative to six of Höltervenhoff et al.
+(Höltervennhoff et al., 2023)’s 26 interview participants.
+Our participants cited the same motivations for exposing unsafe APIs as they did for using unsafe code in any capacity.
 One participant exposed unsafe versions of safe API endpoints which had the same behavior, but they did not include runtime checks.
@@
 Regardless of motivation, 61% of respondents claimed that they would only expose unsafe APIs when they had requirements beyond what Rust’s type system could verify, and they typically documented these requirements.
-Results for Question 6.5 indicate that 90% of respondents who exposed unsafe APIs would document their safety requirements most of the time, if not always.
-However, respondents did not typically add any additional safeguards to prevent unsafe APIs from being used incorrectly, since 71% (6.5) only sometimes inserted run-time checks within these APIs, if at all.
-This was more common when exposing safe APIs; 58% (6.5) inserted checks at least most of the time when exposing a safe API for unsafe code.
-When using unsafe APIs, 85% (6.5) of participants would usually look for documentation to inform their design decisions.
+Results for Question 9 indicate that 90% of respondents who exposed unsafe APIs would document their safety requirements most of the time, if not always.
+However, respondents did not typically add any additional safeguards to prevent unsafe APIs from being used incorrectly, since 71% (9) only sometimes inserted run-time checks within these APIs, if at all.
+This was more common when exposing safe APIs; 58% (9) inserted checks at least most of the time when exposing a safe API for unsafe code.
+When using unsafe APIs, 85% (9) of participants would usually look for documentation to inform their design decisions.
 Interview participants perceived that the Rust community has a strong preference for safety, which affects tool design and development practices.
@@
 An overwhelming majority of survey respondents also preferred safe APIs.
-More than 80% (6.5) of respondents claimed that they would choose a safe API over an unsafe alternative most of the time, if not always—regardless of whether or not they had ever used an unsafe API before.
+More than 80% (9) of respondents claimed that they would choose a safe API over an unsafe alternative most of the time, if not always—regardless of whether or not they had ever used an unsafe API before.
 Table 10.
@@
 Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”.
-The number of respondents in each sample is included in parenthesis on the righthand side.
-{tblr}
-width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
-11a When you expose a safe API for unsafe code, how often are all of its requirements for correctness and safety satisfied by the properties of Rust’s type system? Exposed a safe API for unsafe code
-[Uncaptioned image]62%18%(141)
-11b How often is the Rust community’s guidance and documentation adequate for you to know how to use unsafe correctly? All
-[Uncaptioned image]70%16%(160)
-[…]
+The number of respondents in each sample is included in parentheses on the righthand side.
+ID Question Sample Distribution
+10a When you expose a safe API for unsafe code, how often are all of its requirements for correctness and safety satisfied by the properties of Rust’s type system?
+Exposed a safe API for unsafe code 62%18%(141)
+10b How often is the Rust community’s guidance and documentation adequate for you to know how to use unsafe correctly? All 70%16%(160)
+Always    Most of the time    About half the time    Unsure    Sometimes    Never
 Uncertainty
@@
 When describing their uncertainty about encapsulation, our participants compared the “theory” of safety to the “practice” of how an API would be used.
-When these concepts didn’t align, errors occurred.
+When these concepts were misaligned, errors occurred.
 One participant observed this mismatch with a Rust crate that encapsulated a foreign library.
@@
 Another participant struggled with discrepancies between the interface and the implementation of Rust’s Layout API, which describes the size and alignment of an allocation.
-The design of Layout was changed to “tweak the safety requirements” (P0), which regretfully caused previously correct code to exhibit undefined behavior.
+The design of Layout was changed to “tweak the safety requirements” (P14), which caused code that was previously correct to exhibit undefined behavior.
 Interview participants also had a collective sense of uncertainty about Rust’s semantics.
@@
 Another participant wished to be able to take a reference to the first field of a product type and use it to create valid references to adjacent fields using integer offsets.
-This was considered undefined behavior under Rust’s Stacked Borrows (Jung et al., 2019) aliasing model, but it is accepted under the newer Tree Borrows model (Villani et al., 2023).
+This was considered undefined behavior under Rust’s Stacked Borrows (Jung et al., 2019) aliasing model, but it is accepted under the newer Tree Borrows model (Villani et al., 2025).
 Additionally, multiple participants were uncertain about how to handle the case when a program panics while a value is being dropped.
 Survey respondents were somewhat more certain about their decisions.
-As shown in Table 0, more than 60% (6.5) of respondents who had exposed a safe API for unsafe code were certain at least most of the time, if not always, that its safety properties were satisfied by Rust’s type system.
-Additionally, 70% (6.5) of respondents found that the Rust community’s documentation and guidance was adequate for them to know how to use unsafe code correctly at least most of the time.
+As shown in Table 0, more than 60% (10) of respondents who had exposed a safe API for unsafe code were certain at least most of the time, if not always, that its safety properties were satisfied by Rust’s type system.
+Additionally, 70% (10) of respondents found that the Rust community’s documentation and guidance was adequate for them to know how to use unsafe code correctly at least most of the time.
 However, only 6% felt that it was always adequate, suggesting that improved documentation would be helpful for the overwhelming majority of respondents.
-Similar to the questions in Tables 6.3 and  6.4, we added “Unsure” as a response option for the questions shown in Table 0 since they rely on developers’ subjective notions of certainty and adequacy.
+Similar to the questions in Tables 5 and  8, we added “Unsure” as a response option for the questions shown in Table 0, since they rely on developers’ subjective notions of certainty and adequacy.
 Key Findings - RQ4 (Encapsulation).
@@
 Interview participants derived requirements for unsafe code from external specifications and the inherent guarantees of the borrow checker.
-However, they also relied on ad-hoc reasoning and tacit knowledge.
+However, they also relied on ad hoc reasoning and tacit knowledge.
 Many were uncertain if their encapsulations were sound in all situations.
@@
 Survey respondents who exposed unsafe APIs documented their safety properties, and they relied on documentation to use unsafe APIs correctly.
-We expect that this documentation could be leveraged by tool designers to create formal specifications for these APIs.
-When participants chose to expose unsafe APIs, we found that they were motivated by performance, ergonomics, and necessity to the same extent as for any arbitrary use of unsafe code.
+When our participants chose to expose unsafe APIs, they were motivated by performance, ergonomics, and necessity to the same extent as for any arbitrary use of unsafe code.
 Respondents rarely refactored unsafe code, even if they used it regularly.
-Static analysis tools could encourage refactoring by suggesting safe, idiomatic replacements, but it is also possible the the best practices for writing unsafe code make frequently refactoring unnecessary.
+6. Related Work
+Here, we discuss relevant results from other studies of unsafe Rust code, comparing our results to quantitative findings where relevant.
+Most prior studies have examined unsafe code as it is written, either by surveying Rust’s package ecosystem (Astrauskas et al., 2020; Evans et al., 2020; Ozdemir, 2016) and standard library (Cui et al., 2024), or by examining vulnerabilities caused by unsafe code (Qin et al., 2020; Xu et al., 2021; Zheng et al., 2023; McCormack et al., 2025b; Zhang et al., 2023).
+Other studies have collected blogs, articles, and forum posts to understand the implications of unsafe code for the Rust community.
+Zeng and Crichton (Zeng and Crichton, 2018) examined a set of articles and comments posted on Hacker News and the Rust subreddit to identify potential challenges to Rust’s adoption.
+Zhu et al.
+(Zhu et al., 2022) examined posts on Stack Overflow to determine if developers frequently used unsafe code to resolve compilation errors.
+These studies provide meaningful contextual information that we can use to cross-validate our findings.
+However, none can fully answer our research questions.
+Far fewer studies have used surveys or interviews to directly engage with developers who use unsafe code.
+As an extension to their survey of Rust libraries, Evans et al.
+(Evans et al., 2020) distributed a short survey on the Rust subreddit.
+They asked developers why they use unsafe code, which operations they use, and what steps they take to ensure that their code is correct; it received 20 responses.
+Zeng and Crichton (Zeng and Crichton, 2018) also interviewed three Rust contributors to provide additional context in their analysis.
+Fulton et al.
+(Fulton et al., 2021) interviewed 16 developers who had advocated for using Rust in industry, followed by a survey that received 178 responses from many of the same communities as ours.
+Although they used a similar methodology, Fulton et al.
+(Fulton et al., 2021) mainly focused on evaluating the benefits and drawbacks of Rust as a general-purpose systems programming language instead of examining Rust’s unsafe features.
+These studies provide evidence that can partially answer a subset of our research questions, vut most do not share our direct focus on unsafe code or do not investigate challenges specific to tool use and interoperation.
+Höltervennhoff et al.
+(Höltervennhoff et al., 2023) are the only other group that we are aware of that has used qualitative methods to study how developers use unsafe code.
+Their study was published concurrent to our investigation; after we had completed our interviews but before distributing our survey.
+They interviewed 26 Rust developers who had experience using unsafe code.
+Except for three who piloted the interview protocol, all of their participants had contributed unsafe code to open source projects included in the unofficial “Awesome Rust,” list (Community, 2026a) of applications.
+We cover similar topics, but our research questions reflect different, complementary goals.
+They focused on understanding how developers reason about the security implications of using unsafe code, while we focused specifically on interoperation and tool use.
+We provide additional details on challenges associated with tool use and interoperation that are missing from their work, and we demonstrate that their results can generalize to a broader population of developers.
+Here, we describe how our results compare with these prior studies for each of our research questions.
+RQ1 - Interoperation
+Interoperation has consistently been one of the most common use cases for unsafe code.
+In 2020, Astrauskas et al.
+(Astrauskas et al., 2020) found that 44.6% of all publicly-exported unsafe functions in the Rust ecosystem were static bindings to foreign functions.
+Other surveys of the Rust ecosystem have had similar findings; Ozdemir (Evans et al., 2020) found that 30% of unsafe blocks were entirely dedicated to foreign function calls, and Evans et al.
+(Evans et al., 2020) found that 22.5% of unsafe function calls were to foreign functions.
+Zhang et al.
+(Zhang et al., 2023) manually analyzed 5,946 unsafe blocks from 140 prominent libraries to determine which unsafe operations were necessary and which could have been avoided.
+The majority of all raw pointer dereferences and unsafe function calls were marked as necessary to interoperate with foreign libraries.
+Studies of developers have reached similar conclusions; 70% of Fulton et al.
+(Fulton et al., 2021)’s survey respondents and all but one of Höltervennhoff et al.
+(Höltervennhoff et al., 2023)’s interview participants had called foreign functions from Rust.
+Interoperation was also common among our participants; 70% of survey respondents had called foreign functions.
+However, few studies have identified specific challenges associated with interoperation.
+Cui et al.
+(Cui et al., 2024) explicitly excluded foreign functions from their investigation of Rust’s safety properties.
+Xu et al.
+(Xu et al., 2021) found 12 examples of bugs and vulnerabilities that involved foreign function calls.
+They were related to application-specific invariants, data races, invalid alignment, and inconsistent layouts.
+However, the authors provided few details on the nature of these bugs, indicating that they were “straightforward” and not specific to Rust.
+Fulton et al.
+(Fulton et al., 2021)’s participants had varying experiences with interoperation, but C++ was particularly difficult due to what one participant described as “rampant aliasing.” Höltervennhoff et al.
+(Höltervennhoff et al., 2023) indicated that some participants struggled with foreign function calls that involved “memory management,” but that pure functions were easier to reason about.
+Our interview participants had similar experiences.
+Calling pure functions was typically straightforward, as long as the bindings were correct and ABI-compliant.
+When participants struggled to reason about foreign function calls, it was usually because these APIs had aliasing or concurrency patterns that conflicted with Rust’s idioms.
+Even fewer studies have identified bugs that meaningfully involved Rust’s aliasing model.
+In 2023, a set of anonymous authors classified 22 bugs from five popular Rust encapsulations of foreign libraries (Anonymous, 2023).
+Most of the errors they investigated involved incorrect handling of exceptions and heap objects that were improperly shared across foreign boundaries.
+They identified two bugs related to Rust’s aliasing restrictions, but they did not describe these bugs in terms of Stacked or Tree Borrows.
+McCormack et al.
+(McCormack et al., 2025b) extended Miri to interoperate with an LLVM interpreter and used this hybrid tool to conduct a large-scale evaluation of multilanguage libraries.
+In Rust, packages and libraries are referred to as “crates” and are published on crates.io.
+They found 46 bugs from 37 crates, and one of these crates was maintained by the Rust Project.
+This crate had an aliasing violation that occurred after a foreign API had copied a pointer derived from a “unique” mutable reference.
+This fits a pattern described by one of our interview participants, who was concerned that foreign APIs would “retain a pointer” from Rust and “keep operating on it later” (P11).
+Yu et al.
+(Yu et al., 2025) found additional aliasing violations with CapsLock, an instrumentation tool that uses custom RISC-V hardware capabilities.
+Their approach performs nearly twice as fast as Miri and has support for inline assembly, but its aliasing model is weaker than both Stacked and Tree Borrows, and the capabilities that it relies on have not been implemented in physical hardware yet.
+They evaluated their design on a variety of test cases that called foreign functions and found similar categories of aliasing violations as McCormack et al (McCormack et al., 2025b).
+In particular, certain test cases were “mutating an owned object without borrowing the owner” (Yu et al., 2025).
+RQ2 - Tooling
+Rust is a relatively new programming language, so most research efforts have focused on developing new tools and not evaluating existing ones.
+Zeng and Crichton (Zeng and Crichton, 2018) hypothesized that a lack of promotion for tooling could prevent Rust from being widely adopted as a systems language.
+However, their work was published prior to the creation of most Rust-specific bug-finding tools.
+Lack of promotion could explain why few of our participants had used formal methods tools.
+However, several of the tools listed in our survey, such as Verus (Lattuada et al., 2023), had recently been introduced, and few had support for unsafe code.
+Today, a variety of static verification tools for Rust are being used in industry, and we expect that studying their adoption would be a useful direction for future work.
+Höltervennhoff et al.
+(Höltervennhoff et al., 2023) are the only other group that we are aware of to examine how developers use tooling to assist with writing unsafe code.
+Certain development tools were somewhat underused among their interview population; only seven of their 26 participants had mentioned using Miri.
+Since 61% of our survey respondents had used Miri, this does not seem to indicate a broader trend within the Rust community.
+Likewise, few of their interview participants used automated auditing tools like cargo-audit, while 51% of our survey participants had used one or more of these tools.
+However, our respondents only needed to have used these tools once, while Höltervennhoff et al.
+(Höltervennhoff et al., 2023) indicated that participants who audited their dependencies did so “regularly.” Höltervennhoff et al.
+(Höltervennhoff et al., 2023) also indicated that the majority of their participants had used Clippy, and that some had mixed experiences.
+Although many of our interview participants had also used Clippy, our investigation was primarily concerned with bug-finding tools, and not linting tools, so we did not highlight it in our analysis or survey.
+We also cover debuggers, static verifiers, and tools that generate bindings to foreign functions, which were not mentioned in their study.
+RQ3 - Motivations
+Multiple studies have identified different use cases for unsafe code.
+Qin et al.
+(Qin et al., 2020) collected random samples of operations within unsafe contexts from Rust’s standard library and 10 popular third-party libraries.
+They identified three use cases for these operations; 42% were related to interoperation, 22% were to improve performance, and 14% involved sharing memory between threads.
+Astrauskas et al.
+(Astrauskas et al., 2020) used “anecdotal examples” and snippets of source code to create a taxonomy of six purposes for using unsafe code.
+These included implementing data structures with complex sharing, overcoming the incompleteness of Rust’s type system, emphasizing contracts and invariants, calling foreign functions, using intrinsics or inline assembly, and improving performance.
+The most frequently used unsafe feature was calling unsafe functions.
+Fulton et al.
+(Fulton et al., 2021) found that developers most frequently used unsafe code to call foreign functions, improve performance, and interact directly with hardware or the operating system.
+Unlike these studies, we are concerned with “motivations,” which we define as feature-agnostic reasons for using unsafe code.
+This excludes several of the feature-specific topics mentioned in prior work, such as intrinsics (Astrauskas et al., 2020) or sharing memory between threads (Astrauskas et al., 2020).
+We cannot reliably assign motivations to specific features without knowing more about the decision-making process behind each use case.
+Our perspective is similar to Höltervennhoff et al.
+(Höltervennhoff et al., 2023), who identified necessity, performance, and efficiency as “motivations.” It is unclear whether this distinction was also an explicit consideration for their study.
+Zhang et al.
+(Zhang et al., 2023) were able to classify unsafe operations as “necessary” by applying a strict definition: it had to be impossible to implement a safe alternative.
+In their evaluation of Rust libraries, they identified two reasons why developers chose to use unsafe code when it was not strictly necessary: to improve performance, and because they were unaware of a safer alternative.
+They did not identify whether developers were aware of a safe alternative but had chosen to use unsafe anyway, since it was easier.
+Their strict criteria also excludes certain context-specific perspectives on what it means for unsafe code to be necessary.
+For example, they classified all instances of UnsafeCell as unnecessary, since Rust’s safe RefCell can be used instead.
+However, it may be necessary to avoid the run-time cost of RefCell in performance-critical applications.
+Zhang et al.
+(Zhang et al., 2023) account for this by evaluating the performance impact of several safe alternatives to unsafe operations, but we expect that developers’ subjective experiences will also be helpful in understanding what motivates them to use unsafe code.
+RQ4 - Encapsulation
+Unsafe code is prevalent throughout the Rust ecosystem, but developers use it minimally and hide it under safe interfaces.
+In July of 2016, Ozdemir (Ozdemir, 2016) found that 29% of all published crates used unsafe features, but most unsafe blocks were relatively small and tightly-scoped to include only unsafe operations.
+Evans et al.
+(Evans et al., 2020) observed similar patterns in September of 2018.
+The number of published crates had increased by more than ten times, but 29% still directly used unsafe code.
+Although 38% of crates did not directly use unsafe code, they either directly or indirectly depended on third-party crates that did use these features.
+In January of 2020, Astrauskas et al (Astrauskas et al., 2020) found that 23.6% of third-party crates used unsafe features.
+Most instances of unsafe code were relatively straightforward and hidden beneath safe interfaces.
+The majority of unsafe functions called by libraries were standard (e.g.
+not closures, function pointers, or trait functions) and defined within the same crate or provided by Rust’s standard library.
+In particular, 34.7% of all crates had declared unsafe APIs, but they were not visible to users.
+This aligns with our participants’ experiences; 88% claimed that they would attempt to encapsulate unsafe operations beneath safe APIs, and 61% expected that it would be at least somewhat easy for another developer to understand a random unsafe block or function from their code.
+Developers typically mark functions with the unsafe keyword to indicate that users need to uphold a particular safety property.
+The majority (61%) of the survey respondents who had exposed unsafe APIs only did so when they required an invariant beyond what Rust’s type system could provide.
+However, Astrauskas et al.
+(Astrauskas et al., 2020) found that 36.1% of unsafe functions had completely safe implementations.
+On further investigation, many of these functions had been automatically generated or annotated as unsafe for “legacy reasons.” This could provide an alternative explanation beyond necessity or ease-of-use for why the remaining 39% of our respondents were motivated to expose unsafe APIs.
+While unsafe functions are relatively common, developers rarely use unsafe traits.
+Evans et al.
+(Evans et al., 2020) found that only slightly more than 1% of all crates declared an unsafe trait, and only 6% implemented an unsafe trait.
+Astrauskas et al.
+(Astrauskas et al., 2020) had similar findings; only 2.5% of all traits were unsafe, and 40.4% of these declarations originated from five crates.
+Likewise, Zhang et al.
+(Zhang et al., 2023) only identified 11 instances of unsafe trait implementations.
+Astrauskas et al.
+(Astrauskas et al., 2020) suggested that this lack of use could be caused by a lack of guidance on when a trait should be labeled unsafe.
+One of our interview participants indicated that certain developers may have misconceptions on the role of unsafe traits.
+As mentioned earlier in Section 0, one participant who maintained a library that exposed an unsafe trait found that users who had implemented it incorrectly would blame the library, even though “in theory, the burden of proof is on them” (P7).
+Both the Rust Reference (Rust Community, 2023a) and the Rustonomicon (Rust Community, 2026b) have brief sections on unsafe traits, but additional resources could be helpful.
+The safety properties for unsafe features are varied and complex.
+Cui et al.
+(Cui et al., 2024) manually reviewed 416 unsafe APIs in Rust’s standard library and identified 19 categories of preconditions and postconditions.
+Partly for this reason, discrepancies between the signatures of safe APIs and their interior unsafe implementations have been a significant source of bugs and vulnerabilities for the Rust ecosystem.
+Qin et al.
+(Qin et al., 2020) examined 170 bugs and vulnerabilities from ten high-profile Rust libraries.
+Memory safety errors typically occurred in safe contexts but were caused elsewhere by errors in interior unsafe implementations.
+Qin et al.
+(Qin et al., 2020) also identified 19 examples of incorrect encapsulations of unsafe code from both the third-party libraries in their sample and within Rust’s standard library.
+Xu et al.
+(Xu et al., 2021) exclusively studied registered vulnerabilities and identified similar errors in safe encapsulations.
+Zheng et al.
+(Zheng et al., 2023) conducted the broadest empirical study of security issues in the Rust ecosystem to date.
+They examined 433 unique vulnerabilities that were discovered between 2014 and 2022, identifying 300 affected repositories.
+Vulnerable components were significantly more likely to include unsafe blocks or functions.
+The overwhelming majority of all vulnerabilities affected safe functions.
+For this reason, it is somewhat concerning that our participants were often uncertain about the correctness of their encapsulations.
+Höltervennhoff et al.
+(Höltervennhoff et al., 2023) had similar findings; most of their participants relied on common sense and test coverage to evaluate the security of their code, and few had contributed to projects with formal policies for security or code review.
+However, many of their participants indicated that they would review unsafe code more carefully to ensure that it upheld all necessary safety invariants.
+Few participants had ever encountered security issues caused by unsafe code, and none of these issues were critical.
+Many of our interview participants had encountered memory or thread-safety bugs in unsafe implementations, but few described the security implications of these issues.
+Security was not a central focus of our study, so it is possible that our participants had reflected on the security implications of their decisions but neglected to mention this during interviews.
 7. Discussion
-The majority of our participants attempted to follow several best practices when using unsafe code, such as testing (6.3), writing documentation (6.5), and using safe interfaces whenever possible (6.5).
-As shown in Table 0, more than 70% of respondents were motivated to use unsafe code when they perceived that there was no safe alternative.
-However, only 14% (6.4) of these respondents were always certain that it would be impossible to implement a safe alternative.
-Likewise, 62% (6.5) of respondents who exposed safe APIs for unsafe code were certain “most of the time” that their interfaces were sound, but only 23% were always certain.
-We cannot rely on developers’ perceptions to accurately reflect the soundness of their code, but this uncertainty suggests that there is still a gap between Rust’s promise of static safety and the reality of how developers use unsafe features.
-Multiple factors contribute to this uncertainty.
-Foreign function calls presented unique challenges due to the differences between Rust and other languages.
-Rust’s aliasing model is still evolving, so certain features do not yet have an agreed-upon semantics in every possible situation (Jung, 2020).
-Developers who implement design patterns that sit at the cutting-edge of Rust’s evolving aliasing model will inevitably be uncertain if their implementations are sound.
-However, all developers who write unsafe code will need a reliable way to determine if their implementations follow the rules of whichever aliasing model becomes canonical.
-Miri is not entirely adequate for this task, since 61% of the respondents who had used it were deterred by its slow performance and lack of support for key features.
-However, even a fast, feature-complete version of Miri will not be enough to ensure that developers can always be certain that their implementations are sound, since dynamic analysis is inherently incomplete.
-To address the challenges faced by our participants, the next generation of Rust-specific development tools should be capable of providing guarantees of soundness and should support multi-language applications.
-Static analysis will be particularly helpful for encapsulating foreign functions and guiding developers toward refactoring unsafe code.
-However, any new approaches to tooling should consider the role of performance and ergonomics as motivations for using unsafe features.
-Multi-language Support
-Participants who called foreign functions would pass pointers to heap objects in both directions between Rust and other languages.
-Many used Rust’s memory container types as allocators, and few actively avoided casting foreign pointers into Rust’s safe references.
-Each of these actions can lead to undefined or undesirable behavior, since foreign code can break Rust’s aliasing rules.
-However, no production-ready tools are capable of detecting aliasing violations in this context.
+Our goal in studying interoperation was to describe how developers navigate the differences between Rust and other languages (RQ1) and to determine how tooling could be improved to make this process easier (RQ2).
+Our interview participants found that interoperation was more or less difficult depending on how the characteristics of foreign APIs matched the inherent expectations of Rust’s aliasing model.
+When documentation was missing or certain design idioms conflicted with Rust’s aliasing rules, participants had no choice but to minimize their interaction with foreign function calls to reduce the perceived risk of triggering undefined behavior.
+At the time this survey was conducted, Miri was the only tool capable of finding violations of Rust’s aliasing model—it was both the most popular tool and the most lamented.
+Participants indicated that they had been deterred from using it due to its lack of support for foreign function calls and inline assembly, or its slow performance.
+These challenges are not localized at foreign function boundaries.
+Developers need to reconcile the semantics of foreign APIs with safe, idiomatic Rust encapsulations, which often requires implementing a layer of interior-unsafe operations.
+Developers’ motivations for using these operations (RQ3) and their ability to reason about encapsulating unsafe code (RQ4) are inherently relevant to this process.
+Our population prioritized exposing safe encapsulations whenever possible.
+Necessity was the most predominant motivation for exposing unsafe APIs, and the overwhelming majority (9) of participants indicated that they would always document any necessary preconditions and postconditions.
+However, factors that are orthogonal to correctness, like ergonomics and performance, influenced how participants used unsafe code within safe encapsulations.
+For example, one interview participant indicated that it had been easier to encapsulate a foreign API by transmuting a pointer to a long-lived allocation into a reference with a 'static lifetime.
+The alternative would have been to derive a context-specific lifetime, and another participant indicated that this pattern could make an API more difficult to use.
+These motivations were less common but still significant—51% of participants who exposed unsafe APIs cited performance as a motivation.
+We expect that Rust developers will continue to prioritize exposing safe APIs, at least in part due to community incentives.
+More than 80% (9) of participants who had used an unsafe APIs before would prefer a safe alternative at least most of the time, if not always.
+As one interview participant put it: “everyone wants a safe interface… no one really wants to go use the unsafe heavy ones” (P9).
+This may explain why interview participants were willing to rely on ad hoc, application-specific reasoning to justify their design decisions in the face of domain-specific challenges and limitations in tooling.
+As a result, developers are left uncertain about their design decisions.
+Only 14% (8) of survey respondents who used unsafe code by necessity were always certain that it would be impossible to implement a safe alternative.
+Likewise, 62% (10) of respondents who exposed safe APIs for unsafe code were certain “most of the time” that their interfaces were sound, but only 23% were always certain.
+Participants suggested several potential solutions to this uncertainty.
+Developers who contributed to applications with a significant body of specialized unsafe features—mainly multilanguage applications, JIT compilers, and operating systems—would benefit from new dynamic analysis tools for finding Rust-specific aliasing bugs at-scale.
+Static analysis could also be helpful to generate and validate bindings and encapsulations for unsafe functions.
+However, Rust is also constantly evolving, and interview participants identified several topics that they were uncertain about due to Rust not having a formal, comprehensive specification.
+For this reason, we expect that our population would benefit from more comprehensive documentation on certain unstable language features.
+7.1. Dynamic Tooling
+Our participants also indicated that they needed new tools for finding Rust-specific forms of undefined behavior: “a version of Miri that could go across that FFI boundary” (P18), or “…a unique pointer sanitizer” (P1).
+Miri has experimental support for executing foreign functions from shared libraries.
+However, it is still several orders of magnitude slower than native execution, prohibiting it from being useful in the types of large-scale C and C++ applications where Rust is increasingly being adopted.
+We see two potential routes for implementing new tools in this category.
+The first would be to insert run-time checks for aliasing violations during compilation.
+This would require lowering a subset of Rust’s type information from the MIR level down into the intermediate representations used by these platforms, so that the instrumentation pass could distinguish between references and raw pointers.
+Yu et al.
+(Yu et al., 2025) have had success with this method in CapsLock, but their approach uses an aliasing model that is strictly weaker than Stacked and Tree Borrows.
+In ongoing work, we are creating BorrowSanitizer (McCormack et al., 2025a), which is implemented without hardware acceleration, allowing us to have parity with Miri’s aliasing models.
+Our approach uses many of the same APIs as other popular LLVM-based tools.
+We expect that compile-time instrumentation tools will provide the combination of performance and compatibility that is necessary to support large-scale, multilanguage applications.
+However, participants who contributed to JIT compilers would also benefit from dynamic binary instrumentation.
+Tools in this category, such as Valgrind (Nethercote and Seward, 2007), can encode and instrument native assembly on-the-fly, which would be necessary to support JIT-compiled assembly.
+In 2023, the Krabcake project (Klock and Garza, 2023) proposed extending Valgrind to be capable of finding Stacked Borrows violations.
+Several participants had mentioned Krabcake and indicated that it would be a natural extension to their workflows, since they had already used Valgrind with Rust programs.
+It was also the second most popular dynamic bug-finding tool behind Miri; it was used by 44% of survey respondents.
+However, Krabcake is no longer actively maintained, and it is not yet capable of finding bugs in real-world programs.
+The Rust community should consider reactivating this project.
+7.2. Static Tooling
+Participants also indicated that static analysis tools could be useful for finding undefined behavior across foreign function boundaries.
+Several analyses could be implemented as extensions to binding generation tools at varying levels of sophistication.
+One participant wanted a tool that could tell “whether your FFI conforms to your expectations…Do all my symbols match?
+Do all the types signatures match?” (P5).
 McCormack et al.
-(McCormack et al., 2024) were able to extend Miri with support for calling foreign functions, but the limitations of interpretation will prevent their approach from being effective beyond the context of their investigation.
-A production-ready tool would need to be capable of instrumenting either native binaries or a shared intermediate format.
-One promising solution is the Krabcake (Klock and Garza, 2023) project, which aims to extend Valgrind (Nethercote and Seward, 2007) with a plugin for detecting aliasing violations.
-Valgrind provides a unique advantage compared to other dynamic analysis platforms, since it can instrument native libraries without recompilation.
-This solution would also be appealing for a significant portion of our respondents, since 44% had used Valgrind.
-However, Valgrind incurs significant runtime overhead, so it may also be worthwhile to explore developing a similar tool using LLVM’s sanitizer APIs.
-Multi-language applications also present a challenge for static verification tools.
-Several verifiers rely on Rust’s ownership model to provide aliasing information that can simplify the process of specifying higher-order, functional properties (Lehmann et al., 2023; Astrauskas et al., 2019).
-However, this is only effective when unsafe code is rare.
-Multi-language applications are exposed to significant sources of unsafety, since C and C++ lack static aliasing restrictions.
-Our interview participants would benefit from static analysis and inference tools to ensure that static verification remains usable in this context.
-Techniques such as bi-abduction (Calcagno et al., 2011) would be useful for inferring specifications from unannotated foreign libraries.
-Contributors to Clang have already proposed a form of lifetime inferences for C++, which would provide annotations that could be used to generate Rust encapsulation with equivalent lifetime semantics (Brænne et al., 2022).
-Similar analyses would be helpful for determining if foreign types can implement the Send or Sync traits and to warn when functions retain copies of pointers derived from reference-type arguments.
-Sound Verification
-Despite the need for guarantees of soundness, only 10 survey respondents had used formal methods tools.
-One possible explanation for the lack of responses to this question is that only Verus (Lattuada et al., 2023) and Kani (VanHattum et al., 2022) had any significant support for unsafe code when we distributed the survey, and Verus had been published only several months prior.
-Since then, Verus has added support for programs that manipulate raw pointers, and both Gillian-Rust (Élie Ayoun et al., 2024) and RefinedRust (Gäher et al., 2024) have provided alternative methods of verification that specifically target unsafe code.
-However, none of these tools use formal models that are directly compatible with the aliasing models implemented in Miri, which was used by the majority of our respondents.
-To ensure widespread adoption, verification tools could build on the methods and best practices that the majority of our participants already used.
-Improvements to Miri are likely to be the most effective way to reach our audience.
-For example, if Miri was extended to support symbolic execution, it provide up to 61% of our survey respondents with guarantees of soundness without requiring them to learn to use a new tool.
-The Seer (Renshaw and Canciani, 2018) symbolic execution engine was capable of this at one point, but it was created before Miri had support for detecting aliasing violations (Jung et al., 2019), and the project appears to have been abandoned.
-It would be worthwhile for the Rust community to resurrect this project or implement an equivalent analysis within Miri.
-Additionally, certain approaches to static verification will inevitably require developers to write specifications, which can be burdensome.
-However, 90% (6.5) of our respondents who exposed unsafe APIs claim to have already provided informal specifications as comments or documentation.
-There may be a niche for a lightweight, formal specification language targeted at annotating unsafe APIs with their core safety properties.
-Education
-Only 6% (6.5) of survey respondents indicated that the Rust community’s guidance and documentation were always enough to determine how to use unsafe code correctly.
-While the Rustonomicon (Rust Community, 2023b) is an invaluable entry point for learning about Rust’s unsafe features, we expect that its one-size-fits-all approach will inevitably leave certain users wanting more targeted information.
+(McCormack et al., 2025b) were able to detect errors in foreign function bindings by comparing the size of types in the Rust declarations with the size of types in the LLVM definitions.
+The authors implemented this as part of a dynamic analysis tool, but their method did not depend on dynamic information, so it could be factored out as a static analysis pass.
+Changes in Rust’s newest 2024 edition indicate that Rust will continue to evolve toward improving contract visibility at foreign APIs.
+Rust’s extern blocks now require unsafe annotations (Rust Community, 2024), indicating that developers are responsible for ensuring that foreign function bindings match their definitions.
+Another participant tried to avoid manually reasoning about the lifetime semantics of foreign APIs “in hopes that in future, there will be a better method for doing this correctly” (P4).
+In 2022, contributors to Clang proposed adding lifetime annotations to C++, which would be inferred through a whole-program static analysis pass (Brænne et al., 2022).
+The particular RFC has not had activity since August 2023, but the project was associated with Google’s Crubit analysis tool, which is still in development as of May 2025 (Google, 2025).
+Despite their uncertainty, few of our participants had tried using static verification tools.
+One possible explanation is that when we distributed the survey, only Verus (Lattuada et al., 2023) and Kani (VanHattum et al., 2022) had significant support for verifying unsafe code.
+Since then, both of these tools have added extensive support for unsafe features.
+Verus has been used to verify Vest (Cai et al., 2026), a parser combinator library, as well as several other security-critical application components (Verus Team, 2025).
+Both Kani and Verifast (Jacobs et al., 2025) are being used to verify the entirety of Rust’s standard library as part of a joint initiative between Amazon and the Rust Foundation (Kumar, 2024; Rust Community, 2025g).
+We expect that these new tools will meet our participants’ needs for stronger guarantees of soundness.
+For example, two of our interview participants indicated that refinement typing could be useful—“custom subtyping, so that you can specify I want an integer, but only between one and ten” (P4).
+Flux (Lehmann et al., 2023), a refinement-type checker for Rust, was published concurrent to our study and has since been used to verify several real-world Rust applications (Lehmann et al., 2025; Rindisbacher, 2025).
+As verification continues to become standardized within the Rust community, it would be worthwhile for future studies to examine how these methods are being deployed in practice.
+7.3. Unsafe Code Guidelines
+The majority of survey respondents found that the Rust community’s guidance and documentation were helpful, but only 6% (10) indicated that these resources were always enough to determine how to use unsafe code correctly.
+Interview participants also expressed uncertainty about Rust’s semantics.
+Some were uncertain about the interaction between Rust’s panic!
+macro and Drop implementations, with one participant indicating that “real guidelines for library authors” would be helpful in this area.
+This has been a subject of active discussion since at least 2021 (Jung, 2025), but has yet to be fully documented in community resources (Huss, 2025; Jung, 2025).
+We expect that other, similar topics lack formal documentation—instead, developers need to examine years-long issue threads within repositories for Rust’s compiler and Unsafe Code Guidelines (Rust Community, 2023b).
 We agree with Höltervennhoff et al.
-(Höltervennhoff et al., 2023) that one of the most impactful ways to improve Rust’s documentation would be to provide a dedicated resource that succinctly summarizes points of contention about Rust’s semantics.
-Future interventions should also consider that necessity, performance, and ergonomics can each contribute to a single use case for unsafe code.
-Understanding these motivations may be helpful for effectively communicating safe alternatives.
-For example, a profiling tool that integrates information on the distribution and coverage of unsafe code could make it easier to perform a cost-benefit analysis on the tradeoffs between safety and performance.
-We also expect that lints or static analysis tools could help developers eliminate unnecessary uses of unsafe code, since 88% (6.5) of participants only sometimes refactored their code to remove unsafe features.
-However, refactoring may not have been necessary for these participants if they had only minimally used unsafe code.
+(Höltervennhoff et al., 2023) that developers would benefit from having a dedicated resource summarizing points of contention about Rust’s semantics.
+However, Rust is still a relatively new systems programming language, and certain feature interactions have yet to be fully explored or agreed-upon to the point where writing adequate documentation would even be possible.
+Developers who implement design patterns that sit at the cutting-edge of Rust’s evolving semantics will inevitably be uncertain if their implementations are sound.
+Even so, it would be worthwhile for the Rust team to continue to review popular, long-standing issue threads and flag topics that have not been fully documented in the Rustonomicon (Rust Community, 2026b) or the Rust Reference (Rust Community, 2023a).
+Meanwhile, developers who engage with unsafe code should periodically review relevant issues within the Unsafe Code Guidelines repository to ensure that they are using these features correctly and to have the opportunity to inform compiler developers about their use-cases.
+However, we still expect that improvements to Miri will have the greatest impact on our survey population, since it was used by a majority of survey respondents, and it serves as somewhat of an executable specification for Rust.
 8. Conclusion
@@
 These features can introduce bugs and security vulnerabilities if they are used incorrectly.
-Few Rust-specific development tools provide comprehensive support for unsafe code, and even fewer support multi-language applications.
+Few Rust-specific development tools provide comprehensive support for unsafe code, and even fewer support multilanguage applications.
 Developers will need new solutions for testing, bug finding, and verification in this context.
-We conducted a mixed-methods, exploratory study to determine which interventions will be effective for developers who use unsafe code.
+We conducted a mixed methods, exploratory study to determine which interventions will be effective for developers who use unsafe code.
 We interviewed 19 developers who regularly wrote or edited unsafe code and created a survey that reached an additional 160 developers who engaged with unsafe code in some capacity.
@@
 Certain foreign libraries had aliasing and concurrency patterns that conflicted with the expectations of Rust’s type system, making them difficult to encapsulate.
-However, participants did not have a reliable solution for determining if their encapsulations were sound.
-Miri, a Rust interpreter, was the de facto solution for detecting undefined behavior, but participants who had used it before were deterred from using again it due to its slow performance and lack of support for key features.
+However, participants did not have a reliable method to determine if their encapsulations were sound.
+Miri, a Rust interpreter, was the most popular solution for detecting undefined behavior, but participants who had used it before were deterred from using it again because of its slow performance and lack of support for key features.
 Most participants were motivated to use unsafe code by necessity, but they also cited performance and ergonomics as motivations.
-The majority followed best practices such as encapsulating and documenting unsafe code, but many were somewhat uncertain about their design patterns, and few felt that Rust’s documentation was always adequate for understanding how to use unsafe code correctly.
-To resolve this uncertainty, our participants will need verification and analysis tools that are capable of providing guarantees of soundness and have robust support for multi-language applications.
-We expect that tools which integrate with Miri or use preexisting documentation or can generate specifications from documentation will be uniquely effective at reaching the majority of our audience.
-Any interventions should consider that developers may be motivated by a combination of necessity, performance, and ergonomics.
+The majority followed best practices, such as encapsulating and documenting unsafe code, but many were somewhat uncertain about their design patterns, and few felt that Rust’s documentation was always adequate for understanding how to use unsafe code correctly.
+Our participants would benefit from new static and dynamic analysis tools to assist with interoperation and with finding Rust-specific forms of undefined behavior.
+Dynamic analysis tools could improve on Miri to provide better performance, support for inline assembly, and the capability to detect aliasing violations triggered by operations that span foreign function boundaries.
+The Rust community should also continue to invest in improving and expanding documentation on unsafe code, so that longstanding issue threads can become integrated into standard community resources.
+A useful direction for future work would be to synthesize the methods used by Astrauskas et al.
+(Astrauskas et al., 2020) and Evans et al.
+(Evans et al., 2020), which provide different, complementary perspectives on the distribution of unsafe code.
+Future studies should consider the use of unsafe code in unpublished but publicly available crates, as well as the extent to which Rust libraries are exposed to unsafe operations through foreign function calls.
+The results of such a study would provide an excellent opportunity to triangulate the results described in this paper.
+Acknowledgements.
+We thank the Rust Community for their engagement and support, as well as Jenny Liang and Courtney Miller for their feedback on our study design.